### v1

# Testing identifying assumptions in Tobit Models

Abstract.
This paper develops sharp testable implications for Tobit and IV-Tobit models’ identifying assumptions: linear index specification, (joint) normality of latent errors, and treatment (instrument) exogeneity and relevance. The new sharp testable equalities can detect all possible observable violations of the identifying conditions. We propose a testing procedure for the model’s validity using existing inference methods for intersection bounds. Simulation results suggests proper size for large samples and that the test is powerful to detect large violation of the exogeneity assumption and violations in the error structure. Finally, we review and propose new alternative paths to partially identify the parameters of interest under less restrictive assumptions.
Keywords: Tobit models, hypothesis testing, Testable Implications, Instrumental variables.
JEL subject classification: C12, C24, C26, C34.
1. Introduction
Since the seminal work of Tobin (1958), Tobit models have earned attention in economics, business and social sciences. [^1] According to Google Scholar Tobin’s original paper has more than 10000 citations. Tobin (1958) analyzed household expenditure on durable goods using a regression model that specifically incorporated that expenditure (the dependent variable) cannot be negative. This approach is related to a broader class of models known as censored or truncated regression models depending on whether observations outside a specified range are lost or censored. When applied researchers are interested in modelling limited dependent variables, potentially with mass accumulation points, the Tobit family of models provides structure to identify parameters of interest, such as the average treatment effect (ATE). Identification relies on three primary sources: (i) instrument exogeneity (or exogeneity of the variable of interest itself), (ii) normality of the model’s latent variables, and (iii), in the case of an instrumental variable approach to endogeneity, the relevance condition for the instrument. While researchers recognize the model’s restrictive nature, it remains a valuable tool in the empirical literature.
In this paper, we provide a test for the validity of the Tobit model’s structure and assumptions, providing three main contributions to the literature. The first is to provide sharp testable equalities that can detect all possible observable violations of the Tobit model. Second, we propose a test for the validity of the Tobit model’s identifying assumptions using the sharp equalities that characterize the model to check its falsifiability. Following recent literature, we convert the equalities into conditional moment inequalities, and implement the test by existing inferential methods from Chernozhukov, Lee, and Rosen (2013). The third contribution is to review and propose alternative approaches that can be used when the model is rejected. We explore an alternative path to partially identify the parameter of interest by assuming monotonicity of selection into treatment. Finally, we provide empirical examples illustrating the methodology and its practical relevance. More generally, the current paper contributes to the growing literature on testing identifying assumptions of econometric models.
We focus on two main models: (i) the “classic Tobit” model in which the main variable of interest is assumed to be exogenous and (ii) the IV-Tobit model. In both cases, the proposed test considers all observable violations of the model structure, such as linear index, normality of the latent errors, independence of the treatment, and homoskedasticity; with the addition of the validity of the instrument for the IV-Tobit. Additional results for variants of the Tobit family of models (e.g., type 2 Tobit) are presented in the Appendix.
1.1. Previous Literature
There is a vast literature related to testing the validity of Tobit models and their assumptions, to which we contribute. Most of the preceding work focuses on testing a particular assumption or feature of the model while maintaining other assumptions and structures as valid. Nelson (1981) constructs a Hausman-type test for misspecification of the classic Tobit model (that is, normality, linear index, and homoskedasticity) where the maximum likelihood estimates are compared with method of moment estimates. Nelson’s test compares the sample proportion of non-censored observations with the hypothesized probability of being non-censored in the Tobit model. As stated by Bera, Jarque, and Lee (1984), Nelson (1981) test is equivalent to the Lagrange multiplier test of the Tobit model against Cragg’s model (Lin and Schmidt, 1984). Bera, Jarque, and Lee (1984) proposes a Lagrange Multiplier test for the normality assumption against other distributions of the Pearson family of distributions while the remaining assumptions are maintained. Newey (1987) considers both exogenous and endogenous explanatory variables cases using symmetrically censored least squares estimators to construct specification tests of normality and homoskedasticity assumptions via a Hausman-type specification test. Holden (2004) examines several statistics proposed to test the normality assumption in the Tobit (censored regression) model and reinterpreted them as a version of the Lagrange multiplier (score) test for a common null hypothesis. Reynolds and Shonkwiler (1991) use an information matrix misspecification test to detect violations of the distributional assumptions of the Tobit model.
Other tests include Drukker (2002), which operationalized conditional moment tests developed by Newey (1985) and Tauchen (1985) to the case of misspecification of the distribution of the classic Tobit model. With a similar intuition to our framework, their test writes down conditional moment restrictions which should have zero conditional expected values under the null. The intuition behind the conditional moments test developed by Drukker (2002) is that since the model was estimated by maximum likelihood, the assumed data-generating process specifies all moments of disturbances conditional on the covariates to be the ones of a normal distribution. Drukker (2002) use these moment-based methods based on the third and fourth moments of the normal distribution. Our procedure, while intuitively similar detects all possible violations of the model, not only those evident from deviations in the third and fourth moments. Smith and Blundell (1986) propose a test of the treatment variable’s exogeneity in the IV-Tobit Model by a control function approach exploiting the joint normality of the latent variables.
Most of these approaches focus on testing a particular subset of assumptions, or a specific class of alternatives to consider. We consider the full set of assumptions with a general alternative hypothesis, which is not a specific subset of models.
The developments proposed in this paper consider all observable violations of the general Tobit model structure and its assumptions, serving as a useful test for empirical researchers constructing models for censored/truncated data. This work contributes to the growing literature on the testability of the identifying assumptions in various econometric models. Those include tests for the validity of assumptions in instrumental variable models, local average treatment effects, regression discontinuity designs, bivariate probit, etc (Mourifié and Wan, 2017; Kitagawa, 2015; Huber and Mellace, 2015; Gunsilius, 2021; Arai et al., 2022; Acerenza, Bartalotti, and Kédagni, 2023). Our testing procedures are connected to kedagni2020generalized which provides tests for assumptions related to the instrumental variables in the model, while we also consider the implications of the parametric structure on the outcome and latent error structure which are inherent to Tobit models.
The remainder of the paper is organized as follows. Section 2 presents the model and identifying assumptions. Section 3 derives the sharp testable implications. Section 4 discusses heuristically the identification of the model’s parameters. Section 5 outlines the testing procedure. Section 6 include simulation evidence about the test’s size and power. Section 7 discusses how to relax the assumptions in case of rejection. Section 8 and 9 provide empirical illustrations. Finally, Section 10 concludes. Additional results and details are collected in Appendices A-E.
2. Models
This section describes two popular “two-part” models for truncated data for which we derive testable implications in Section 3: the classic Tobit, and the instrumental variable Tobit. [^2] We omit conditioning on other exogenous covariates $X$ for simplicity of exposition.
Let $Y=\max(0,Y^{*})$ be the observed outcome taking values in $\mathcal{Y}\subset\mathbb{R}^{+}$ with $Y^{*}$ a latent continuous dependent variable taking values in $\mathcal{Y^{*}}\subset\mathbb{R}$. Both the treatment of interest, $D$, and the instrumental variable, $Z$, can be discrete or continuous and take values in $\mathcal{D}\subset\mathbb{R}$ and $\mathcal{Z}\subset\mathbb{R}$, respectively. We normalize the coefficients by the standard deviation of the convenient unobservable structure for the exposition. [^3] Identification for the normalized and non-normalized parameters is discussed in section 5.
2.1. Classic Tobit model
The classic Tobit model considers the case in which the researcher is interested in the effect of an exogenous treatment on a non-negative outcome that has a mass point at zero:
$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ \\ Y^{*}&=&\alpha_{0}+\alpha_{1}D+U\end{array}\right.$ (2.4)
Where $U$ is an unobservable (latent) error. In addition to the model structure in Equation 2.4, the classic Tobit also considers restrictions on the distribution of the error term and its relationship to the treatment.
Assumption 1.
$D$ is independent of $U$.
Assumption 2.
$U$ is distributed as $N(0,1)$.
Assumption 1 states that the treatment of interest is independent of the model’s unobservables. Assumption 2 imposes that the latent error has a standard normal distribution. [^4] The error could take $N(0,\sigma^{2})$ but after scale normalizations of the coefficients of $Y^{*}$ we can model it as $N(0,1)$.
Despite its simplicity, the classic Tobit model can accommodate treatment that is continuous or discrete, and the normality assumption for the latent error could be replaced by other distributional assumptions that preserve the identification of the parameters.
2.2. IV Tobit model
The IV Tobit considers the case in which the treatment of interest is endogenous, but an instrumental variable is available to identify its effect. Let the IV Tobit model be: [^5] Results for the IV-Tobit model are presented for continuous $D$ but can accommodate a binary treatment, $D=1\{\gamma_{0}+\gamma_{1}Z-V>0\}$. In that case, the parameters are identified only up to scale.
$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ \\ Y^{*}&=&\alpha_{0}+\alpha_{1}D+U\\ \\ D&=&\gamma_{0}+\gamma_{1}Z+V\end{array}\right.$ (2.10)
Alternatively, in its reduced form representation:
$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ \\ Y^{*}&=&\beta_{0}+\beta_{1}Z+W\\ \\ D&=&\gamma_{0}+\gamma_{1}Z+V\end{array}\right.$ (2.16)
Where $U,V$ are the latent structural error terms. To that model structure, we add restrictions on the joint distribution of the latent variables and their relationship to the instrumental variable.
Assumption 3.
Assume that: $Z$ is independent of $(U,V)$ and $\gamma_{1}\neq 0$.
Assumption 4.
Let $U,V$ follow a bivariate normal distribution with covariance $\rho$, i.e., $\begin{pmatrix}U\\ V\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma=\begin{pmatrix}\sigma^{2}_{U}&\rho_{UV}\\ \rho_{UV}&\sigma^{2}_{V}\end{pmatrix}$.
Assumption 3(i) states that the instrumental variable is independent of the model’s structural latent variables. The second part of Assumption 3 is the usual instrument relevance in determining the treatment. This assumption also implies that $Z$ is independent of $W,V$, since $W=\alpha_{1}V+U$.
Assumption 4 characterises the distribution of the latent vector of error terms as bivariate normal. Then, $W,V$ also follows a bivariate normal distribution and $\alpha_{0},\alpha_{1},\gamma_{0},\gamma_{1}$ could be scale normalized so $(W,V)^{\prime}$ follows the standard bivariate normal distribution with covariance $\rho$, i.e., $\begin{pmatrix}W\\ V\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\end{pmatrix}$, and $\Sigma=\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}$. [^6] Note that if $U,V$ follow a bivariate normal distribution with covariance $\rho$, i.e., $\begin{pmatrix}U\\ V\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma=\begin{pmatrix}\sigma^{2}_{U}&\rho_{UV}\\ \rho_{UV}&\sigma^{2}_{V}\end{pmatrix}$, then $W,V$ follow a bivariate normal distribution $\begin{pmatrix}W\\ V\end{pmatrix}\sim\mathcal{N}(\mu_{1},\Sigma_{1})$, where $\mu_{1}=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma_{1}=\begin{pmatrix}\alpha_{1}^{2}\sigma^{2}_{V}+\sigma^{2}_{U}+2\rho_{UV}\alpha_{1}\sigma_{V}\sigma_{U}&\alpha_{1}\sigma_{V}^{2}+\rho_{UV}\\ \alpha_{1}\sigma_{V}^{2}+\rho_{UV}&\sigma^{2}_{V}\end{pmatrix}$. Then the reduced form system can be scale normalized. The latter characterisation aids the development of the sharp testable implications.
3. Sharp testable equalities
This section presents the sharp testable equalities for the models described in Section 2. Deviations from these equalities imply violations of the “null hypothesis” of the Tobit model (linear latent index, independence, instrument validity, and normality). The equalities are conditional on parameters being identified under the model’s assumptions, and a discussion on identification is postponed to Section 4.
3.1. Classic Tobit model
A defining characteristic of Tobit and similar models is the mass accumulation at zero for the distribution of the non-negative outcome, $Y$. Thus, the model’s testable implications require characterizing the distribution at both the mass point and beyond it.
Starting at the continuous part of the distribution of $Y$, the conditional probabilities of $c_{0}\leq Y\leq c_{1}$ for a $c_{1},c_{0}>0$ are observed. For any value of the treatment variables $d\in\mathcal{D}$:
$P(c_{0}\leq Y\leq c_{1}|D=d)$ $=$ $P(c_{0}\leq Y\leq c_{1},Y^{*}\geq 0|D=d)$ (3.1)
$+$ $P(c_{0}\leq Y\leq c_{1},Y^{*}<0|D=d)$
$=$ $P(c_{0}\leq Y\leq c_{1},Y^{*}\geq 0|D=d)$
$=$ $P(c_{0}\leq Y^{*}\leq c_{1},Y^{*}\geq 0|D=d)$
$=$ $P(c_{0}\leq Y^{*}\leq c_{1}|D=d)$
$=$ $P(c_{0}-\alpha_{0}-\alpha_{1}d\leq U\leq c_{1}-\alpha_{0}-\alpha_{1}d|D=d)$
$=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)$
where $\Phi(.)$ is the standard normal CDF. The first equality follows from the law of total probability. The second through fourth equalities follow from the model’s structure described in Equation 2.4, namely $P(Y>0,Y^{*}<0|D=d)=0$ and $Y^{*}=Y$ for $Y^{*}>0$. The fifth one is given by the latent linear model structure of $Y^{*}$, and finally, the last step follows from assumptions 1 and 2. To cover all possible cases of positive outcomes, note that for any $c_{1}>0$:
$P(c_{1}\leq Y|D=d)$ $=$ $1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d).$ (3.2)
Turning to the accumulation point, the observed event of $Y=0$,
$P(Y=0|D=d)=P(Y^{*}\leq 0|D=d)=1-\Phi(\alpha_{0}+\alpha_{1}d).$ (3.3)
The first equality follows from the model’s structure in Equation 2.4, and the second one applies assumptions 1 and 2. The equalities described in equations 3.1-3.3 fully characterize the distribution of $Y$ conditional on $D$, connecting the probabilities in the observed data to those implied by the Tobit model.
Remark 1 (Sharpness).
In the context of model 2.4, equalities 3.1-3.3 are sharp in the sense that whenever they hold, it is possible to construct a vector of $(\tilde{Y},D,\tilde{U})$ that satisfies model 2.4, Assumptions 1 and 2, and induces the observed distribution on the data $(Y,D)$. For the proof see Appendix A.
Then, the equalities 3.1-3.3 are sharp testable implications for the validity of the classic Tobit Model in 2.4 coupled with assumptions 1-2. They serve as the basis for the test procedure described in Section 5.
Remark 2 (Non-learnability).
The testable implications and sharpness discussed above show that the classic Tobit model is generally refutable. However, the model is non-verifiable. To prove this, we show that for any joint distribution on the data $(Y,D)$ that satisfies equalities 3.1-3.3, there exists a vector of $(\bar{Y},D,\bar{U})$ that induces the observed distribution on the data $(Y,D)$, but the vector $\bar{U}$ is not normally distributed. See Appendix E for the proof.
Remark 3 (Extensions).
The previous derivation can be adjusted for different variations of Tobit models, such as generalizations of the distributional assumptions (barros2018generalized), different thresholds (carson2007Tobit), dynamic Tobit models, or including individual specific effects (wooldridge2005simple; honore2000estimation; honore1993orthogonality). In appendix C we derive the equalities for the type 2 Tobit model and similar logic can be applied to other two-part models. Additional testable results for the aforementioned models are discussed in Appendix D.
3.2. IV Tobit model
We turn our attention to the Instrumental Variable Tobit case and propose testable implications that can be used to test the model described in 2.10 and the associated assumptions 3 and 4.
The observed data includes $(Y,D,Z)$, and we characterise the joint distribution of $(Y,D)\in\mathcal{Y}\times\mathcal{D}$ conditional on the instrument, $Z$, to obtain the model’s testable implications. The mapping from observed probabilities to its corresponding quantities implied by the IV Tobit model requires jointly evaluating the continuous (interior) support for the outcome and treatment variables, as well as the accumulation points and distribution tails conditional on $Z$. Consider any $0<c_{0}<c_{1},d_{0}<d_{1}$, and $z\in\mathcal{Y}\times\mathcal{D}\times\mathcal{Z}$,
$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.4)
-- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
$\equiv$ $\Phi^{1}_{c_{1},c_{0},d_{1},d_{0}},$
$P(c_{0}\leq Y\leq c_{1},D\leq d_{0}|Z=z)$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.5)
-- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
$\equiv$ $\Phi^{1}_{c_{1},c_{0},d_{0}},$
$P(c_{0}\leq Y\leq c_{1},D\geq d_{1}|Z=z)$ $=$ $\Phi_{W}(c_{1}-\beta_{0}-\beta_{1}z)-\Phi_{W}(c_{0}-\beta_{0}-\beta_{1}z)$ (3.6)
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
$\equiv$ $\Phi^{1}_{c_{1},c_{0},d_{1}},$
$P(c_{1}\leq Y,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z;\rho)-\Phi_{V}(d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.7)
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
$+$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
$\equiv$ $\Phi^{1}_{c_{1},d_{1},d_{0}},$
$P(c_{1}\leq Y,D\leq d_{0}|Z=z)$ $=$ $\Phi_{V}(d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.8)
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
$\equiv$ $\Phi^{1}_{c_{1},d_{0}},$
$P(c_{1}\leq Y,D\geq d_{1}|Z=z)$ $=$ $1-\Phi_{W}(c_{1}-\beta_{0}-\beta_{1}z)-\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.9)
$+$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
$\equiv$ $\Phi^{1}_{c_{1},d_{1}}.$
where $\Phi_{W,V}(w,v;\rho)$ denotes the joint c.d.f. of $(W,V)$, a bivariate normal with coefficient of correlation $\rho$. Similarly, considering the case that the observed $Y$ equals zero,
$P(Y=0,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.10)
-- $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)\equiv\Phi^{2}_{d_{1},d_{0}},$
$P(Y=0,D\leq d_{0}|Z=z)$ $=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)\equiv\Phi^{2}_{d_{0}},$ (3.11)
$P(Y=0,D\geq d_{1}|Z=z)$ $=$ $\Phi_{W}(-\beta_{0}-\beta_{1}z)-\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.12)
$\equiv$ $\Phi^{2}_{d_{1}}.$
Remark 4 (Sharpness).
In the context of model 2.10 or 2.16, equalities 3.4-3.12 for all $c,d,z\in\mathcal{Y}\times\mathcal{D}\times\mathcal{Z}$ are sharp in the sense that whenever they hold, it is possible to construct a vector of $(\tilde{Y},\tilde{D},\tilde{U},\tilde{V},Z)$ or equivalently $(\tilde{Y},\tilde{D},\tilde{W},\tilde{V},Z)$ that satisfies model 2.10 and 2.16, Assumptions 3 and 4, and induces the observed distribution on the data $(Y,D,Z)$. See Appendix B for the proof.
Then, the equalities 3.4-3.12 are sharp testable implications for the validity of the Instrumental Variable Tobit Model in 2.10 coupled with assumptions 3-4. They serve as the basis for the test procedure described in Section 5. A similar claim of non-learnability holds for the iv-tobit. The demonstration follows a similar logic.
4. Identification
In this section, we provide heuristic arguments for the identification of the parameters of the respective models, expanding on Han and Vytlacil (2017) and Acerenza, Bartalotti, and Kédagni (2023).
4.1. Classic Tobit model
Equation (3.3) established that in the classic Tobit model,
$1-P(Y=0|D)=\Phi(\alpha_{0}+\alpha_{1}D).$
Since the standard normal CDF, $\Phi(.)$, is monotonic, we use its inverse to obtain:
$\Phi^{-1}(1-P(Y=0|D))=\alpha_{0}+\alpha_{1}D.$
Thus, $\alpha_{1}=\frac{Cov(\Phi^{-1}(1-P(Y=0|D)),D)}{Var(D)}$, and $\alpha_{0}=E(\Phi^{-1}(1-P(Y=0|D)))-\alpha_{1}E(D)$.
The coefficients have been normalized by dividing by the latent variable’s ($U$) square root of the variance, denoted by $\sigma$. Naturally, this implies that the ratio of the original parameters, say $\tilde{\alpha}_{0},\tilde{\alpha}_{1}$ and $\sigma$ is identified since $P(Y=0|D=d)=\Phi(\alpha_{0}+\alpha_{1}D)=\Phi(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D)$.
Similarly to the case analysed by Han and Vytlacil (2017), this allows the identification of the inverse mills ratio which is a known function of the standard normal density, $\lambda(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D))$. Then,
$E(Y|D,Y>0)=\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D+\sigma\lambda(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D))$
Which in turn implies:
$Y=\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D+\sigma\lambda(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D))+e$
with $e$ mean independent of $D,\lambda(\cdot)$. Since $D$ is known and $\lambda(\cdot)$ was already identified then by considering the moment conditions for $Cov(.,D)$, $Cov(.,\lambda(\cdot))$ and $E(.)$ we get a three by three linear system with a unique solution for $\tilde{\alpha}_{0},\tilde{\alpha}_{1},\sigma$. In other words, the normality assumption allows identification by leveraging the relationship of the variable of interest and the outcome at the accumulation point ($Y=0$) to obtain the scaled parameters, and the same relationship when $Y>0$ to overcome the normalization. However, the reliance on the normality and linearity assumptions underlines the importance of adequately testing the model’s structure and assumptions in empirical applications.
4.2. IV Tobit model
Turning the focus to the IV-Tobit model, the first stage is readily identified from the linear structure, $\gamma_{1}=\frac{Cov(Z,D)}{Var(Z)}$, $\gamma_{0}=E(D)-\gamma_{1}E(Z)$. Now note that calculating $P(Y=0|Z)$:
$1-P(Y=0|Z)=\Phi(-\beta_{0}-\beta_{1}Z)$
Similarly to subsection 4.1, by inverting the normal CDF we obtain $-\beta_{1}=\frac{Cov(\Phi^{-1}(1-P(Y=0|Z)),Z)}{Var(Z)}$, $-\beta_{0}=E(\Phi^{-1}(1-P(Y=0|Z)))+\beta_{1}E(Z)$. The relationship between $\beta_{0},\beta_{1},\gamma_{0},\gamma_{1}$ identifies $(\alpha_{0},\alpha_{1})$. Finally, to identify $\rho$, let $s=c-\beta_{0}-\beta_{1}z$ and $t=d-\gamma_{0}-\gamma_{1}z$. Then, by using equation 3.7:
$P(Y\leq c,D\leq d|Z=z)$ $=$ $\Phi_{W,V}(s,t;\rho)$
Which by classic results of bivariate normal random variables if we differentiate with respect to $\rho$ is:
$\frac{\partial P(Y\leq c,D\leq d|Z=z)}{\partial\rho}$ $=$ $\phi_{W,V}(s,t;\rho)$
Where $\phi_{a,b}$ is the bivariate normal probability density, which is positive for any $s,t$. Hence, $\Phi_{W,V}(.,.;\rho)$ is monotonic in $\rho$ and thus invertible. So $\rho$ can be identified.
Recall the previous coefficients are normalized, a two-step procedure can be followed to identify the full original covariance matrix and the reduced form covariance matrix (Wooldridge, 2010, p. 683-84).
5. Testing procedure
To test the sharp equalities, we rewrite each of these equalities as two inequalities (Mourifié and Wan, 2017; Acerenza, Bartalotti, and Kédagni, 2023). For a concrete example, we first note that for example
$P(c_{1}\leq Y|D=d)=1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)\iff E[1\{c_{1}\leq D\}-1+\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]=0.$
We then we rewrite each of these moment equalities implied by the restrictions on the empirical distribution as moment inequalities
$E[1\{c_{1}\leq D\}-1+\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]\leq 0,E[-1\{c_{1}\leq D\}+1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]\leq 0.$
In doing this, for all the restrictions on the empirical distribution we can implement a test relying on existing intersection bounds inferential methods such as Chernozhukov, Lee, and Rosen (2013) which is specifically suited to test conditional moment inequalities.
Note that the equalities for the classic Tobit need to hold for any pair of constants $(c_{0},c_{1})$, and the ones from the IV Tobit need to hold for pairs of $(c_{0},c_{1})$ and $(d_{0},d_{1})$. We propose a partition of $\mathcal{Y}\times\mathcal{D}$ to test sufficient conditions of these sharp equalities. Rejection of any of the null hypotheses that the equalities hold at these particular levels will imply violations of the model.
Classic Tobit model
The sharp testable equalities for every $c_{0},c_{1}\in\mathcal{Y}$ are given by equations 3.1 and 3.3. For a partition of the support of $Y$ into $K$ arbitrary chosen sets $C_{k}=(0,c_{k})$ such that $C_{k}\in C_{k+1}$, the following set of sufficient inequalities are, for some chosen values of $c_{k},c_{k+1}\in\mathcal{Y}$, related to the components of 3.1-3.3.
The formulation of the inequalities considered will depend on each partition’s location on the support of the outcome variable $Y$. Let $c_{1}=0$ and $W_{k}$ be
$W_{k}=\left\{\begin{array}[]{cc}1\{Y=0\}-(1-\Phi(\alpha_{0}+\alpha_{1}D)),&\text{ if }k=0\\ 1\{c_{k}<Y\leq c_{k+1}\}-\Phi(c_{k+1}-\alpha_{0}-\alpha_{1}D)+\Phi(c_{k}-\alpha_{0}-\alpha_{1}D),&\text{ if }1\leq k<K\\ 1\{c_{K}<Y\}-(1-\Phi(c_{K}-\alpha_{0}-\alpha_{1}D)),&\text{ if }k=K\\ \end{array}\right.$
The intersection bounds framework considers the following $2(K+1)$ inequalities
$\sup_{d}E[W_{k}|D=d]$$\leq$ 00 0
$\sup_{d}E[-W_{k}|D=d]$$\leq$ $0,\text{ for }k=0,\dots,K$
We can write more compactly as
$\max_{k}\sup_{D}\theta_{k}(D)\leq 0$ (5.2)
where $\theta_{k}(D)$ collects all the inequalities being tested. The decision rule for the test is given by Chernozhukov, Lee, and Rosen (2013), we reject $H_{0}$ if
$\hat{\theta}_{1-\alpha}\equiv\max_{k}\sup_{D}\left\{\hat{\theta}(D,k)-\kappa_{1-\alpha}\hat{s}(D,k)\right\}>0,$ (5.3)
where $\hat{\theta}(D,k)$ is a nonparametric estimator for $\theta(D,k)$, $\hat{s}(D,k)$ its standard error, and $\kappa_{1-\alpha}$ is a critical value at the significance level $\alpha$.
IV Tobit model
For the instrumental variable Tobit model, the continuous support for both the outcome and treatment poses challenges to the implementation of the test. The sharp testable equalities for every $c_{0},c_{1}\in\mathcal{Y}$ and $d_{0},d_{1}\in\mathcal{D}$ are given by equations 3.4-3.12.
Consider a partition of the support of $Y$ into $K$ arbitrary chosen sets $C_{k}=(0,c_{k})$ such that $C_{k}\in C_{k+1}$ and of the support of $D$ into $Q$ arbitrary chosen sets $D_{q}=(0,d_{q})$ such that $D_{q}\in D_{q+1}$, the following set of sufficient inequalities are related to the components of (3.4)-(3.12). Analogous to the classic Tobit case, the formulation of the inequalities considered will depend on each partition’s location on the joint support of the outcome and treatment variables. Let $W_{kq}$ be given by,
$W_{kq}=\left\{\begin{array}[]{cc}1\{Y=0\}1\{D\leq d_{0}\}-\Phi^{2}_{d_{0}},&\text{ if }k=0,q=0\\ 1\{Y=0\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{2}_{d_{q},d_{q+1}},&\text{ if }k=0,1\leq q<Q\\ 1\{Y=0\}1\{D\geq d_{Q}\}|-\Phi^{2}_{d_{Q}},&\text{ if }k=0,q=Q\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{D\leq d_{0}\}-\Phi^{1}_{c_{k+1},c_{k},d_{0}},&\text{ if }1\leq k<K,q=0\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{1}_{c_{k+1},c_{k},d_{q+1},d_{q}},&\text{ if }1\leq k<K,1\leq q<Q\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{D\geq d_{Q}\}-\Phi^{1}_{c_{k+1},c_{k},d_{Q}},&\text{ if }1\leq k<K,q=Q\\ 1\{Y\geq c_{K}\}1\{D\leq d_{0}\}-\Phi^{1}_{c_{K},d_{0}},&\text{ if }k=K,q=0\\ 1\{Y\geq c_{K}\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{1}_{c_{K},d_{q+1},d_{q}},&\text{ if }k=K,1\leq q<Q\\ 1\{Y\geq c_{K}\}1\{D\geq d_{Q}\}-\Phi^{1}_{c_{K},d_{Q}},&\text{ if }k=K,q=Q.\\ \end{array}\right.$
Where we used the simplifying notation defined in equations (3.4)-(3.12) for each partition of the support for $Y$ and $D$. The intersection bounds framework considers the following $2(K+1)(Q+1)$ inequalities
$\sup_{z}E[W_{kq}|Z=z]$$\leq$ 00 0
$\sup_{z}E[-W_{kq}|Z=z]$$\leq$ $0,\text{ for }k=0,\dots,K;q=0,\dots,Q.$
We can write more compactly as
$\max_{k,q}\sup_{Z}\theta_{k}q(Z)\leq 0$ (5.5)
where $\theta_{k}q(Z)$ collects all the inequalities being tested. The decision rule for the test is given by Chernozhukov, Lee, and Rosen (2013), we reject $H_{0}$ if
$\hat{\theta}_{1-\alpha}\equiv\max_{k,q}\sup_{Z}\left\{\hat{\theta}(Z,k,q)-\kappa_{1-\alpha}\hat{s}(D,k,q)\right\}>0,$ (5.6)
where $\hat{\theta}(D,k)$ is a nonparametric estimator for $\theta(D,k)$, $\hat{s}(D,k)$ its standard error, and $\kappa_{1-\alpha}$ is a critical value at the significance level $\alpha$.
To implement the test within the Chernozhukov, Lee, and Rosen (2013) intersection bounds inferential method, we use the CLR Stata package described in Chernozhukov et al. (2015). The parameters in the relevant model, for example, $\beta_{0},\beta_{1},\gamma_{0},\gamma_{1},\rho$, are replaced by their maximum likelihood estimators (MLE), and asymptotic validity of this “plug-in” test follows from the argument described by (Acerenza, Bartalotti, and Kédagni, 2023, Appendix B). Some additional details on the test implementation are discussed in Section 6.
Remark 5.
Intuitively, the testable conditions derived above consider if the empirical conditional distribution of the observed outcome variable - in both the mass accumulation and non-truncated parts of the support of $Y$ is consistent with a random variable(s) following the (bivariate) normal distribution for different sections of the distribution and values of the independent instrument, $Z$.
The proposed test procedures are intended to detect violations of the model due to:
1. Misspecification of the latent structure that makes the coefficient estimates biased as estimates of the true coefficients of $Y^{*}$;
2. Violations arising from the empirical distribution of $Y$ being inconsistent with the implied distributions from the parametric structure (that is if the proportion of residuals in different parts of its support deviate from the normality assumptions);
3. Violations due to the empirical distributions of the residuals differing from the implied distributions in certain values of the treatment (Classic Tobit) or instrument (IV Tobit), which indicate violations of the exogeneity of treatment or instrument (kedagni_generalized_2020).
Remark 6 (On the inclusion of covariates).
The existing procedure can be extended to include exogenous covariates, $X$, within the linear index model in 2.10. Then, the testable equalities for both classic and IV Tobit models could be derived with the additional conditioning on $X$. To implement the test with covariates, we suggest generating a partition of the covariate space, say in $J$ grids, similarly to the partition of the exogenous variable used in Section 5. For every grid in the partition, one could compute the test statistic 5.3 or 5.6 for the classic or the IV-tobit, respectively. Then obtain critical values that account for multiple testing via a Bonferroni correction. In particular, for every grid, set the critical value at the significance level $\frac{\alpha}{J}$, namely $\kappa_{1-\frac{\alpha}{J}}$.
A theoretically interesting approach would partition the joint support of all the exogenous variables and compute the test statistics across all parts of the grid. However, this approach would entail significant implementation challenges, since the statistical packages for Chernozhukov et al. (2015) allow for only one exogenous covariate.
6. Simulations
In this section, we provide simulation exercises for the size and power of the proposed test for the classical Tobit model assumptions.
The testing procedure described relies on testing sharp equalities that should hold for any arbitrary partition. Since both the outcome and treatment are continuous, evaluating all possible equalities is technically challenging. We focus on a non-sharp set of the equalities by evaluating them at different grid partitions of their support, in a similar spirit to honore2020selection. Naturally, this choice makes the test less powerful as we don’t consider the continuum of equalities derived in Section 4, but is justified by the ease of implementation of the procedure based on intersection bounds and the performance of the test on the simulations below.
For the simulations presented below, we considered a partition of the support of the observed outcome variable into the accumulation point ($Y=0$) and four quartiles on the (untruncated) positive range. The choice of the number and location of the partitions/evaluation points balances the implementation computational requirements, availability of data for different parts of the joint support of the outcome, treatment and exogenous instrumental variable. For the procedure in Section 5 to be feasible, we must have data on both the outcome and the exogenous variables within each partition. Using the empirical quantiles of the non-truncated outcomes to determine the partitions guarantees a reasonable number of observations for each part of the grid. Larger sample sizes might allow finer partitions for the outcome support to be used, but the added computational requirements created by an increased number of equalities being checked, coupled with the larger datasets, can substantially increase computing time. [^7] The simulations presented in Section 6 have limited sample sizes and a relatively small number of equalities being tested due to the long running time and computational constraints when repeating the test procedure thousands of times.
When implementing the intersection bounds in STATA using the package clrbound Chernozhukov et al. (2015), the researcher must determine the range of values of the exogenous variable for which each equality will be evaluated. To guarantee the feasibility of the procedure, we adjust the evaluation points for $D$ ($Z$) to the first and $99^{th}$ percentiles of the exogenous variable in each partition of the support for the outcome $Y$ [^8] See details on implementation on the simulation replication STATA code.
6.1. Classic Tobit model
6.1.1. Size
For the classic Tobit model, we consider a continuous treatment $D$ following the data generating process described in Equation 6.8.
$\left\{\begin{array}[]{lcl}Y&=&\max\{Y^{*},0\}\\ Y^{*}&=&D+U\\ D&=&2Z-V\\ \begin{pmatrix}U\\ V\\ Z\end{pmatrix}&\sim&\mathcal{N}(\mathbf{0}_{3},\mathbf{\Sigma})\\ \mathbf{\Sigma}&=&\begin{pmatrix}1&\rho_{uv}&\rho_{uz}\\ \rho_{uv}&1&\rho_{vz}\\ \rho_{uz}&\rho_{vz}&1\end{pmatrix}\\ \rho_{uz}&=&\rho_{vz}=0\\ \rho_{uv}&=&\rho\end{array}\right.$ (6.8)
Where $\mathbf{0}_{p}$ is a $p\times 1$ vector of zeroes. The parameter $\rho$ determines the intensity of dependence between the latent variables jointly determining the treatment and outcome. Under the treatment exogeneity condition described in Assumption 1, $\rho=0$. Table 1 presents the empirical test sizes for this scenario, under different significance levels $\alpha$. The results indicate that while the test over rejects the null hypothesis for small to mid-sized samples, the test’s empirical coverage approaches its desired nominal benchmark as samples larger than 5,000 are used.
Table 1. Test Size
N $\alpha=10\%$ $\alpha=5\%$ $\alpha=1\%$
[…]
Based on $500$ replications.
6.1.2. Power
To consider the test’s performance under violations of the exogeneity assumption, we use the DGP described in Equation 6.8 with different values for $\rho$, reflecting various degrees of endogeneity. Larger values of $\rho$ produce more acute violations of the null hypothesis. Table 2 presents results for these scenarios: columns 3-4 present the MSE while the remaining columns present the empirical rejection rates. As expected, the power of the test increases with larger $\rho$ and bigger sample sizes, presenting substantial power for $\alpha=10\%$ and presenting a decreasing pattern for $\alpha=5\%$ and $\alpha=1\%$.
Table 2. Test power for violations in exogeneity
$N$ $\rho$ MSE$(\beta_{0})$ MSE$(\beta_{1})$ $\alpha=10\%$ $\alpha=5\%$ $\alpha=1\%$
5000 0.10 0.0005 0.0005 8.60% 6.00% 3.20%
0.50 0.0005 0.0104 16.03% 10.82% 5.21%
0.75 0.0005 0.0227 36.00% 27.80% 14.20%
0.80 0.0005 0.0259 42.00% 30.60% 16.80%
0.90 0.0004 0.0326 62.40% 48.40% 27.60%
0.95 0.0004 0.0360 71.34% 59.32% 35.87%
8000 0.10 0.0003 0.0005 8.40% 5.00% 2.00%
0.50 0.0003 0.0102 17.43% 12.22% 5.61%
0.75 0.0003 0.0225 45.58% 36.14% 20.08%
0.80 0.0003 0.0256 53.80% 39.80% 18.00%
0.90 0.0003 0.0323 79.00% 65.40% 36.40%
0.95 0.0002 0.0363 85.57% 75.15% 49.70%
10000 0.10 0.0003 0.0005 10.04% 7.43% 2.41%
0.50 0.0002 0.0101 15.80% 10.00% 3.80%
0.75 0.0002 0.0226 44.80% 32.80% 16.00%
0.80 0.0002 0.0256 56.60% 41.00% 21.60%
0.90 0.0002 0.0326 79.60% 66.60% 36.60%
0.95 0.0002 0.0361 90.56% 79.52% 51.61%
Based on $500$ replications.
Turning our attention to the test’s power under violations of normality of the latent error, we consider the modified data generating process in Equation 6.15. In this DGP, $F(\cdot)$ denotes an alternative density for the latent error $U$, which we vary across scenarios. Table 3 present the empirical coverage of the tests when $F(\cdot)$ is uniform, lognormal and t-student distributions.
$\left\{\begin{array}[]{lcl}Y&=&\max\{Y^{*},0\}\\ Y^{*}&=&D+U\\ D&=&2Z-V\\ \begin{pmatrix}V\\ Z\end{pmatrix}&\sim&\mathcal{N}(\mathbf{0}_{2},\mathbf{\Sigma})\\ \mathbf{\Sigma}&=&\begin{pmatrix}1&0\\ 0&1\\ \end{pmatrix}\\ U\sim F(\cdot)\end{array}\right.$ (6.15)
As expected, the rejection rate for the case of the t-student with 80 degrees of freedom is close to the nominal size of the test, since it represents a very mild violation of the normality assumption. In the other three cases, there is a strong violation of the normality assumption and the test rejects the null hypothesis in most instances. This example provides indication that the proposed test can be useful in detecting violations of the distributional assumptions in the classic Tobit model.
Table 3. Test power for violations in error structure
$N$ $U$ MSE$(\beta_{0})$ MSE$(\beta_{1})$ $\alpha=10\%$ $\alpha=5\%$ $\alpha=1\%$
5000 t-student (df=80) 0.9987 1.0018 7.21% 4.01% 1.40%
t-student (df=5) 1.027 1.0624 92.20% 81.80% 48.80%
LogNormal 1.3336 0.0896 100.00% 100.00% 100.00%
Uniform 0.892 0.785 100.00% 100.00% 100.00%
8000 t-student (df=80) 0.9999 1.0034 7.00% 4.80% 1.20%
t-student (df=5) 1.0251 1.0653 99.20% 97.80% 87.58%
LogNormal 1.3385 0.0866 100.00% 100.00% 100.00%
Uniform 0.8932 0.7896 100.00% 100.00% 100.00%
10000 t-student (df=80) 1.0016 1.0059 5.42% 2.41% 1.00%
t-student (df=5) 1.0241 1.0642 99.20% 98.40% 90.58%
LogNormal 1.3348 0.0868 100.00% 100.00% 100.00%
Uniform 0.8908 0.7862 100.00% 100.00% 100.00%
Based on $500$ replications.
7. Relaxation of the assumptions
When the test proposed in Section 5 rejects the null hypothesis of the Tobit model’s validity, researchers must pursue alternative models and less restrictive assumptions to learn confidently about the parameter of interest.
7.1. Existing Solutions
There is a vast literature on alternatives to the Tobit Model that can be implemented in the presence of censored dependent variables. Most approaches consider changes or relaxations of one of the two main assumptions associated with the Tobit model. The first assumption is the parametric distribution of the error terms and latent index form connecting treatment (and covariates) to the outcome. The second assumption is the exogeneity of the treatment or potential instrument. Here, we provide a non-exhausting survey of existing work.
cragg1971some maintains the normality of the errors, linearity of the index and exogeneity but relaxes the way censoring occurs in comparison to the latent structure of the censored outcome. Specifically, while the latent outcome is still modelled by $Y^{*}=\alpha_{0}+\alpha_{1}D+U$, they allow for the censoring to depend on a different linear index, $P(Y^{*}>0)=P(\gamma_{0}+\gamma_{1}D+e)$, increasing the model’s flexibility.
POWELL1984303 relaxes the parametric structure of the errors while maintaining the latent linear index and treatment exogeneity, and estimates the parameters of interest by least absolute deviations. powell1986symmetrically also maintains linearity of the latent index and treatment exogeneity, but relaxes normality by imposing symmetrical distributions to the latent errors, which leads to estimation by symmetrically censored least squares. newey1987efficient relaxes exogeneity of the treatment, relying on normality and an instrumental variable to identify the model, which is estimated by generalized least squares.
honore1994pairwise relax linearity and mean independence of the unobservable with respect to the treatment to exploit the idea that, although $Y^{*}_{i}-\alpha_{0}-\alpha_{1}D_{i}$ is not mean-independent of $D_{i}$, one can trim any pair of residuals $Y^{*}_{i}-\alpha_{0}-\alpha_{1}D_{i}$ and $Y^{*}_{j}-\alpha_{0}-\alpha_{1}D_{j}$, and the trimmed residuals are independent and identically distributed conditional on $D_{i},D_{j}$. They estimate the model by identically censored least absolute deviations and identically censored least squares (ICLS). das2002estimators estimates a model using symmetrically censored least squares that relaxes exogeneity of the treatment and normality of the errors. To achieve that they rely on instrumental variables, linearity of the mean of the structural error conditional on the reduced form error, and mean independence of the reduced form error.
blundell2007censored proposes a control variable approach that relaxes exogeneity and normality but maintains the latent linear structure ($\alpha_{0}+\alpha_{1}D+U$). Crucially, they impose that the distribution (or quantiles) of the latent error conditional on the treatment and instrument is only a function of the control variable $V=D-\pi(Z)$, which isolates the endogenous variation on the treatment. [^9] This assumption is weaker than independence of all errors and instruments since it does not impose $V$ independent of $Z$ but is neither stronger nor weaker than independence of $U$ and $Z$, since it permits $Z$ to affect $U$ through $V$. This allows them to estimate the effect of the treatment by censored quantile instrumental variable regression augmented by a control variable given by the quantiles of $U$ conditional on $V$ at the quantile of interest. In a similar spirit, chernozhukov2015quantile focuses on conditional quantile functions and flexible approaches to estimate the control variable in the first-stage.
Finally, chesher2023iv provide partial identification results for a general alternative by relaxing exogeneity of the treatment and instrument, linearity of the latent index and imposing no parametric structure of the error term. They characterize the identified set for the parameters of interest following the Generalized Instrumental Variables framework (chesher2017generalized) , relying on the assumption that the relationship of $Y^{*}$ to treatment and errors is continuous and monotonic in the errors. Their approach uses the residual sets associated with the structure of the latent function and conditional probability of the error term given potential instruments.
7.2. Partial identification under monotonicity
In this subsection, we present an approach that partially identifies the effect of an endogenous treatment variable by replacing the normality and exogeneity assumptions with a monotonicity in treatment selection constraint. While less general than chesher2023iv, this approach is easy to implement and could be useful to empirical researchers.
Consider the model that maintains linearity (or a known structure of $Y^{*}$ up to a finite number of parameters),
$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\alpha_{0}+\alpha_{1}D+U\end{array}\right.$ (7.3)
As an alternative to treatment exogeneity and normality, consider a constraint on the direction of the endogenous relationship between the treatment and the unobservables that affect the outcome.
Assumption 5 (Monotone Treatment Selection - MTS).
Let $E(U|D=d,Y>0)\equiv\Gamma(d)$. Then, for any $d>d^{*}$ we either have $\Gamma(d)<\Gamma(d^{*})$ or $\Gamma(d)>\Gamma(d^{*})$.
Assumption 5 is common in the partial identification literature (jiang2014monotone; manski1997monotone; manski1998monotone; okumura2014concave). In this context, we restrict the latent selection to behave monotonically with the treatment. This assumption is embedded in the classic Tobit model since the inverse mills ratio, $\lambda(\cdot)$, is monotonic (and decreasing) in the treatment variable. Furthermore, independence between $D$ and $U$ restricts the sign of the coefficient of the selection term directly, as the derivative of the inverse mills-ratio is $\lambda^{\prime}(\alpha_{0}+\alpha_{1}D)\alpha_{1}$. Thus, without imposing independence or a parametric latent structure, we maintain a relevant property of the Tobit model that aids identification. Since it is not as restrictive as imposing a parametric structure and independence, we can only partially identify the parameter of interest.
Under Assumption 5 and the model described in 7.3, treatment and outcome are not independent. Note that,
$E(Y|D=d,Y>0)$ $=$ $\alpha_{0}+\alpha_{1}d+\Gamma(d)$ (7.4)
$E(Y|D=d,Y>0)$displaystyle-- $\alpha_{0}-\alpha_{1}d=\Gamma(d)$ (7.5)
Then, for any two $d>d^{*}$ we have, by Assumption 5:
$\begin{array}[]{lcl}\Gamma(d)&<&\Gamma(d^{*})\\ &\Leftrightarrow&\\ E(Y|D=d,Y>0)-\alpha_{0}-\alpha_{1}d&<&E(Y|D=d^{*},Y>0)-\alpha_{0}-\alpha_{1}d^{*}\\ &\Leftrightarrow&\\ \alpha_{1}&>&\frac{E(Y|D=d^{*},Y>0)-E(Y|D=d,Y>0)}{d^{*}-d},\end{array}$ (7.11)
which implies a lower bound on parameter interest.
For a binary treatment $D\in\{0,1\}$ the lower bound is, intuitively, the difference in average outcomes between treated and untreated individuals away from the mass point at zero:
$\begin{array}[]{lcl}\alpha_{1}&>&E(Y|D=1,Y>0)-E(Y|D=0,Y>0).\end{array}$ (7.13)
The bound can be more informative in the case of a continuous or multi-valued treatment. If $\Gamma(D)$ is differentiable we have:
$\alpha_{1}$ $>$ $\frac{\partial E(Y|D=d,Y>0)}{\partial d}\text{, for all d.}$. (7.14)
Since the inequality holds for any $d$ in the continuous case or for any $d,d^{*}$ for multi-valued discrete treatment, the linear index structure with constant parameters implies that tighter bounds for $\alpha_{1}$ are given by the maximum value of $\frac{\partial E(Y|D=d,Y>0)}{\partial d}$ across all possible points in the support for $D$. Analogous results with the inequalities reverted can be derived for any $d>d^{*}$, as we have $\Gamma(d)>\Gamma(d^{*})$.
One-sided simple confidence regions can be computed based on these outer sets of the treatment effect. One can estimate $E(Y|D=d,Y>0)$ using its sample analogs, and rely on their asymptotic normality. For example, let the estimators be given by $\widehat{E}(Y|D=d,Y>0)$. By the continuous mapping theorem, $\frac{\widehat{E}(Y|D=d^{*},Y>0)-\widehat{E}(Y|D=d,Y>0)}{d^{*}-d}$ is asymptotically normal. Thus, a one-sided confidence interval for $\frac{E(Y|D=d^{*},Y>0)-E(Y|D=d,Y>0)}{d^{*}-d}$ can be computed via bootstrap which implies a conservative estimate for the lower bound for $\alpha_{1}$.
A bootstrap procedure could be used for a nonparametric estimator of $\frac{\partial E(Y|D=d,Y>0)}{\partial d}$ using a local polynomial regression. In this case the estimator is asymptotically normal and converges at a nonparametric rate that depends on the bandwidth $h$. Recent developments in calonico2018effect; calonico2022coverage for estimation and optimal coverage error bandwidth and kernel selection methods that are nonparametric robust bias-corrected (RBC) can be used through their convenient implementation using the package nprobust (calonico2019nprobust).
8. Empirical illustration: Olson1998
In this section, we implement the test of model validity described in Section 5 and the alternative method for partial identification under monotonicity of treatment selection using data from Olson1998.
Employer-provided health insurance is one of the most important mechanisms by which individuals under the age of 65 gain access to health care in the USA. Health insurance coverage insures individuals and their families against the financial risk associated with the costs of treating health conditions. Olson1998 hypothesize that women who lack health insurance coverage through their partner’s employer are more likely to work full-time to obtain health insurance for themselves and other family members. Using data from the Current Population Survey (CPS) from March 1993, Olson1998 finds that women without spousal coverage were less likely to be out of the labor force than those with partner’s coverage.
The 1993 CPS refers to hours worked and health insurance benefits provided by the longest job held during 1992. Olson1998 created the study’s sample by matching spouse’s records on household ID codes to create a single record for each married couple in the survey. Next, they excluded couples for which the wife was over the age of 64 or her labor force status at the time of the interview in March 1993 was different from her status in 1992. Finally, households were eliminated if respondents were assigned imputed values for the relevant health insurance questions or the husband’s total earnings in 1992. This produced a sample of 22,272 families. For 15,620 families, the wife’s usual weekly hours worked in 1992 were greater than zero.
Olson1998 is interested in the effect of having health insurance coverage through the partner’s employment ($D$) on women’s labor supply measured as the usual hours worked per week ($Y$). A Tobit model was estimated using this labor supply measure, and a dummy indicating insurance coverage through the husband’s job. Table 4 presents the results for model 2.4 under assumptions 1-2.
Table 4. Tobit specification for Olson1998
MLE
Hours Worked
Wife covered through partner’s job $-0.6876071^{*}$
$(0.3637667)$
Constant $21.70212^{***}$
$(0.2866817)$
$N$ $22,272$
Standard errors (in parentheses); ***: significant at 1% level, * significant at 10% level.
The first column presents estimates indicating that having health insurance through the husband’s employment has a negative effect on women’s hours worked. However, the model is rejected at conventional significance levels when we test for the Tobit model’s validity as proposed in Section 5 ($\hat{\theta}_{0.99}=0.4491>0$, $\hat{\theta}_{0.95}=0.4517>0$, and $\hat{\theta}_{0.90}=0.4529>0$).
Confidence Bounds under Assumption 5 and latent linear index
Given the rejection of the Tobit model in this case, we relax the distributional and exogeneity of treatment assumptions, and construct lower bounds on the treatment effect for the effect of receiving health insurance through partner’s employment on women’s labor supply under the MTS Assumption 5 and latent linear index only. In the context of this application, Assumption 5 imposes that average unobservables affecting women’s preferences related to hours worked away from home, such as schedule flexibility and reservation wages, are monotonically increasing in characteristics that drive husbands to take jobs that cover spouses. For example, households that prioritize flexible schedules for women might similarly prioritize partner jobs that provide more generous non-wage benefits such as family health insurance coverage.
Table 5. Confidence sets for parameter of interest
Parameter Tobit estimates 95% conf. sets for the LB for $\alpha_{1}$
$\alpha_{1}$0.6876071 [-0.8146481, -0.5687435]
conf.: confidence; LB: lower bound; UB: upper bound.
The first column of Table 5 repeats the estimate for $\alpha_{1}$ from Table 4. The second column reports the 95% confidence set for the lower bound for $\alpha_{1}$, obtained under Assumption 5 and latent linear index only. Recall that the lower end of the confidence set is the lower bound for the treatment effect of interest $\alpha_{1}$. Even after relaxing the normality of errors and exogeneity of the treatment, we obtain informative bounds for the size of the effect of having employer-provided health insurance through their partner’s employment on the female labor supply. These estimates indicate an effect lower bound of -0.8146 hours worked per week for married women who work positive hours - the intensive margin of labor supply. While we can rule out a reduction of one hour per week in female labor supply, we cannot rule out that the effect is zero.
9. Empirical Illustration: lee1995semi
In this section we implement the proposed test to the data from lee1995semi. Using the 1987 cross-section of the Michigan Panel Study of Income Dynamics, the authors study the responses of married women’s labor supply ($Y$) - measured in hours per year - to “other” household income ($D$), which is endogenous. The potential instrumental variables explored are dummy variables for the husband’s occupation ($Z$), which implies the restrictive identifying assumption that the wife’s labor supply is affected by the husbands’ occupation only through their income.
Following lee1995semi, we proceed with the analysis focusing on the data for married couples with non-negative family total income and where the wife was of working age (18-64) and not self-employed. The number of observations is 3382 and 26 percent of wage observations are censored. Table LABEL:Table:Lee presents the estimates obtained using IV Tobit model.
Table 6. IV Tobit specification for lee1995semi
MLE
Other household income Hours per year worked
Husband’s occupation: manager or professional $198.3894^{***}$
$(10.43154)$
Other household income $0.7029415^{***}$
(.$(.237801)$
Constant $240.4661^{***}$ $752.9173^{***}$
$(5.563514)$ $(74.01555)$
$\rho$ $-0.3069913^{***}$
(.$(.0572254)$
$n$ $3,382$ $3,382$
Standard errors (in parentheses); ***: significant at 1% level.
The first column presents the first-stage estimates indicating the relevance of the potential $IV$. The second column presents the structural equation reflecting the effect the treatment variable has on the outcome equation. The parameter $\rho$ shows evidence of a negative correlation between the unobservables driving the other household income and hours worked. The estimated coefficient of interest indicates that other household income has a positive effect on the wife’s labor supply. However, the model is rejected at conventional significance levels when we test for the IV-Tobit model’s validity, ($\hat{\theta}_{0.99}=0.0247848>0$, $\hat{\theta}_{0.95}=0.0270805>0$, and $\hat{\theta}_{0.90}=0.0283793>0$).
10. Conclusion
In this paper, we develop sharp testable equalities for Tobit-type “two-part” models that can detect all observable violations of the model’s assumptions. By converting these sharp equalities into conditional moment inequalities, we propose a testing procedure that detects violations of the Tobit model assumptions on a grid on the joint support of the outcome (and treatment) variables leveraging inference results from Chernozhukov, Lee, and Rosen (2013) and the implementation from Chernozhukov et al. (2015).
Simulation results suggest the test performs well for reasonably sized samples (larger than 5000 observations). For smaller samples, the test is conservative, overejecting the null hypothesis of model validity. Simulations indicate that the test is powerful to detect violations of the exogeneity assumption for the treatment/instrument that affect the point estimates and inference. Finally, the proposed test exhibits good performance for violations of the distributional assumptions about the error structure.
We provide a brief review of existing models that could be implemented under weak/alternative assumptions when the Tobit model is rejected. Furthermore, we propose a simple model that partially identifies the parameter of interest by relying solely on linear index and monotone treatment selection restrictions, a standard assumption from the partial identification literature (manski2000amonotone).
We illustrate our methods on data from Olson1998 and lee1995semi. We conclude that stylized versions (excluding covariates) of both empirical strategies reject the null hypothesis of a Tobit model and an IV-Tobit model respectively. Furthermore when illustrating partial identification strategy on Olson1998 we obtain informative lower bounds for the effect of having employer-provided health insurance through their partner’s employment on the female labor supply. These estimates indicate an effect lower bound of $-0.8146$ hours worked per week for married women working positive hours - the intensive margin of labor supply. While we can rule out a reduction of one hour per week in female labor supply, we cannot rule out that the effect is zero.
Appendix A Sharpness for the classic Tobit case
Let $\alpha_{1},\alpha_{0}$ be identified. Suppose 3.1- 3.3 hold. Let $f_{\tilde{U}|D}=\frac{1}{\sqrt{2\pi}}e^{-1/2\tilde{U}^{2}}$. Let $\tilde{Y}^{*}=\alpha_{0}+\alpha_{1}D+\tilde{U}$ and $\tilde{Y}=\max(0,\tilde{Y}^{*})$.
From $f_{\tilde{U}|D}$ we can see assumption 1 holds, since $f_{\tilde{U}|D}=f_{\tilde{U}}$. Also note $f_{\tilde{U}}$ is the $N(0,1)$ density, thus assumption 2 holds. Considering the positive values for $\tilde{Y}$, for any constants $c_{0},c_{1}$ such that $0<c_{0}<c_{1}$:
$P$ $(c_{0}\leq\tilde{Y}\leq c_{1}|D=d)=P(c_{0}\leq\tilde{Y}^{*}\leq c_{1}|D=d)=P(c_{0}\leq\alpha_{0}+\alpha_{1}d+\tilde{U}\leq c_{1}|D=d)$
$=$ $P(c_{0}-\alpha_{0}-\alpha_{1}d\leq\tilde{U}\leq c_{1}-\alpha_{0}-\alpha_{1}d|D=d)=P(c_{0}-\alpha_{0}-\alpha_{1}d\leq\tilde{U}\leq c_{1}-\alpha_{0}-\alpha_{1}d)$
$=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)=P(c_{0}\leq Y\leq c_{1}|D=d).$
The first three equalities follow from the definitions of $\tilde{Y}^{*}$ and $\tilde{Y}$. The fourth and fifth steps use the specific choice for the probability density of $\tilde{U}$, which implies independence from D and normality respectively. The last step uses the relationship between the observed data and the model, established in 3.1. Similar reasoning applies for $P(c_{1}\leq\tilde{Y}|D=d)=P(c_{1}\leq\alpha_{0}+\alpha_{1}d+\tilde{U}|D=d)=1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)=P(c_{1}\leq Y|D=d)$.
Furthermore, for $P(\tilde{Y}=0|D=d)$:
$P$ $(\tilde{Y}=0|D=d)=P(\tilde{Y}^{*}\leq 0|D=d)=P(\alpha_{0}+\alpha_{1}d+\tilde{U}\leq 0|D=d)$
$=$ $P(\tilde{U}\leq-\alpha_{0}-\alpha_{1}d|D=d)=P(\tilde{U}\leq-\alpha_{0}-\alpha_{1}d)$
$=$ $\Phi(-\alpha_{0}-\alpha_{1}d)=1-\Phi(\alpha_{0}+\alpha_{1}d)=P(Y=0|D=d)$
The last step uses the equality established in 3.3. Thus, we characterized the distribution of $\tilde{Y}|D$, which is equal to the one of $Y|D$, with $D$ given. Thus, we can induce the observed distribution $Y,D$.
Appendix B Details of the IV-Tobit case
B.1. Derivation of equation 3.4
$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z)$ (B.1)
$=$ $P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1},Y^{*}\geq 0|Z=z)+P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1},Y^{*}<0|Z=z)$
$=$ $P(c_{0}\leq Y^{*}\leq c_{1},d_{0}\leq D\leq d_{1},Y^{*}\geq 0|Z=z)$
$=$ $P(c_{0}\leq\beta_{0}+\beta_{1}z+W\leq c_{1},d_{0}\leq\gamma_{0}+\gamma_{1}z+V\leq d_{1},Y^{*}\geq 0|Z=z)$
$=$ $P(c_{0}-\beta_{0}-\beta_{1}z\leq W\leq c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq V\leq d_{1}-\gamma_{0}-\gamma_{1}z,Y^{*}\geq 0|Z=z)$
$=$ $P(c_{0}-\beta_{0}-\beta_{1}z\leq W\leq c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq V\leq d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
$=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)-\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)+\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
The first equality follows from the law of total probability. The second through fourth equalities are in consequence of the model structure in 2.16. The fifth step uses Assumption 3 and $c_{0}\geq 0$. The final equality is by the properties of probabilities and the joint normality for $W,V$ (Assumption 4).
B.2. Derivation of equation 3.10
By a similar approach to the derivation of 3.4:
$P(Y=0,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $P(Y^{*}\leq 0,d_{0}\leq D\leq d_{1}|Z=z)$ (B.2)
$=$ $P(W\leq-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}\leq V\leq d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
$=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
-- $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
B.3. Proof of Sharpness
Let $\beta_{0},\beta_{1},\gamma_{0},\gamma_{1},\rho$ be identified and equalities 3.4-3.12 hold. Define the joint density of $\left(\tilde{W},\tilde{V}\right)$ conditional on $Z$ as
$f_{\left(\tilde{W},\tilde{V}|Z\right)}(w,v|z)=\frac{1}{\sqrt{1-\rho^{2}}}\phi\left(\frac{w-\rho v}{\sqrt{1-\rho^{2}}}\right)\phi(v),$
where $\phi(t)=\exp(-t^{2}/2)$, and define
$\left\{\begin{array}[]{lcl}\tilde{Y}&=&\max(0,\tilde{Y}^{*})\\ \\ \tilde{Y}^{*}&=&\beta_{0}+\beta_{1}Z+\tilde{W}\\ \\ \tilde{D}&=&\gamma_{0}+\gamma_{1}Z+\tilde{V}\end{array}\right.$
Note that $f_{\left(\tilde{W},\tilde{V}|Z\right)}=f_{\left(\tilde{W},\tilde{V}\right)}$, thus assumption 3 holds. Similarly, $\left(\tilde{W},\tilde{V}\right)$ follow a bivariate normal distribution as $\begin{pmatrix}\tilde{W}\\ \tilde{V}\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\end{pmatrix}$, and $\Sigma=\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}$. Let $\tilde{U}=\tilde{W}-\alpha_{1}\tilde{V}$, which implies $\tilde{U},\tilde{V}$ satisfies assumption 4, given the scale-location normalizations.
Then, for any constants $0<c_{0}\leq c_{1}$ and $d_{0}<d_{1}$:
$P$ $(c_{0}\leq\tilde{Y}\leq c_{1},d_{0}\leq\tilde{D}\leq d_{1}|Z=z)=P(c_{0}\leq\tilde{Y}^{*}\leq c_{1},d_{0}\leq\tilde{D}\leq d_{1}|Z=z)$
$=$ $P(c_{0}\leq\beta_{0}+\beta_{1}z+\tilde{W}\leq c_{1},d_{0}\leq\gamma_{0}+\gamma_{1}z+\tilde{V}\leq d_{1}|Z=z)$
$=$ $P(c_{0}-\beta_{0}-\beta_{1}z\leq\tilde{W}\leq c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq\tilde{V}\leq d_{1}-\gamma_{0}-\gamma_{1}z|Z=z)$
$=$ $P(c_{0}-\beta_{0}-\beta_{1}z\leq\tilde{W}\leq c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq\tilde{V}\leq d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
$=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)-\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)+\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
$=$ $P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z)$
The first through third equalities follow from the definitions of $\tilde{Y},\tilde{Y}^{*}$ and $\tilde{D}$. The fourth and fifth steps are consequences of the particular choice for joint density for $\left(\tilde{W},\tilde{V}\right)$ conditional on $Z$. The final equality is given by the relationship between observable data and the latent model structure in Equation 3.4. Similar derivations hold for 3.5-3.9.
For the accumulation point, at $Y=0$:
$P$ $(\tilde{Y}=0,d_{0}\leq\tilde{D}\leq d_{1}|Z=z)=P(\tilde{Y}^{*}\leq 0,d_{0}\leq\tilde{D}\leq d_{1}|Z=z)$
$=$ $P(\beta_{0}+\beta_{1}z+\tilde{W}\leq 0,d_{0}\leq\gamma_{0}+\gamma_{1}z+\tilde{V}\leq d_{1}|Z=z)$
$=$ $P(\tilde{W}\leq-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq\tilde{V}\leq d_{1}-\gamma_{0}-\gamma_{1}z|Z=z)$
$=$ $P(\tilde{W}\leq-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq\tilde{V}\leq d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
$=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)-\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
$=$ $P(Y=0,d_{0}\leq D\leq d_{1}|Z=z).$
The steps of the proof are similar to previous cases and the last equality is given by the relationship between observable data and the latent model structure in Equation 3.10. Similar derivations hold for equalities 3.11-3.12.
Thus, we characterized the distribution of $(\tilde{Y},\tilde{D})|Z$, which coincides with the joint distribution of $(Y,D)|Z$, for given $Z$. Thus, we induced the observed distribution of the data $Y,D,Z$.
Appendix C Type 2 Tobit testable implications
In this section we derive the results of the Type 2 tobit model also known as selection models or Heckman selection type of models (heckman1979sample). The basic setup with no covariates (which can be extended in several directions and with different distributional assumptions as well as to incorporate treatment endogeneity, as we pointed out for the classic Tobit model in remark 3) is:
$\left\{\begin{array}[]{lcl}Y&=&Y^{*}\quad\text{if}\quad S=1\\ \\ Y&=&\text{missing}\quad\text{if}\quad S=0\\ \\ Y^{*}&=&\alpha_{0}+\alpha_{1}D+U\\ \\ S&=&1\{\gamma_{0}+\gamma_{1}Z+V\geq 0\}\end{array}\right.$ (C.8)
Where $U,V$ are the latent structural error terms.
Assumption 6.
$D,Z$ be independent of $U,V$. Furthermore, let $\gamma_{1}\neq 0$
Assumption 7.
Let $U,V$ follow a bivariate normal distribution with covariance $\rho$, i.e., $\begin{pmatrix}U\\ V\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma=\begin{pmatrix}\sigma^{2}_{U}&\rho_{UV}\\ \rho_{UV}&\sigma^{2}_{V}\end{pmatrix}$.
Note that $Y$ is missing at $S=0$. Thus, fully characterizing the distribution implies observing the behaviour at the missing point and beyond it.
From the observed data the conditional probabilities of $c_{0}\leq Y\leq c_{1}$ for an $c_{1},c_{0}$ can be computed. Then note for any $d\in\mathcal{D}$ and $z\in\mathcal{Z}$ :
$P(Y=\text{missing}\quad|D=d,Z=z)$ $=$ $P(S=0|D=d,Z=z)$ (C.9)
$=$ $P(\gamma_{0}+\gamma_{1}Z+V<0|D=d,Z=z)$
$=$ $P(\gamma_{0}+\gamma_{1}z+V<0)$
$=$ $P(V<-\gamma_{0}-\gamma_{1}z)$
$=$ $\Phi_{v}(\frac{-\gamma_{0}-\gamma_{1}z}{\sigma_{V}})$
Where the first and the second equality is due to the structure of the model described in equation C.8, the third one is due to Assumption 6 and the last one is due to the normalization for normal random variables and Assumption 7.
Additionally,
$P(c_{0}\leq Y\leq c_{1}|D=d,Z=z)$ $=$ $P(c_{0}\leq Y\leq c_{1},S=0|D=d,Z=z)$ (C.10)
$+$ $P(c_{0}\leq Y\leq c_{1},S=1|D=d,Z=z)$
$=$ $P(c_{0}\leq Y^{*}\leq c_{1},S=1|D=d,Z=z)$
$=$ $P(c_{0}\leq\alpha_{0}+\alpha_{1}D+U\leq c_{1},\gamma_{0}+\gamma_{1}Z+V\geq 0|D=d,Z=z)$
$=$ $P(c_{0}\leq\alpha_{0}+\alpha_{1}d+U\leq c_{1},\gamma_{0}+\gamma_{1}z+V\geq 0|D=d,Z=z)$
$=$ $P(c_{0}-\alpha_{0}-\alpha_{1}d\leq U\leq c_{1}-\alpha_{0}-\alpha_{1}d,V\geq-\gamma_{0}-\gamma_{1}z|D=d,Z=z)$
$=$ $P(c_{0}-\alpha_{0}-\alpha_{1}d\leq U\leq c_{1}-\alpha_{0}-\alpha_{1}d,V\geq-\gamma_{0}-\gamma_{1}z)$
$=$ $\Phi_{U}(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi_{U}(c_{0}-\alpha_{0}-\alpha_{1}d)$
-- $\Phi_{U,V}(c_{1}-\alpha_{0}-\alpha_{1}d,-\gamma_{0}-\gamma_{1}z)$
$+$ $\Phi_{U,V}(c_{0}-\alpha_{0}-\alpha_{1}d,-\gamma_{0}-\gamma_{1}z)$
The first equality follows from the law of total probability. The second through fifth equalities follow from the model’s structure described in Equation C.8, namely $P(Y>0,Y^{*}<0|D=d)=0$ and $Y^{*}=Y,\text{not missing}$ for $S=1$ and the structure of $S$ as well as the latent linear model. The sixth one is due to Assumption 6 The seventh one is due to Assumption 7 as well as properties of normal random variables as well as bivariate cummulative distribution functions. Thus, we can test in a similar fashion as we did in the classic Tobit case and the IV-Tobit case but using the previous two type of equalities in addition to the ones of the form $P(Y\geq c_{2}|D=d,Z=z)$ which can be easily derived.
Appendix D Results from Remark 3
In this appendix we derive the testable implications of the extensions discussed in Remark 3. We will discuss testable implications assuming identification of the relevant parameters or distributions.
D.1. barros2018generalized and carson2007Tobit
barros2018generalized proposes a variant of the tobit model with Elliptically contoured distributions and a non-zero threshold. At the same time carson2007Tobit proposes a tobit model with a non-zero threshold. In this section we combine both type of results and report the testable implications with a generic non-zero threshold and a generic known or identifiable parametric distribution function. In addition, extend the latent structure to be also known or an identifiable function up to a vector of parameters but invertible in $U$. In this context, let:
$\left\{\begin{array}[]{lcl}Y&=&\max(\tau,Y^{*})\\ \\ Y^{*}&=&g(\alpha,D,U)\end{array}\right.$ (D.4)
Assumption 8.
$D$ is independent of $U$.
Assumption 9.
$U$ is distributed according to distribution $F_{H}(.)$ with parameters $\omega$.
In this context, the testable implications are, starting at the continuous part of the distribution of $Y$, the conditional probabilities of $c_{0}\leq Y\leq c_{1}$ for a $c_{1},c_{0}>\tau$ are observed. For any value of the treatment variables $d\in\mathcal{D}$:
$P(c_{0}\leq Y\leq c_{1}|D=d)$ $=$ $P(c_{0}\leq Y\leq c_{1},Y^{*}\geq\tau|D=d)+P(c_{0}\leq Y\leq c_{1},Y^{*}<\tau|D=d)$ (D.5)
$=$ $P(c_{0}\leq Y\leq c_{1},Y^{*}\geq\tau|D=d)$
$=$ $P(c_{0}\leq Y^{*}\leq c_{1},Y^{*}\geq\tau|D=d)$
$=$ $P(c_{0}\leq Y^{*}\leq c_{1}|D=d)$
$=$ $P(c_{0}\leq g(\alpha,D,U)\leq c_{1}|D=d)$
$=$ $P(g^{-1}(\alpha,d,c_{0})\leq U\leq g^{-1}(\alpha,d,c_{1})|D=d)$
$=$ $F_{H}(g^{-1}(\alpha,d,c_{1});\omega)-F_{H}(g^{-1}(\alpha,d,c_{0});\omega)$
Turning to the accumulation point, the observed event of $Y=0$,
$P(Y=0|D=d)=P(Y^{*}\leq\tau|D=d)=F_{H}(g^{-1}(\alpha,d,\tau);\omega).$ (D.6)
Which can then be used to test in a similar way as did for the classic tobit or iv-tobit by adding the type of equalities $P(Y\geq c_{2}|D=d)$ which can be derived in a similar fashion. Also, note it is easy to extend the previous development to cases where the truncation takes the form $Y=Y^{*}$ if $\tau_{0}\leq Y^{*}\leq\tau_{1}$, $Y=\tau_{0}$ if $\tau_{0}\geq Y^{*}$ and $Y=\tau_{1}$ if $Y^{*}\geq\tau_{1}$.
D.2. wooldridge2005simple; honore2000estimation; honore1993orthogonality
Consider the following dynamic version of the tobit model which is related to wooldridge2005simple; honore2000estimation; honore1993orthogonality. Here we specify the conditional behaviour of $c_{i}$ in the spirit of wooldridge2005simple and others.
$\left\{\begin{array}[]{lcl}Y_{i,t}&=&\max(0,Y_{i,t}^{*})\\ \\ Y_{i,t}^{*}&=&\alpha_{0}+\alpha_{1}D_{i,t}+\alpha_{2}g(Y_{i,t-1})+c_{i}+U_{i,t}\end{array}\right.$ (D.10)
Assumption 10.
$D_{i,t}$ is independent of $U_{i,t}$ given $Y_{i,t-1},c_{i}$.
Assumption 11.
$U_{i,t}|Y_{i,t-1},c_{i}$ is distributed $N(0,1)$.
Assumption 12.
$c_{i}|Y_{i,t-1},D_{i,t-1}$ has a known distribution such as $N(0,1)$.
Then, starting at the continuous part of the distribution of $Y$, the conditional probabilities of $c_{0}\leq Y\leq c_{1}$ for a $c_{1},c_{0}>0$ are observed. For any value of the treatment variables $d\in\mathcal{D}$, $P(c_{0}\leq Y_{i,t}\leq c_{1}|D_{i,t}=d,Y_{i,t-1}=y)$:
$=$ $\int P(c_{0}\leq Y_{i,t}\leq c_{1}|D_{i,t}=d,Y_{i,t-1}=y,c_{i})f(c_{i}|D_{i,t}=d,Y_{i,t-1}=y)dc_{i}$ (D.11)
$=$ $\int P(c_{0}\leq Y_{i,t}\leq c_{1}|Y_{i,t-1}=y,c_{i})f(c_{i}|D_{i,t}=d,Y_{i,t-1}=y)dc_{i}$
$=$ $\int P(c_{0}\leq Y^{*}_{i,t}\leq c_{1}|Y_{i,t-1}=y,c_{i})f(c_{i}|D_{i,t}=d,Y_{i,t-1}=y)dc_{i}$
$=$ $\int P(c_{0}-\alpha_{0}-\alpha_{1}d-\alpha_{2}g(y)-c_{i}\leq U_{i,t}\leq c_{1}-\alpha_{0}-\alpha_{1}d-\alpha_{2}g(y)-c_{i}|Y_{i,t-1}=y,c_{i})f(c_{i}|d,y)dc_{i}$
[…]
Similarly,
$P(Y_{i,t}=0|D_{i,t}=d,Y_{i,t-1}=y)$ $=$ $\int P(Y_{i,t}=0|D_{i,t}=d,Y_{i,t-1}=y,c_{i})f(c_{i}|D_{i,t}=d,Y_{i,t-1}=y)dc_{i}$ (D.12)
$=$ $\int P(Y_{i,t}=0|Y_{i,t-1}=y,c_{i})f(c_{i}|D_{i,t}=d,Y_{i,t-1}=y)dc_{i}$
$=$ $\int P(Y^{*}_{i,t}\leq 0|Y_{i,t-1}=y,c_{i})f(c_{i}|D_{i,t}=d,Y_{i,t-1}=y)dc_{i}$
$=$ $\int P(U_{i,t}\leq-\alpha_{0}-\alpha_{1}d-\alpha_{2}g(y)-c_{i}|Y_{i,t-1}=y,c_{i})f(c_{i}|d,y)dc_{i}$
[…]
Which can then be used to test in a similar way as did for the classic tobit or iv-tobit by adding the type of equalities $P(Y_{i,t}\geq c_{2}|D_{i,t}=d,Y_{i,t-1}=y)$, with the caveat that the right hand side will need to be numerically integrated or approximated by an estimator when no close form is available. Such estimator should assure that the whole left hand side converges at root-$N$ in order for the estimation of the the null model to be asymptotically negligble(Acerenza, Bartalotti, and Kédagni, 2023, Appendix B).
Appendix E Non-learnability
E.1. Non-Learnability for the classic Tobit model
As mentioned in Remark 2, the standard Tobit model is non-verifiable, that is, we cannot learn if the maintained Tobit model is the true data generating process based on the sharp equalities proposed in Section 3.1. One can show that by finding an alternative model that is compatible with the equalities (3.1)-(3.3) in all cases in which the Tobit model could not be disregarded. Suppose that,
$\left\{\begin{array}[]{lcl}\bar{Y}&=&\max(0,\bar{Y}^{*})\\ \\ \bar{Y}^{*}&=&\alpha_{0}+\alpha_{1}D+\bar{U}\end{array}\right.$ (E.4)
For $c_{0}\leq Y\leq c_{1}$ for a $c_{1},c_{0}>0$ and for any value of the treatment variables $d\in\mathcal{D}$, assume that the equalities (3.1)-(3.3) hold:
$P(c_{0}\leq Y\leq c_{1}|D=d)$ $=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)$
$P(c_{1}\leq Y|D=d)$ $=$ $1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)$
$P(Y=0|D=d)$ $=$ $1-\Phi(\alpha_{0}+\alpha_{1}d).$
Now define $\bar{Y}_{d}=\alpha_{0}+\alpha_{1}d+\bar{U}$ and define the following distribution function:
$P(\bar{Y}_{d}\leq c_{1},\bar{U}\leq 0|D=d)=\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)\Lambda(d)$
$P(\bar{Y}_{d}\leq c_{1},\bar{U}>0|D=d)=\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)(1-\Lambda(d))$
Where $\Lambda(d)=\frac{e^{d}}{1+e^{d}}$. In the above proposed DGP, $\bar{U}$ is not normally distributed. However, the DGP is compatible with the data since,
$P(\bar{Y}=0|D=d)$ $=$ $P(\bar{Y}^{*}\leq 0|D=d)=P(\alpha_{0}+\alpha_{1}d+\bar{U}\leq 0|D=d)$ (E.5)
$=$ $P(\alpha_{0}+\alpha_{1}d+\bar{U}\leq 0,\bar{U}\leq 0|D=d)$
$+$ $P(\alpha_{0}+\alpha_{1}d+\bar{U}\leq 0,\bar{U}>0|D=d)$
$=$ $P(Y_{d}\leq 0,\bar{U}\leq 0|D=d)+P(Y_{d}\leq 0,\bar{U}>0|D=d)$
$=$ $\Phi(-\alpha_{0}-\alpha_{1}d)\Lambda(d)+\Phi(-\alpha_{0}-\alpha_{1}d)(1-\Lambda(d))$
$=$ $\Phi(-\alpha_{0}-\alpha_{1}d)=1-\Phi(\alpha_{0}+\alpha_{1}d)=P(Y=0|D=d)$
where the last equality is due to the testable equalities holding. Similarly, for any $c_{1}>0$,
$P(c_{1}<\bar{Y}|D=d)$ $=$ $P(c_{1}<\bar{Y}^{*}|D=d)=P(c_{1}<\alpha_{0}+\alpha_{1}d+\bar{U}|D=d)$ (E.6)
$=$ $P(c_{1}<\bar{Y}_{d}|D=d)=1-P(\bar{Y}_{d}<c_{1}|D=d)$
$=$ $1-P(\bar{Y}_{d}<c_{1},\bar{U}\leq 0|D=d)-P(\bar{Y}_{d}<c_{1}\leq 0,\bar{U}>0|D=d)$
$=$ $1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)\Lambda(d)+\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)(1-\Lambda(d))$
$=$ $1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)=P(c_{1}<Y|D=d).$
Finally,
$P(c_{0}\leq\bar{Y}\leq c_{1}|D=d)$ $=$ $P(c_{0}\leq\bar{Y}^{*}\leq c_{1}|D=d)=P(c_{0}<\alpha_{0}+\alpha_{1}d+\bar{U}<c_{1}|D=d)$ (E.7)
$=$ $P(c_{0}<\bar{Y}_{d}<c_{1}|D=d)=P(\bar{Y}_{d}<c_{1}|D=d)P-(\bar{Y}_{d}<c_{0}|D=d)$
$=$ $P(\bar{Y}_{d}<c_{1},\bar{U}\leq 0|D=d)+P(\bar{Y}_{d}<c_{1}\leq 0,\bar{U}>0|D=d)$
-- $P(\bar{Y}_{d}<c_{0},\bar{U}\leq 0|D=d)+P(\bar{Y}_{d}<c_{0}\leq 0,\bar{U}>0|D=d)$
$=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)\Lambda(d)+\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)(1-\Lambda(d))$
-- $\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)\Lambda(d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)(1-\Lambda(d))$
$=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)$
$=$ $P(c_{0}\leq Y\leq c_{1}|D=d).$
Hence, when the sharp equalities hold, there exists an alternative model to the classic tobit that can induce the observed data distribution. Indeed, the observed distribution of $Y$ and the non refuted marginal distribution of $U$, do not imply that we can learn the true distribution of $Y_{1},Y_{0},U$. A similar phenomena occurs with the iv-tobit mutatis mutandis.

### v2

@@
 We propose a testing procedure for the model’s validity using existing inference methods for intersection bounds.
-Simulation results suggests proper size for large samples and that the test is powerful to detect large violation of the exogeneity assumption and violations in the error structure.
-Finally, we review and propose new alternative paths to partially identify the parameters of interest under less restrictive assumptions.
+Simulation results suggest adequate test size for large samples and reasonable power in detecting violations of the exogeneity assumption and error structure.
+We review and propose new alternative paths to partially identify the parameters of interest under less restrictive assumptions.
+We revisit a study of married women’s labor supply in Lee (1995) to demonstrate the practical implementation and usefulness of the test.
+The present version is as of January 17, 2025, the first version is from December 2022.
+All errors are ours.
+Email address: otavio.bartalotti@monash.edu.
+We want to thank participants at seminars in the 2024 Jornadas Anuales de Economía del Banco Central del Uruguay and RMIT University, the 2024 IAAE and Southern Economic Association meetings.
 Keywords: Tobit models, hypothesis testing, Testable Implications, Instrumental variables.
@@
 Tobin (1958) analyzed household expenditure on durable goods using a regression model that specifically incorporated that expenditure (the dependent variable) cannot be negative.
-This approach is related to a broader class of models known as censored or truncated regression models depending on whether observations outside a specified range are lost or censored.
+This approach is related to a broader class of censored or truncated regression models, depending on whether observations outside a specified range are lost or censored.
 When applied researchers are interested in modelling limited dependent variables, potentially with mass accumulation points, the Tobit family of models provides structure to identify parameters of interest, such as the average treatment effect (ATE).
-Identification relies on three primary sources: (i) instrument exogeneity (or exogeneity of the variable of interest itself), (ii) normality of the model’s latent variables, and (iii), in the case of an instrumental variable approach to endogeneity, the relevance condition for the instrument.
+Identification relies on three primary sources: (i) instrument exogeneity (or exogeneity of the variable of interest itself), (ii) normality of the model’s latent variables, and in the case of an instrumental variable approach to endogeneity, (iii) the relevance condition for the instrument.
 While researchers recognize the model’s restrictive nature, it remains a valuable tool in the empirical literature.
-In this paper, we provide a test for the validity of the Tobit model’s structure and assumptions, providing three main contributions to the literature.
+In this paper, we develop a test for the validity of the Tobit model’s structure and assumptions, providing three main contributions to the literature.
 The first is to provide sharp testable equalities that can detect all possible observable violations of the Tobit model.
 Second, we propose a test for the validity of the Tobit model’s identifying assumptions using the sharp equalities that characterize the model to check its falsifiability.
-Following recent literature, we convert the equalities into conditional moment inequalities, and implement the test by existing inferential methods from Chernozhukov, Lee, and Rosen (2013).
+Following recent literature, we convert the equalities into conditional moment inequalities and implement the test by existing inferential methods from Chernozhukov, Lee, and Rosen (2013).
+The Tobit family of models involves continuous outcomes with accumulation points and potentially continuous treatment, implying a large number of moment equalities to be tested.
+This creates significant challenges to test implementation.
+We propose a discretization of the space of the treatment and the outcome that balances the computational requirements and data availability for different parts of their joint support.
+This simplifies the implementation, making it easy to compute and providing an asymptotically valid testing procedure.
 The third contribution is to review and propose alternative approaches that can be used when the model is rejected.
-We explore an alternative path to partially identify the parameter of interest by assuming monotonicity of selection into treatment.
-Finally, we provide empirical examples illustrating the methodology and its practical relevance.
+We explore an alternative path to partially identify the parameter of interest by assuming the monotonicity of selection into treatment.
+Finally, we provide an empirical example illustrating the methodology’s practical relevance.
 More generally, the current paper contributes to the growing literature on testing identifying assumptions of econometric models.
-We focus on two main models: (i) the “classic Tobit” model in which the main variable of interest is assumed to be exogenous and (ii) the IV-Tobit model.
+We focus on two main models: (i) the “classic Tobit” model in which the main variable of interest is assumed to be exogenous and (ii) the instrumental variable (IV) Tobit model.
 In both cases, the proposed test considers all observable violations of the model structure, such as linear index, normality of the latent errors, independence of the treatment, and homoskedasticity; with the addition of the validity of the instrument for the IV-Tobit.
-Additional results for variants of the Tobit family of models (e.g., type 2 Tobit) are presented in the Appendix.
+Additional results for variants of the Tobit family of models are presented in the Appendix.
 1.1. Previous Literature
-There is a vast literature related to testing the validity of Tobit models and their assumptions, to which we contribute.
+There is a vast literature related to testing the validity of Tobit models and their assumptions, which we contribute to.
 Most of the preceding work focuses on testing a particular assumption or feature of the model while maintaining other assumptions and structures as valid.
@@
 Nelson’s test compares the sample proportion of non-censored observations with the hypothesized probability of being non-censored in the Tobit model.
-As stated by Bera, Jarque, and Lee (1984), Nelson (1981) test is equivalent to the Lagrange multiplier test of the Tobit model against Cragg’s model (Lin and Schmidt, 1984).
-Bera, Jarque, and Lee (1984) proposes a Lagrange Multiplier test for the normality assumption against other distributions of the Pearson family of distributions while the remaining assumptions are maintained.
-Newey (1987) considers both exogenous and endogenous explanatory variables cases using symmetrically censored least squares estimators to construct specification tests of normality and homoskedasticity assumptions via a Hausman-type specification test.
-Holden (2004) examines several statistics proposed to test the normality assumption in the Tobit (censored regression) model and reinterpreted them as a version of the Lagrange multiplier (score) test for a common null hypothesis.
+Bera, Jarque, and Lee (1984) state that the test is equivalent to the Lagrange multiplier (LM) test of the Tobit model against Cragg’s model (Lin and Schmidt, 1984) and propose an alternative LM test for the normality assumption against other distributions of the Pearson family of distributions while the remaining assumptions are maintained.
+Newey (1987b) considers both exogenous and endogenous explanatory variables cases using symmetrically censored least squares estimators to construct specification tests of normality and homoskedasticity assumptions via a Hausman-type specification test.
+Holden (2004) examines several statistics proposed to test the normality assumption in the Tobit (censored regression) model and reinterpreted them as a version of the LM (score) test for a common null hypothesis.
 Reynolds and Shonkwiler (1991) use an information matrix misspecification test to detect violations of the distributional assumptions of the Tobit model.
 Other tests include Drukker (2002), which operationalized conditional moment tests developed by Newey (1985) and Tauchen (1985) to the case of misspecification of the distribution of the classic Tobit model.
-With a similar intuition to our framework, their test writes down conditional moment restrictions which should have zero conditional expected values under the null.
-The intuition behind the conditional moments test developed by Drukker (2002) is that since the model was estimated by maximum likelihood, the assumed data-generating process specifies all moments of disturbances conditional on the covariates to be the ones of a normal distribution.
+With a similar intuition to our framework, their test writes down conditional moment restrictions, which should have zero conditional expected values under the null.
+Since the model was estimated by maximum likelihood, the assumed data-generating process specifies the moments of disturbances conditional on the covariates to be the ones of a normal distribution.
 Drukker (2002) use these moment-based methods based on the third and fourth moments of the normal distribution.
-Our procedure, while intuitively similar detects all possible violations of the model, not only those evident from deviations in the third and fourth moments.
+While intuitively similar, our procedure detects all possible violations of the model, not only those evident from deviations in the third and fourth moments.
 Smith and Blundell (1986) propose a test of the treatment variable’s exogeneity in the IV-Tobit Model by a control function approach exploiting the joint normality of the latent variables.
-Most of these approaches focus on testing a particular subset of assumptions, or a specific class of alternatives to consider.
-We consider the full set of assumptions with a general alternative hypothesis, which is not a specific subset of models.
+Most of these approaches focus on testing a particular subset of assumptions or a specific class of alternatives to consider.
 The developments proposed in this paper consider all observable violations of the general Tobit model structure and its assumptions, serving as a useful test for empirical researchers constructing models for censored/truncated data.
 This work contributes to the growing literature on the testability of the identifying assumptions in various econometric models.
-Those include tests for the validity of assumptions in instrumental variable models, local average treatment effects, regression discontinuity designs, bivariate probit, etc (Mourifié and Wan, 2017; Kitagawa, 2015; Huber and Mellace, 2015; Gunsilius, 2021; Arai et al., 2022; Acerenza, Bartalotti, and Kédagni, 2023).
-Our testing procedures are connected to kedagni2020generalized which provides tests for assumptions related to the instrumental variables in the model, while we also consider the implications of the parametric structure on the outcome and latent error structure which are inherent to Tobit models.
+Those include tests for the validity of assumptions in instrumental variable models, local average treatment effects, regression discontinuity designs, bivariate probit, encouragement designs, among others.
+(Mourifié and Wan, 2017; Kitagawa, 2015; Huber and Mellace, 2015; Gunsilius, 2021; Arai et al., 2022; Acerenza, Bartalotti, and Kédagni, 2023; Bai and Tabord-Meehan, 2024).
+Our testing procedures are connected to Kédagni and Mourifié (2020), which provides tests for assumptions related to the instrumental variables in the model, while we also consider the implications of the parametric structure on the outcome and latent error structure, which are inherent to Tobit models.
+The two papers closest to ours are Acerenza, Bartalotti, and Kédagni (2023) and Goff, Kédagni, and Wu (2024).
+The first focuses on bivariate probit models, leading to a finite set of moment equalities as testable restrictions.
+As mentioned above, this study builds upon their work by tackling the case of the Tobit family of estimators, which provides additional technical and practical challenges given the continuous nature of outcome and treatment, including accumulation points.
+Goff, Kédagni, and Wu (2024) considers separable parametric instrumental variable models that induce a different set of moment equalities to be tested while our approach focuses on Tobit and IV-Tobit models, which can potentially be non-separable.
+[^2] While Goff, Kédagni, and Wu (2024) briefly mention the possibility of extending their results to non-separable models, they obtain different equalities from ours and do not discuss identification or sharpness.
+Implementation would also differ from our approach.
 The remainder of the paper is organized as follows.
@@
 Section 5 outlines the testing procedure.
-Section 6 include simulation evidence about the test’s size and power.
+Section 6 include simulation evidence about the test’s size while discussions about power are relegated to the Appendix.
 Section 7 discusses how to relax the assumptions in case of rejection.
-Section 8 and 9 provide empirical illustrations.
+Section 8 provides an empirical illustration revisiting the study of married women’s labor supply from Lee (1995) to demonstrate the practical implementation and usefulness of the test.
 Finally, Section 9 concludes.
-Additional results and details are collected in Appendices A-E.
+Additional results and details are collected in Appendices A-F.
 2. Models
-This section describes two popular “two-part” models for truncated data for which we derive testable implications in Section 3: the classic Tobit, and the instrumental variable Tobit.
-[^2] We omit conditioning on other exogenous covariates $X$ for simplicity of exposition.
+This section describes two popular “two-part” models for truncated data for which we derive testable implications in Section 3: the classic Tobit, and the IV-Tobit.
+[^3] We omit conditioning on other exogenous covariates ($X$) for simplicity of exposition.
+The results including covariates are presented in Section 5.
 Let $Y=\max(0,Y^{*})$ be the observed outcome taking values in $\mathcal{Y}\subset\mathbb{R}^{+}$ with $Y^{*}$ a latent continuous dependent variable taking values in $\mathcal{Y^{*}}\subset\mathbb{R}$.
@@
 $\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ \\ Y^{*}&=&\alpha_{0}+\alpha_{1}D+U\end{array}\right.$ (2.4)
-Where $U$ is an unobservable (latent) error.
-In addition to the model structure in Equation 2.4, the classic Tobit also considers restrictions on the distribution of the error term and its relationship to the treatment.
+where $U$ is an unobservable (latent) error. In addition to the model structure in system 2.4, the classic Tobit restricts the distribution of the error term and its relationship to the treatment.
 Assumption 1.
@@
 Assumption 2 imposes that the latent error has a standard normal distribution.
-[^4] The error could take $N(0,\sigma^{2})$ but after scale normalizations of the coefficients of $Y^{*}$ we can model it as $N(0,1)$.
-Despite its simplicity, the classic Tobit model can accommodate treatment that is continuous or discrete, and the normality assumption for the latent error could be replaced by other distributional assumptions that preserve the identification of the parameters.
+[^5] The error could take $N(0,\sigma^{2})$ but after scale normalizations we can model it as $N(0,1)$.
+Despite its simplicity, the classic Tobit model can accommodate continuous or discrete treatment, and the normality assumption for the latent error could be replaced by other distributional assumptions that preserve the identification of the parameters.
 2.2. IV Tobit model
 The IV Tobit considers the case in which the treatment of interest is endogenous, but an instrumental variable is available to identify its effect.
-Let the IV Tobit model be: [^5] Results for the IV-Tobit model are presented for continuous $D$ but can accommodate a binary treatment, $D=1\{\gamma_{0}+\gamma_{1}Z-V>0\}$.
+Let the IV Tobit model be: [^6] Results for the IV-Tobit model are presented for continuous $D$ but hold for a binary treatment, $D=1\{\gamma_{0}+\gamma_{1}Z-V>0\}$.
 In that case, the parameters are identified only up to scale.
-$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ \\ Y^{*}&=&\alpha_{0}+\alpha_{1}D+U\\ \\ D&=&\gamma_{0}+\gamma_{1}Z+V\end{array}\right.$ (2.10)
+$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\alpha_{0}+\alpha_{1}D+U\\ D&=&\gamma_{0}+\gamma_{1}Z+V\end{array}\right.$ (2.8)
 Alternatively, in its reduced form representation:
-$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ \\ Y^{*}&=&\beta_{0}+\beta_{1}Z+W\\ \\ D&=&\gamma_{0}+\gamma_{1}Z+V\end{array}\right.$ (2.16)
-Where $U,V$ are the latent structural error terms. To that model structure, we add restrictions on the joint distribution of the latent variables and their relationship to the instrumental variable.
+$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\beta_{0}+\beta_{1}Z+W\\ D&=&\gamma_{0}+\gamma_{1}Z+V\end{array}\right.$ (2.12)
+Where $U,V$ are the latent structural error terms.
+To that model structure, restrictions on the joint distribution of the latent variables and their relationship to the instrumental variable are added.
 Assumption 3.
-Assume that: $Z$ is independent of $(U,V)$ and $\gamma_{1}\neq 0$.
+$Z$ is independent of $(U,V)$ and $\gamma_{1}\neq 0$.
 Assumption 4.
@@
 This assumption also implies that $Z$ is independent of $W,V$, since $W=\alpha_{1}V+U$.
-Assumption 4 characterises the distribution of the latent vector of error terms as bivariate normal.
+Assumption 4 characterizes the distribution of the latent vector of error terms as bivariate normal.
 Then, $W,V$ also follows a bivariate normal distribution and $\alpha_{0},\alpha_{1},\gamma_{0},\gamma_{1}$ could be scale normalized so $(W,V)^{\prime}$ follows the standard bivariate normal distribution with covariance $\rho$, i.e., $\begin{pmatrix}W\\ V\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\end{pmatrix}$, and $\Sigma=\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}$.
-[^6] Note that if $U,V$ follow a bivariate normal distribution with covariance $\rho$, i.e., $\begin{pmatrix}U\\ V\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma=\begin{pmatrix}\sigma^{2}_{U}&\rho_{UV}\\ \rho_{UV}&\sigma^{2}_{V}\end{pmatrix}$, then $W,V$ follow a bivariate normal distribution $\begin{pmatrix}W\\ V\end{pmatrix}\sim\mathcal{N}(\mu_{1},\Sigma_{1})$, where $\mu_{1}=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma_{1}=\begin{pmatrix}\alpha_{1}^{2}\sigma^{2}_{V}+\sigma^{2}_{U}+2\rho_{UV}\alpha_{1}\sigma_{V}\sigma_{U}&\alpha_{1}\sigma_{V}^{2}+\rho_{UV}\\ \alpha_{1}\sigma_{V}^{2}+\rho_{UV}&\sigma^{2}_{V}\end{pmatrix}$.
+[^7] Note that if $U,V$ follow a bivariate normal distribution, $\begin{pmatrix}U\\ V\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma=\begin{pmatrix}\sigma^{2}_{U}&\rho_{UV}\\ \rho_{UV}&\sigma^{2}_{V}\end{pmatrix}$, then $W,V$ follow a bivariate normal distribution, $\begin{pmatrix}W\\ V\end{pmatrix}\sim\mathcal{N}(\mu_{1},\Sigma_{1})$, where $\mu_{1}=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma_{1}=\begin{pmatrix}\alpha_{1}^{2}\sigma^{2}_{V}+\sigma^{2}_{U}+2\rho_{UV}\alpha_{1}\sigma_{V}\sigma_{U}&\alpha_{1}\sigma_{V}^{2}+\rho_{UV}\\ \alpha_{1}\sigma_{V}^{2}+\rho_{UV}&\sigma^{2}_{V}\end{pmatrix}$.
 Then the reduced form system can be scale normalized.
-The latter characterisation aids the development of the sharp testable implications.
+The latter characterization aids the development of the sharp testable implications.
 3. Sharp testable equalities
 This section presents the sharp testable equalities for the models described in Section 2.
-Deviations from these equalities imply violations of the “null hypothesis” of the Tobit model (linear latent index, independence, instrument validity, and normality).
+Deviations from these equalities imply violations of the “null hypothesis” that the Tobit model (linear latent index, independence, instrument validity, and normality) is valid.
 The equalities are conditional on parameters being identified under the model’s assumptions, and a discussion on identification is postponed to Section 4.
@@
 Thus, the model’s testable implications require characterizing the distribution at both the mass point and beyond it.
-Starting at the continuous part of the distribution of $Y$, the conditional probabilities of $c_{0}\leq Y\leq c_{1}$ for a $c_{1},c_{0}>0$ are observed.
+Starting at the continuous part of the distribution of $Y$, the conditional probabilities of $c_{0}\leq Y\leq c_{1}$ are observed for $c_{1},c_{0}>0$.
 For any value of the treatment variables $d\in\mathcal{D}$:
@@
 $P(c_{1}\leq Y|D=d)$ $=$ $1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d).$ (3.2)
-Turning to the accumulation point, the observed event of $Y=0$,
-$P(Y=0|D=d)=P(Y^{*}\leq 0|D=d)=1-\Phi(\alpha_{0}+\alpha_{1}d).$ (3.3)
+Turning to the accumulation point, the observed event of $Y=0$ has probability,
+$P(Y=0|D=d)=P(Y^{*}\leq 0|D=d)=\Phi(-\alpha_{0}-\alpha_{1}d).$ (3.3)
 The first equality follows from the model’s structure in Equation 2.4, and the second one applies assumptions 1 and 2.
@@
 In the context of model 2.4, equalities 3.1-3.3 are sharp in the sense that whenever they hold, it is possible to construct a vector of $(\tilde{Y},D,\tilde{U})$ that satisfies model 2.4, Assumptions 1 and 2, and induces the observed distribution on the data $(Y,D)$.
-For the proof see Appendix A.
+For the proof, see Appendix A.
 Then, the equalities 3.1-3.3 are sharp testable implications for the validity of the classic Tobit Model in 2.4 coupled with assumptions 1-2.
@@
 The testable implications and sharpness discussed above show that the classic Tobit model is generally refutable.
-However, the model is non-verifiable.
-To prove this, we show that for any joint distribution on the data $(Y,D)$ that satisfies equalities 3.1-3.3, there exists a vector of $(\bar{Y},D,\bar{U})$ that induces the observed distribution on the data $(Y,D)$, but the vector $\bar{U}$ is not normally distributed.
+However, the model is non-verifiable in the sense that we can always construct a joint probability law of $(\bar{Y},D,\bar{U})$ that violates the Tobit model validity but satisfies equalities 3.1-3.3.
 See Appendix E for the proof.
 Remark 3 (Extensions).
-The previous derivation can be adjusted for different variations of Tobit models, such as generalizations of the distributional assumptions (barros2018generalized), different thresholds (carson2007Tobit), dynamic Tobit models, or including individual specific effects (wooldridge2005simple; honore2000estimation; honore1993orthogonality).
-In appendix C we derive the equalities for the type 2 Tobit model and similar logic can be applied to other two-part models.
+The previous derivation can be adjusted for different variations of Tobit models, such as generalizations of the distributional assumptions (Barros et al., 2018), different thresholds (Carson and Sun, 2007), dynamic Tobit models, or including individual specific effects (Wooldridge, 2005; Honore, Kyriazidou, and Powell, 2000; Honoré, 1993).
+In appendix C, we derive the equalities for the type 2 Tobit model, and similar logic can be applied to other two-part models.
 Additional testable results for the aforementioned models are discussed in Appendix D.
 3.2. IV Tobit model
-We turn our attention to the Instrumental Variable Tobit case and propose testable implications that can be used to test the model described in 2.10 and the associated assumptions 3 and 4.
-The observed data includes $(Y,D,Z)$, and we characterise the joint distribution of $(Y,D)\in\mathcal{Y}\times\mathcal{D}$ conditional on the instrument, $Z$, to obtain the model’s testable implications.
-The mapping from observed probabilities to its corresponding quantities implied by the IV Tobit model requires jointly evaluating the continuous (interior) support for the outcome and treatment variables, as well as the accumulation points and distribution tails conditional on $Z$.
+We turn our attention to the instrumental variable Tobit case and propose testable implications that can be used to test the model described in 2.8 and the associated assumptions 3 and 4.
+The observed data includes $(Y,D,Z)$, and we characterize the joint distribution of $(Y,D)\in\mathcal{Y}\times\mathcal{D}$ conditional on the instrument, $Z$, to obtain the model’s testable implications.
+The mapping from observed probabilities to their corresponding quantities implied by the IV Tobit model requires jointly evaluating the continuous (interior) support for the outcome and treatment variables, as well as the accumulation points and distribution tails conditional on $Z$.
 Consider any $0<c_{0}<c_{1},d_{0}<d_{1}$, and $z\in\mathcal{Y}\times\mathcal{D}\times\mathcal{Z}$,
@@
 $\equiv$ $\Phi^{1}_{c_{1},c_{0},d_{1}},$
-$P(c_{1}\leq Y,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z;\rho)-\Phi_{V}(d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.7)
+$P(c_{1}\leq Y,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z)-\Phi_{V}(d_{0}-\gamma_{0}-\gamma_{1}z)$ (3.7)
 -- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
@@
 $\equiv$ $\Phi^{1}_{c_{1},d_{1},d_{0}},$
-$P(c_{1}\leq Y,D\leq d_{0}|Z=z)$ $=$ $\Phi_{V}(d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.8)
+$P(c_{1}\leq Y,D\leq d_{0}|Z=z)$ $=$ $\Phi_{V}(d_{0}-\gamma_{0}-\gamma_{1}z)$ (3.8)
 -- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
 $\equiv$ $\Phi^{1}_{c_{1},d_{0}},$
-$P(c_{1}\leq Y,D\geq d_{1}|Z=z)$ $=$ $1-\Phi_{W}(c_{1}-\beta_{0}-\beta_{1}z)-\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.9)
+$P(c_{1}\leq Y,D\geq d_{1}|Z=z)$ $=$ $1-\Phi_{W}(c_{1}-\beta_{0}-\beta_{1}z)-\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z)$ (3.9)
 $+$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
 $\equiv$ $\Phi^{1}_{c_{1},d_{1}}.$
-where $\Phi_{W,V}(w,v;\rho)$ denotes the joint c.d.f. of $(W,V)$, a bivariate normal with coefficient of correlation $\rho$. Similarly, considering the case that the observed $Y$ equals zero,
+where $\Phi_{W,V}(w,v;\rho)$ denotes the joint c.d.f. of $(W,V)$, a standard bivariate normal with coefficient of correlation $\rho$. Similarly, considering the case that the observed $Y$ equals zero,
 $P(Y=0,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.10)
@@
 Remark 4 (Sharpness).
-In the context of model 2.10 or 2.16, equalities 3.4-3.12 for all $c,d,z\in\mathcal{Y}\times\mathcal{D}\times\mathcal{Z}$ are sharp in the sense that whenever they hold, it is possible to construct a vector of $(\tilde{Y},\tilde{D},\tilde{U},\tilde{V},Z)$ or equivalently $(\tilde{Y},\tilde{D},\tilde{W},\tilde{V},Z)$ that satisfies model 2.10 and 2.16, Assumptions 3 and 4, and induces the observed distribution on the data $(Y,D,Z)$.
+In the context of model 2.8 or 2.12, equalities 3.4-3.12 for all $c,d,z\in\mathcal{Y}\times\mathcal{D}\times\mathcal{Z}$ are sharp in the sense that whenever they hold, it is possible to construct a vector of $(\tilde{Y},\tilde{D},\tilde{U},\tilde{V},Z)$ or equivalently $(\tilde{Y},\tilde{D},\tilde{W},\tilde{V},Z)$ that satisfies model 2.8 and 2.12, Assumptions 3 and 4, and induces the observed distribution on the data $(Y,D,Z)$.
 See Appendix B for the proof.
-Then, the equalities 3.4-3.12 are sharp testable implications for the validity of the Instrumental Variable Tobit Model in 2.10 coupled with assumptions 3-4.
+Remark 5 (Non-learnability).
+The testable implications and sharpness discussed above show that the IV-Tobit model is generally refutable.
+However, the model is non-verifiable.
+The demonstration follows a similar logic as in the classic Tobit case.
+See Appendix E for the proof.
+Then, the equalities 3.4-3.12 are sharp testable implications for the validity of the instrumental variable Tobit Model in 2.8 coupled with assumptions 3-4.
 They serve as the basis for the test procedure described in Section 5.
-A similar claim of non-learnability holds for the iv-tobit.
-The demonstration follows a similar logic.
 4. Identification
@@
 The coefficients have been normalized by dividing by the latent variable’s ($U$) square root of the variance, denoted by $\sigma$.
-Naturally, this implies that the ratio of the original parameters, say $\tilde{\alpha}_{0},\tilde{\alpha}_{1}$ and $\sigma$ is identified since $P(Y=0|D=d)=\Phi(\alpha_{0}+\alpha_{1}D)=\Phi(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D)$.
-Similarly to the case analysed by Han and Vytlacil (2017), this allows the identification of the inverse mills ratio which is a known function of the standard normal density, $\lambda(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D))$.
+Naturally, this implies that the ratio of the original parameters is identified since $P(Y=0|D=d)=\Phi(\alpha_{0}+\alpha_{1}D)=\Phi(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D)$.
+This allows the identification of the inverse mills ratio, which is a known function of the standard normal density, $\lambda(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D))$.
 Then,
@@
 5. Testing procedure
-To test the sharp equalities, we rewrite each of these equalities as two inequalities (Mourifié and Wan, 2017; Acerenza, Bartalotti, and Kédagni, 2023).
-For a concrete example, we first note that for example
-$P(c_{1}\leq Y|D=d)=1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)\iff E[1\{c_{1}\leq D\}-1+\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]=0.$
-We then we rewrite each of these moment equalities implied by the restrictions on the empirical distribution as moment inequalities
-$E[1\{c_{1}\leq D\}-1+\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]\leq 0,E[-1\{c_{1}\leq D\}+1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]\leq 0.$
-In doing this, for all the restrictions on the empirical distribution we can implement a test relying on existing intersection bounds inferential methods such as Chernozhukov, Lee, and Rosen (2013) which is specifically suited to test conditional moment inequalities.
-Note that the equalities for the classic Tobit need to hold for any pair of constants $(c_{0},c_{1})$, and the ones from the IV Tobit need to hold for pairs of $(c_{0},c_{1})$ and $(d_{0},d_{1})$.
+To test the sharp equalities, we rewrite each of these equalities as two inequalities (Mourifié and Wan, 2017; Acerenza, Bartalotti, and Kédagni, 2023). For a concrete example, we first note that
+$P(c_{1}\leq Y|D=d)=1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)\iff E[1\{c_{1}\leq Y\}-1+\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]=0.$
+We rewrite each of these moment equalities implied by the restrictions on the empirical distribution as moment inequalities
+$E[1\{c_{1}\leq Y\}-1+\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]\leq 0,E[-1\{c_{1}\leq Y\}+1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]\leq 0.$
+In doing this for all the restrictions on the empirical distribution, we can implement a test relying on existing intersection bounds inferential methods such as Chernozhukov, Lee, and Rosen (2013), which is specifically suited to test conditional moment inequalities.
+Note that the equalities for the classic Tobit hold for any pair of constants $(c_{0},c_{1})$, and the ones from the IV Tobit hold for pairs of $(c_{0},c_{1})$ and $(d_{0},d_{1})$.
 We propose a partition of $\mathcal{Y}\times\mathcal{D}$ to test sufficient conditions of these sharp equalities.
-Rejection of any of the null hypotheses that the equalities hold at these particular levels will imply violations of the model.
+Rejection of any of the null hypotheses that the equalities hold at these particular levels implies violations of the model.
 Classic Tobit model
@@
 $\hat{\theta}_{1-\alpha}\equiv\max_{k}\sup_{D}\left\{\hat{\theta}(D,k)-\kappa_{1-\alpha}\hat{s}(D,k)\right\}>0,$ (5.3)
-where $\hat{\theta}(D,k)$ is a nonparametric estimator for $\theta(D,k)$, $\hat{s}(D,k)$ its standard error, and $\kappa_{1-\alpha}$ is a critical value at the significance level $\alpha$.
+where $\hat{\theta}(D,k)$ is a nonparametric estimator for $\theta_{k}(D)$, $\hat{s}(D,k)$ its standard error, and $\kappa_{1-\alpha}$ is a critical value at the significance level $\alpha$.
 IV Tobit model
 For the instrumental variable Tobit model, the continuous support for both the outcome and treatment poses challenges to the implementation of the test.
+[^8] Note that a discrete treatment can also be accommodated as an intermediate step between the current derivation and the derivations from Acerenza, Bartalotti, and Kédagni (2023).
 The sharp testable equalities for every $c_{0},c_{1}\in\mathcal{Y}$ and $d_{0},d_{1}\in\mathcal{D}$ are given by equations 3.4-3.12.
@@
 Let $W_{kq}$ be given by,
-$W_{kq}=\left\{\begin{array}[]{cc}1\{Y=0\}1\{D\leq d_{0}\}-\Phi^{2}_{d_{0}},&\text{ if }k=0,q=0\\ 1\{Y=0\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{2}_{d_{q},d_{q+1}},&\text{ if }k=0,1\leq q<Q\\ 1\{Y=0\}1\{D\geq d_{Q}\}|-\Phi^{2}_{d_{Q}},&\text{ if }k=0,q=Q\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{D\leq d_{0}\}-\Phi^{1}_{c_{k+1},c_{k},d_{0}},&\text{ if }1\leq k<K,q=0\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{1}_{c_{k+1},c_{k},d_{q+1},d_{q}},&\text{ if }1\leq k<K,1\leq q<Q\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{D\geq d_{Q}\}-\Phi^{1}_{c_{k+1},c_{k},d_{Q}},&\text{ if }1\leq k<K,q=Q\\ 1\{Y\geq c_{K}\}1\{D\leq d_{0}\}-\Phi^{1}_{c_{K},d_{0}},&\text{ if }k=K,q=0\\ 1\{Y\geq c_{K}\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{1}_{c_{K},d_{q+1},d_{q}},&\text{ if }k=K,1\leq q<Q\\ 1\{Y\geq c_{K}\}1\{D\geq d_{Q}\}-\Phi^{1}_{c_{K},d_{Q}},&\text{ if }k=K,q=Q.\\ \end{array}\right.$
+$W_{kq}=\left\{\begin{array}[]{cc}1\{Y=0\}1\{D\leq d_{0}\}-\Phi^{2}_{d_{0}},&\text{ if }k=0,q=0\\ 1\{Y=0\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{2}_{d_{q},d_{q+1}},&\text{ if }k=0,1\leq q<Q\\ 1\{Y=0\}1\{D\geq d_{Q}\}-\Phi^{2}_{d_{Q}},&\text{ if }k=0,q=Q\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{D\leq d_{0}\}-\Phi^{1}_{c_{k+1},c_{k},d_{0}},&\text{ if }1\leq k<K,q=0\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{1}_{c_{k+1},c_{k},d_{q+1},d_{q}},&\text{ if }1\leq k<K,1\leq q<Q\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{D\geq d_{Q}\}-\Phi^{1}_{c_{k+1},c_{k},d_{Q}},&\text{ if }1\leq k<K,q=Q\\ 1\{Y\geq c_{K}\}1\{D\leq d_{0}\}-\Phi^{1}_{c_{K},d_{0}},&\text{ if }k=K,q=0\\ 1\{Y\geq c_{K}\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{1}_{c_{K},d_{q+1},d_{q}},&\text{ if }k=K,1\leq q<Q\\ 1\{Y\geq c_{K}\}1\{D\geq d_{Q}\}-\Phi^{1}_{c_{K},d_{Q}},&\text{ if }k=K,q=Q.\\ \end{array}\right.$
 Where we used the simplifying notation defined in equations (3.4)-(3.12) for each partition of the support for $Y$ and $D$.
@@
 $\hat{\theta}_{1-\alpha}\equiv\max_{k,q}\sup_{Z}\left\{\hat{\theta}(Z,k,q)-\kappa_{1-\alpha}\hat{s}(D,k,q)\right\}>0,$ (5.6)
-where $\hat{\theta}(D,k)$ is a nonparametric estimator for $\theta(D,k)$, $\hat{s}(D,k)$ its standard error, and $\kappa_{1-\alpha}$ is a critical value at the significance level $\alpha$.
+where $\hat{\theta}(Z,k,q)$ is a nonparametric estimator for $\theta_{kq}(Z)$, $\hat{s}(D,k,q)$ its standard error, and $\kappa_{1-\alpha}$ is a critical value at the significance level $\alpha$.
 To implement the test within the Chernozhukov, Lee, and Rosen (2013) intersection bounds inferential method, we use the CLR Stata package described in Chernozhukov et al.
@@
 Remark 6.
-Intuitively, the testable conditions derived above consider if the empirical conditional distribution of the observed outcome variable - in both the mass accumulation and non-truncated parts of the support of $Y$ is consistent with a random variable(s) following the (bivariate) normal distribution for different sections of the distribution and values of the independent instrument, $Z$.
+Intuitively, the testable conditions derived above consider whether the empirical conditional distribution of the observed outcome variable — in both the mass accumulation and non-truncated parts of the support of $Y$ — is consistent with a random variable(s) following the (bivariate) normal distribution for different sections of the distribution and values of the independent instrument, $Z$.
 The proposed test procedures are intended to detect violations of the model due to:
@@
 3.
-Violations due to the empirical distributions of the residuals differing from the implied distributions in certain values of the treatment (Classic Tobit) or instrument (IV Tobit), which indicate violations of the exogeneity of treatment or instrument (kedagni_generalized_2020).
-Remark 6 (On the inclusion of covariates).
-The existing procedure can be extended to include exogenous covariates, $X$, within the linear index model in 2.10.
-Then, the testable equalities for both classic and IV Tobit models could be derived with the additional conditioning on $X$.
-To implement the test with covariates, we suggest generating a partition of the covariate space, say in $J$ grids, similarly to the partition of the exogenous variable used in Section 5.
-For every grid in the partition, one could compute the test statistic 5.3 or 5.6 for the classic or the IV-tobit, respectively.
+Violations due to the empirical distributions of the residuals differing from the implied distributions in certain values of the treatment (Classic Tobit) or instrument (IV Tobit), which indicate violations of the exogeneity of treatment or instrument (Kédagni and Mourifié, 2020).
+Including Covariates
+The proposed procedure can be extended to include exogenous covariates, $X$, within the linear index model in 2.8.
+Then, the testable equalities for classic and IV Tobit models can be derived with the additional conditioning on $X$.
+The test with covariates could be implemented by generating a partition of the covariate space, say in $J$ grids, similar to the partition of the exogenous variable used in Section 5.
+For every grid in the partition, one computes the test statistic 5.3 or 5.6 for the classic or the IV-tobit, respectively.
 Then obtain critical values that account for multiple testing via a Bonferroni correction.
 In particular, for every grid, set the critical value at the significance level $\frac{\alpha}{J}$, namely $\kappa_{1-\frac{\alpha}{J}}$.
-A theoretically interesting approach would partition the joint support of all the exogenous variables and compute the test statistics across all parts of the grid.
-However, this approach would entail significant implementation challenges, since the statistical packages for Chernozhukov et al.
-(2015) allow for only one exogenous covariate.
+[^9] A theoretically interesting approach would partition the joint support of all the exogenous variables and compute the test statistics across all parts of the grid.
+However, this approach would entail significant implementation challenges since the statistical package for Chernozhukov et al.
+(2015) allows for only one exogenous covariate.
+This procedure can be cumbersome when there are many covariates or when they are continuous.
+Another route that is less computationally intensive follows Acerenza, Bartalotti, and Kédagni (2023).
+We illustrate it for the $IV$-tobit case, and a similar logic could be applied to the classic Tobit case.
+Thus, with covariates added to the linear index function, we have,
+$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\beta_{0}+\beta_{1}Z+\beta_{2}X+W\\ D&=&\gamma_{0}+\gamma_{1}Z+\gamma_{2}X+V\end{array}\right.$ (5.10)
+We extend Assumption 3 to formalize covariates’ exogeneity:
+Assumption 5 (full independence).
+[…]
+Under Assumptions 4-5, by a similar argument to that used in Section 3, any given generic testable implication becomes
+$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z,X=x)$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{1}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
+-- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{1}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
+-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{0}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
+$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{0}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
+for all $z\in\mathcal{Z}$ and $x\in\mathcal{X}$, where $x$ can be a vector. Equivalent conditions for the case that $Y=0$ and other parts of the support of $Y$ and $D$ can be similarly obtained.
+Implementation
+Since the test is carried out nonparametrically, we face challenges when the dimension of $X$ is large, especially with continuous covariates.
+Another operational limitation is that the Stata’s clrbound package only allows for one conditioning variable at a time.
+For those practical reasons, we propose implementing a weaker, non-sharp, version of the testable equalities.
+Continuing with the example of Equality 5, we can integrate over the covariates $X$ (or alternatively, $Z$), taking advantage of the fact that $\mathbb{E}[W|Z=z,X=x]=0$ implies $\mathbb{E}[W|Z=z]=0$ and $\mathbb{E}[W|X=x]=0$ for all random variables $W$.
+Implementation becomes very similar to the IV-Tobit discussed above by redefining the simplifying notation in equations (3.4)-(3.12), with the only difference being the inclusion of $X$ on the linear indexes in $\Phi_{W}(\cdot)$ and $\Phi_{W,V}(\cdot,\cdot,\cdot)$.
+For example,
+$\Phi^{1}_{c_{1},c_{0},d_{1},d_{0}}$$\equiv$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{1}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
+-- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{1}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
+-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{0}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
+$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{0}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho),$
+and similarly for the other terms.
+Then, compute the new $W_{kq}$ in the same manner as in Section 5, and the intersection bounds framework considers the similar $2(K+1)(Q+1)$ inequalities, based on the partition of the support of $Y$ and $D$.
+$\sup_{z}E[W_{kq}|Z=z]$$\leq$ 00 0
+$\sup_{z}E[-W_{kq}|Z=z]$$\leq$ $0,\text{ for }k=0,\dots,K;q=0,\dots,Q.$
+This is the test procedure that we implement for the empirical application in Section 8.
+Under Assumption 5, one could base the test in conditioning on a particular covariate $X_{C}$ instead of $Z$ by integrating the sharp Equality 5 over $Z$ and then implementing the intersection bounds procedure using $X_{C}$ as the sole conditioning variable.
+[^10] Assumption 5 constraints the relationship between the covariates and the latent error terms in such a way that one could combine the information on all $X$ and $Z$ in an index and construct non-sharp testable equalities that would hold conditional on this index.
+For example, having $\sup_{t}E[W_{kq}|Z+X_{1}+...+X_{C}=t]=0$, for $k=0,\dots,K;q=0,\dots,Q$.
 6. Simulations
-In this section, we provide simulation exercises for the size and power of the proposed test for the classical Tobit model assumptions.
-The testing procedure described relies on testing sharp equalities that should hold for any arbitrary partition.
-Since both the outcome and treatment are continuous, evaluating all possible equalities is technically challenging.
-We focus on a non-sharp set of the equalities by evaluating them at different grid partitions of their support, in a similar spirit to honore2020selection.
+In this section, we provide simulation exercises for the proposed tests.
+The testing procedure described relies on testing sharp equalities that should hold for any arbitrary partition of the outcome, treatment and instrument support.
+When both the outcome and treatment are continuous, evaluating all possible equalities is technically challenging.
+We focus on a non-sharp set of the equalities by evaluating them at different grid partitions of their support, in a similar spirit to Honoré and Hu (2020) which is particularly well suited for continuous supports.
 Naturally, this choice makes the test less powerful as we don’t consider the continuum of equalities derived in Section 4, but is justified by the ease of implementation of the procedure based on intersection bounds and the performance of the test on the simulations below.
-For the simulations presented below, we considered a partition of the support of the observed outcome variable into the accumulation point ($Y=0$) and four quartiles on the (untruncated) positive range.
-The choice of the number and location of the partitions/evaluation points balances the implementation computational requirements, availability of data for different parts of the joint support of the outcome, treatment and exogenous instrumental variable.
+For the simulations related to the classic Tobit, we partition the support of the observed outcome variable into the accumulation point ($Y=0$) and four quartiles on the (untruncated) positive range, while for the IV-Tobit case we also partition the support of the treatment variable ($D$) to create a grid based on both the outcome and treatment.
+[^11] Note that if the treatment variable was discrete, the respective grid points could be set naturally to the possible countable values $D$ takes.
+The choice of the number and location of the partitions/evaluation points balances the implementation computational requirements, data availability for different parts of the joint support of the outcome, treatment and exogenous instrumental variable.
 For the procedure in Section 5 to be feasible, we must have data on both the outcome and the exogenous variables within each partition.
-Using the empirical quantiles of the non-truncated outcomes to determine the partitions guarantees a reasonable number of observations for each part of the grid.
-Larger sample sizes might allow finer partitions for the outcome support to be used, but the added computational requirements created by an increased number of equalities being checked, coupled with the larger datasets, can substantially increase computing time.
-[^7] The simulations presented in Section 6 have limited sample sizes and a relatively small number of equalities being tested due to the long running time and computational constraints when repeating the test procedure thousands of times.
+Using the empirical quantiles of the non-truncated outcomes to determine the partitions guarantees a reasonable number of observations for each grid part.
+Larger sample sizes might allow finer partitions for the outcome support.
+Still, the added computational requirements created by an increased number of equalities being checked, coupled with the larger datasets, can substantially increase computing time.
+[^12] The simulations presented in Section 6 have limited sample sizes and a relatively small number of equalities being tested due to the long-running time and computational constraints when repeating the test procedure thousands of times.
 When implementing the intersection bounds in STATA using the package clrbound Chernozhukov et al.
@@
 To guarantee the feasibility of the procedure, we adjust the evaluation points for $D$ ($Z$) to the first and $99^{th}$ percentiles of the exogenous variable in each partition of the support for the outcome $Y$ [^13] See details on implementation on the simulation replication STATA code.
-6.1. Classic Tobit model
-6.1.1. Size
-For the classic Tobit model, we consider a continuous treatment $D$ following the data generating process described in Equation 6.8.
+In evaluating the finite sample performance of the proposed tests, we consider a continuous treatment $D$ following the data generating process described in Equation 6.8.
 $\left\{\begin{array}[]{lcl}Y&=&\max\{Y^{*},0\}\\ Y^{*}&=&D+U\\ D&=&2Z-V\\ \begin{pmatrix}U\\ V\\ Z\end{pmatrix}&\sim&\mathcal{N}(\mathbf{0}_{3},\mathbf{\Sigma})\\ \mathbf{\Sigma}&=&\begin{pmatrix}1&\rho_{uv}&\rho_{uz}\\ \rho_{uv}&1&\rho_{vz}\\ \rho_{uz}&\rho_{vz}&1\end{pmatrix}\\ \rho_{uz}&=&\rho_{vz}=0\\ \rho_{uv}&=&\rho\end{array}\right.$ (6.8)
@@
 Under the treatment exogeneity condition described in Assumption 1, $\rho=0$.
-Table 1 presents the empirical test sizes for this scenario, under different significance levels $\alpha$.
+Table 1 presents the empirical test sizes for the Classic Tobit test this scenario, for different significance levels $\alpha$.
 The results indicate that while the test over rejects the null hypothesis for small to mid-sized samples, the test’s empirical coverage approaches its desired nominal benchmark as samples larger than 5,000 are used.
-Table 1. Test Size
+Table 1. Classic Tobit Test Size
 N $\alpha=10\%$ $\alpha=5\%$ $\alpha=1\%$
@@
 Based on $500$ replications.
-6.1.2. Power
-To consider the test’s performance under violations of the exogeneity assumption, we use the DGP described in Equation 6.8 with different values for $\rho$, reflecting various degrees of endogeneity.
+To consider the test’s performance under violations of the exogeneity assumption, we modify the DGP in Equation 6.8 with different values for $\rho$, reflecting various degrees of treatment endogeneity.
 Larger values of $\rho$ produce more acute violations of the null hypothesis.
-Table 2 presents results for these scenarios: columns 3-4 present the MSE while the remaining columns present the empirical rejection rates.
-As expected, the power of the test increases with larger $\rho$ and bigger sample sizes, presenting substantial power for $\alpha=10\%$ and presenting a decreasing pattern for $\alpha=5\%$ and $\alpha=1\%$.
-Table 2. Test power for violations in exogeneity
+Table 2 presents the results for the Classic Tobit test: columns 3-4 report the MSE, while the remaining columns show empirical rejection rates.
+As expected, the power of the test increases with larger $\rho$ and bigger sample sizes.
+[^14] Simulations for test performance when violations of normality of the latent errors occur are presented in Appendix F.
+Table 2. Classic Tobit test power for violations in exogeneity
 $N$ $\rho$ MSE$(\beta_{0})$ MSE$(\beta_{1})$ $\alpha=10\%$ $\alpha=5\%$ $\alpha=1\%$
@@
 Based on $500$ replications.
-Turning our attention to the test’s power under violations of normality of the latent error, we consider the modified data generating process in Equation 6.15.
-In this DGP, $F(\cdot)$ denotes an alternative density for the latent error $U$, which we vary across scenarios.
-Table 3 present the empirical coverage of the tests when $F(\cdot)$ is uniform, lognormal and t-student distributions.
-$\left\{\begin{array}[]{lcl}Y&=&\max\{Y^{*},0\}\\ Y^{*}&=&D+U\\ D&=&2Z-V\\ \begin{pmatrix}V\\ Z\end{pmatrix}&\sim&\mathcal{N}(\mathbf{0}_{2},\mathbf{\Sigma})\\ \mathbf{\Sigma}&=&\begin{pmatrix}1&0\\ 0&1\\ \end{pmatrix}\\ U\sim F(\cdot)\end{array}\right.$ (6.15)
-As expected, the rejection rate for the case of the t-student with 80 degrees of freedom is close to the nominal size of the test, since it represents a very mild violation of the normality assumption.
-In the other three cases, there is a strong violation of the normality assumption and the test rejects the null hypothesis in most instances.
-This example provides indication that the proposed test can be useful in detecting violations of the distributional assumptions in the classic Tobit model.
-Table 3. Test power for violations in error structure
-$N$ $U$ MSE$(\beta_{0})$ MSE$(\beta_{1})$ $\alpha=10\%$ $\alpha=5\%$ $\alpha=1\%$
-5000 t-student (df=80) 0.9987 1.0018 7.21% 4.01% 1.40%
-t-student (df=5) 1.027 1.0624 92.20% 81.80% 48.80%
-LogNormal 1.3336 0.0896 100.00% 100.00% 100.00%
-Uniform 0.892 0.785 100.00% 100.00% 100.00%
-8000 t-student (df=80) 0.9999 1.0034 7.00% 4.80% 1.20%
-t-student (df=5) 1.0251 1.0653 99.20% 97.80% 87.58%
-LogNormal 1.3385 0.0866 100.00% 100.00% 100.00%
-Uniform 0.8932 0.7896 100.00% 100.00% 100.00%
-10000 t-student (df=80) 1.0016 1.0059 5.42% 2.41% 1.00%
-t-student (df=5) 1.0241 1.0642 99.20% 98.40% 90.58%
-LogNormal 1.3348 0.0868 100.00% 100.00% 100.00%
-Uniform 0.8908 0.7862 100.00% 100.00% 100.00%
+Naturally, researchers concerned about treatment endogeneity should consider the IV-Tobit model and implement the test of its identifying assumptions proposed in Section 5.
+Table 3 presents the empirical coverage for the test of the IV-Tobit model for different levels of treatment endogeneity ($\rho=\{0,0.5,0.8\}$) for sample sizes 5,000 and 8,000.
+The test produces adequate empirical coverage, in line with the results for the classic Tobit test.
+Table 3. IV Tobit - Test Size
+N $\rho_{uv}$ $\alpha=10\%$ $\alpha=5\%$ $\alpha=1\%$
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Based on $500$ replications.
 7. Relaxation of the assumptions
-When the test proposed in Section 5 rejects the null hypothesis of the Tobit model’s validity, researchers must pursue alternative models and less restrictive assumptions to learn confidently about the parameter of interest.
-7.1. Existing Solutions
+When the test proposed in Section 5 rejects the null hypothesis of the model’s validity, researchers must pursue alternative models and less restrictive assumptions to learn confidently about the parameters of interest.
+7.1. Alternative Approaches
 There is a vast literature on alternatives to the Tobit Model that can be implemented in the presence of censored dependent variables.
-Most approaches consider changes or relaxations of one of the two main assumptions associated with the Tobit model.
+Most approaches consider changes or relaxations of one of the two main assumptions associated with the model.
 The first assumption is the parametric distribution of the error terms and latent index form connecting treatment (and covariates) to the outcome.
@@
 Here, we provide a non-exhausting survey of existing work.
-cragg1971some maintains the normality of the errors, linearity of the index and exogeneity but relaxes the way censoring occurs in comparison to the latent structure of the censored outcome.
+Cragg (1971) maintains the normality of the errors, linearity of the index and exogeneity but relaxes the way censoring occurs in comparison to the latent structure of the censored outcome.
 Specifically, while the latent outcome is still modelled by $Y^{*}=\alpha_{0}+\alpha_{1}D+U$, they allow for the censoring to depend on a different linear index, $P(Y^{*}>0)=P(\gamma_{0}+\gamma_{1}D+e)$, increasing the model’s flexibility.
-POWELL1984303 relaxes the parametric structure of the errors while maintaining the latent linear index and treatment exogeneity, and estimates the parameters of interest by least absolute deviations.
-powell1986symmetrically also maintains linearity of the latent index and treatment exogeneity, but relaxes normality by imposing symmetrical distributions to the latent errors, which leads to estimation by symmetrically censored least squares.
-newey1987efficient relaxes exogeneity of the treatment, relying on normality and an instrumental variable to identify the model, which is estimated by generalized least squares.
-honore1994pairwise relax linearity and mean independence of the unobservable with respect to the treatment to exploit the idea that, although $Y^{*}_{i}-\alpha_{0}-\alpha_{1}D_{i}$ is not mean-independent of $D_{i}$, one can trim any pair of residuals $Y^{*}_{i}-\alpha_{0}-\alpha_{1}D_{i}$ and $Y^{*}_{j}-\alpha_{0}-\alpha_{1}D_{j}$, and the trimmed residuals are independent and identically distributed conditional on $D_{i},D_{j}$.
+Powell (1984) relaxes the parametric structure of the errors while maintaining the latent linear index and treatment exogeneity, and estimates the parameters of interest by least absolute deviations.
+Powell (1986) also maintains linearity of the latent index and treatment exogeneity, but relaxes normality by imposing symmetrical distributions to the latent errors, which leads to estimation by symmetrically censored least squares.
+Newey (1987a) relaxes exogeneity of the treatment, relying on normality and an instrumental variable to identify the model, which is estimated by generalized least squares.
+Honoré and Powell (1994) relax linearity and mean independence of the unobservable with respect to the treatment to exploit the idea that, although $Y^{*}_{i}-\alpha_{0}-\alpha_{1}D_{i}$ is not mean-independent of $D_{i}$, one can trim any pair of residuals $Y^{*}_{i}-\alpha_{0}-\alpha_{1}D_{i}$ and $Y^{*}_{j}-\alpha_{0}-\alpha_{1}D_{j}$, and the trimmed residuals are independent and identically distributed conditional on $D_{i},D_{j}$.
 They estimate the model by identically censored least absolute deviations and identically censored least squares (ICLS).
-das2002estimators estimates a model using symmetrically censored least squares that relaxes exogeneity of the treatment and normality of the errors.
+Das (2002) estimates a model using symmetrically censored least squares that relaxes exogeneity of the treatment and normality of the errors.
 To achieve that they rely on instrumental variables, linearity of the mean of the structural error conditional on the reduced form error, and mean independence of the reduced form error.
-blundell2007censored proposes a control variable approach that relaxes exogeneity and normality but maintains the latent linear structure ($\alpha_{0}+\alpha_{1}D+U$).
+Blundell and Powell (2007) proposes a control variable approach that relaxes exogeneity and normality but maintains the latent linear structure ($\alpha_{0}+\alpha_{1}D+U$).
 Crucially, they impose that the distribution (or quantiles) of the latent error conditional on the treatment and instrument is only a function of the control variable $V=D-\pi(Z)$, which isolates the endogenous variation on the treatment.
@@
 This allows them to estimate the effect of the treatment by censored quantile instrumental variable regression augmented by a control variable given by the quantiles of $U$ conditional on $V$ at the quantile of interest.
-In a similar spirit, chernozhukov2015quantile focuses on conditional quantile functions and flexible approaches to estimate the control variable in the first-stage.
-Finally, chesher2023iv provide partial identification results for a general alternative by relaxing exogeneity of the treatment and instrument, linearity of the latent index and imposing no parametric structure of the error term.
-They characterize the identified set for the parameters of interest following the Generalized Instrumental Variables framework (chesher2017generalized) , relying on the assumption that the relationship of $Y^{*}$ to treatment and errors is continuous and monotonic in the errors.
+In a similar spirit, Chernozhukov, Fernández-Val, and Kowalski (2015) focuses on conditional quantile functions and flexible approaches to estimate the control variable in the first stage.
+Finally, Chesher, Kim, and Rosen (2023) provides partial identification results for a general alternative by relaxing the exogeneity of the treatment and instrument, linearity of the latent index and imposing no parametric structure of the error term.
+They characterize the identified set for the parameters of interest following the Generalized Instrumental Variables framework (Chesher and Rosen, 2017), relying on the assumption that the relationship of $Y^{*}$ to treatment and errors is continuous and monotonic in the errors.
 Their approach uses the residual sets associated with the structure of the latent function and conditional probability of the error term given potential instruments.
@@
 In this subsection, we present an approach that partially identifies the effect of an endogenous treatment variable by replacing the normality and exogeneity assumptions with a monotonicity in treatment selection constraint.
-While less general than chesher2023iv, this approach is easy to implement and could be useful to empirical researchers.
+While less general than Chesher, Kim, and Rosen (2023), this approach is easy to implement and could be useful to empirical researchers.
 Consider the model that maintains linearity (or a known structure of $Y^{*}$ up to a finite number of parameters),
@@
 Let $E(U|D=d,Y>0)\equiv\Gamma(d)$. Then, for any $d>d^{*}$ we either have $\Gamma(d)<\Gamma(d^{*})$ or $\Gamma(d)>\Gamma(d^{*})$.
-Assumption 5 is common in the partial identification literature (jiang2014monotone; manski1997monotone; manski1998monotone; okumura2014concave).
-In this context, we restrict the latent selection to behave monotonically with the treatment.
+Assumption 6 is common in the partial identification literature (Jiang, Chiba, and VanderWeele, 2014; Manski, 1997; Manski and Pepper, 2000; Okumura and Usui, 2014).
+In this context, we restrict the latent selection to be monotonic with respect to the treatment.
 This assumption is embedded in the classic Tobit model since the inverse mills ratio, $\lambda(\cdot)$, is monotonic (and decreasing) in the treatment variable.
@@
 Under Assumption 6 and the model described in 7.3, treatment and outcome are not independent. Note that,
-$E(Y|D=d,Y>0)$ $=$ $\alpha_{0}+\alpha_{1}d+\Gamma(d)$ (7.4)
-$E(Y|D=d,Y>0)$displaystyle-- $\alpha_{0}-\alpha_{1}d=\Gamma(d)$ (7.5)
+$E(Y|D=d,Y>0)$ $=$ $\alpha_{0}+\alpha_{1}d+\Gamma(d)$
+$E(Y|D=d,Y>0)$displaystyle-- $\alpha_{0}-\alpha_{1}d=\Gamma(d)$
 Then, for any two $d>d^{*}$ we have, by Assumption 6:
 $\begin{array}[]{lcl}\Gamma(d)&<&\Gamma(d^{*})\\ &\Leftrightarrow&\\ E(Y|D=d,Y>0)-\alpha_{0}-\alpha_{1}d&<&E(Y|D=d^{*},Y>0)-\alpha_{0}-\alpha_{1}d^{*}\\ &\Leftrightarrow&\\ \alpha_{1}&>&\frac{E(Y|D=d^{*},Y>0)-E(Y|D=d,Y>0)}{d^{*}-d},\end{array}$ (7.9)
-which implies a lower bound on parameter interest.
+which implies a lower bound on the parameter interest.
 For a binary treatment $D\in\{0,1\}$ the lower bound is, intuitively, the difference in average outcomes between treated and untreated individuals away from the mass point at zero:
@@
 One-sided simple confidence regions can be computed based on these outer sets of the treatment effect.
-One can estimate $E(Y|D=d,Y>0)$ using its sample analogs, and rely on their asymptotic normality.
+One can estimate $E(Y|D=d,Y>0)$ using its sample analogs and rely on their asymptotic normality.
 For example, let the estimators be given by $\widehat{E}(Y|D=d,Y>0)$.
 By the continuous mapping theorem, $\frac{\widehat{E}(Y|D=d^{*},Y>0)-\widehat{E}(Y|D=d,Y>0)}{d^{*}-d}$ is asymptotically normal.
-Thus, a one-sided confidence interval for $\frac{E(Y|D=d^{*},Y>0)-E(Y|D=d,Y>0)}{d^{*}-d}$ can be computed via bootstrap which implies a conservative estimate for the lower bound for $\alpha_{1}$.
+Thus, a one-sided confidence interval for $\frac{E(Y|D=d^{*},Y>0)-E(Y|D=d,Y>0)}{d^{*}-d}$ can be computed via bootstrap, which implies a conservative estimate for the lower bound for $\alpha_{1}$.
 A bootstrap procedure could be used for a nonparametric estimator of $\frac{\partial E(Y|D=d,Y>0)}{\partial d}$ using a local polynomial regression.
-In this case the estimator is asymptotically normal and converges at a nonparametric rate that depends on the bandwidth $h$.
-Recent developments in calonico2018effect; calonico2022coverage for estimation and optimal coverage error bandwidth and kernel selection methods that are nonparametric robust bias-corrected (RBC) can be used through their convenient implementation using the package nprobust (calonico2019nprobust).
-8. Empirical illustration: Olson1998
-In this section, we implement the test of model validity described in Section 5 and the alternative method for partial identification under monotonicity of treatment selection using data from Olson1998.
-Employer-provided health insurance is one of the most important mechanisms by which individuals under the age of 65 gain access to health care in the USA.
-Health insurance coverage insures individuals and their families against the financial risk associated with the costs of treating health conditions.
-Olson1998 hypothesize that women who lack health insurance coverage through their partner’s employer are more likely to work full-time to obtain health insurance for themselves and other family members.
-Using data from the Current Population Survey (CPS) from March 1993, Olson1998 finds that women without spousal coverage were less likely to be out of the labor force than those with partner’s coverage.
-The 1993 CPS refers to hours worked and health insurance benefits provided by the longest job held during 1992.
-Olson1998 created the study’s sample by matching spouse’s records on household ID codes to create a single record for each married couple in the survey.
-Next, they excluded couples for which the wife was over the age of 64 or her labor force status at the time of the interview in March 1993 was different from her status in 1992.
-Finally, households were eliminated if respondents were assigned imputed values for the relevant health insurance questions or the husband’s total earnings in 1992.
-This produced a sample of 22,272 families.
-For 15,620 families, the wife’s usual weekly hours worked in 1992 were greater than zero.
-Olson1998 is interested in the effect of having health insurance coverage through the partner’s employment ($D$) on women’s labor supply measured as the usual hours worked per week ($Y$).
-A Tobit model was estimated using this labor supply measure, and a dummy indicating insurance coverage through the husband’s job.
-Table 4 presents the results for model 2.4 under assumptions 1-2.
-Table 4. Tobit specification for Olson1998
-MLE
-Hours Worked
-Wife covered through partner’s job $-0.6876071^{*}$
-$(0.3637667)$
-Constant $21.70212^{***}$
-$(0.2866817)$
-$N$ $22,272$
-Standard errors (in parentheses); ***: significant at 1% level, * significant at 10% level.
-The first column presents estimates indicating that having health insurance through the husband’s employment has a negative effect on women’s hours worked.
-However, the model is rejected at conventional significance levels when we test for the Tobit model’s validity as proposed in Section 5 ($\hat{\theta}_{0.99}=0.4491>0$, $\hat{\theta}_{0.95}=0.4517>0$, and $\hat{\theta}_{0.90}=0.4529>0$).
-Confidence Bounds under Assumption 5 and latent linear index
-Given the rejection of the Tobit model in this case, we relax the distributional and exogeneity of treatment assumptions, and construct lower bounds on the treatment effect for the effect of receiving health insurance through partner’s employment on women’s labor supply under the MTS Assumption 5 and latent linear index only.
-In the context of this application, Assumption 5 imposes that average unobservables affecting women’s preferences related to hours worked away from home, such as schedule flexibility and reservation wages, are monotonically increasing in characteristics that drive husbands to take jobs that cover spouses.
-For example, households that prioritize flexible schedules for women might similarly prioritize partner jobs that provide more generous non-wage benefits such as family health insurance coverage.
-Table 5. Confidence sets for parameter of interest
-Parameter Tobit estimates 95% conf. sets for the LB for $\alpha_{1}$
-$\alpha_{1}$0.6876071 [-0.8146481, -0.5687435]
-conf.: confidence; LB: lower bound; UB: upper bound.
-The first column of Table 5 repeats the estimate for $\alpha_{1}$ from Table 4.
-The second column reports the 95% confidence set for the lower bound for $\alpha_{1}$, obtained under Assumption 5 and latent linear index only.
-Recall that the lower end of the confidence set is the lower bound for the treatment effect of interest $\alpha_{1}$.
-Even after relaxing the normality of errors and exogeneity of the treatment, we obtain informative bounds for the size of the effect of having employer-provided health insurance through their partner’s employment on the female labor supply.
-These estimates indicate an effect lower bound of -0.8146 hours worked per week for married women who work positive hours - the intensive margin of labor supply.
-While we can rule out a reduction of one hour per week in female labor supply, we cannot rule out that the effect is zero.
-9. Empirical Illustration: lee1995semi
-In this section we implement the proposed test to the data from lee1995semi.
-Using the 1987 cross-section of the Michigan Panel Study of Income Dynamics, the authors study the responses of married women’s labor supply ($Y$) - measured in hours per year - to “other” household income ($D$), which is endogenous.
-The potential instrumental variables explored are dummy variables for the husband’s occupation ($Z$), which implies the restrictive identifying assumption that the wife’s labor supply is affected by the husbands’ occupation only through their income.
-Following lee1995semi, we proceed with the analysis focusing on the data for married couples with non-negative family total income and where the wife was of working age (18-64) and not self-employed.
-The number of observations is 3382 and 26 percent of wage observations are censored.
-Table LABEL:Table:Lee presents the estimates obtained using IV Tobit model.
-Table 6. IV Tobit specification for lee1995semi
+In this case, the estimator is asymptotically normal and converges at a nonparametric rate that depends on the bandwidth $h$.
+Recent developments in Calonico, Cattaneo, and Farrell (2018, 2022) for estimation and optimal coverage error bandwidth and kernel selection methods that are nonparametric robust bias-corrected (RBC) can be used through their convenient implementation using the package nprobust (Calonico, Cattaneo, and Farrell, 2019).
+Remark 7 (On including covariates).
+If we impose that $Y^{*}=\alpha_{0}+\alpha_{1}D+\alpha_{2}X+U$ the procedure can include exogenous covariates by modifying Assumption 6 to hold conditional on $X$. Then,
+$\Gamma(d,X)$ $<$ $\Gamma(d^{*},X)$
+$\Leftrightarrow$
+$E(Y|D=d,X,Y>0)-\alpha_{0}-\alpha_{1}d-\alpha_{2}X$ $<$ $E(Y|D=d^{*},X,Y>0)-\alpha_{0}-\alpha_{1}d^{*}-\alpha_{2}X$
+$\Leftrightarrow$
+$\alpha_{1}$ $>$ $\frac{E(Y|D=d^{*},X,Y>0)-E(Y|D=d,X,Y>0)}{d^{*}-d}.$
+A similar argument holds when the treatment variable is continuous.
+Tighter bounds are achieved by considering the maximum value of $\frac{\partial E(Y|D=d,X=x,Y>0)}{\partial d}$ across all possible points in the support for $D$ and $X$.
+8. Empirical Illustration: Lee (1995)
+In this section, we implement the proposed test to the data from Lee (1995).
+Using the 1987 cross-section of the Michigan Panel Study of Income Dynamics, the authors study the responses of married women’s labor supply ($Y$) - measured in hours per year - to hundreds of dollars in “other” household income ($D$), which is endogenous.
+The instrumental variables explored are dummy variables for the husband’s occupation ($Z$), which implies the restrictive identifying assumption that the wife’s labor supply is affected by the husband’s occupation only through their income.
+Following the original study, we add other covariates ($X$) in the linear index for both the outcome and treatment equations, controlling for factors that could impact women’s labor supply.
+Those include a quadratic on the person’s age, their years of completed education, the number of children coded in three categories (children up to 5 years old, ages 6 to 13, ages 14 to 17), the local unemployment rate in percentage points, and indicators for race (0 if white, 1 otherwise), homeownership (1 if owner, 0 otherwise) and if the couple has a mortgage on their home (1 if yes, 0 otherwise).
+Our empirical illustration considers as an instrument the binary variable indicating if the husband’s occupation is classified as manager or professional.
+[^16] The 1987 PSID uses 3-digit occupation codes from the 1970 U.S.
+Census, and this dummy variable seems to include workers listed in the categories “1-195 Professional, Technical, and Kindred Workers,” and “201–245 Managers and Administrators, except Farm.”
+Following Lee (1995), we proceed with the analysis focusing on the data for married couples with non-negative family total income or “other” household income and where the wife was of working age (18-64) and not self-employed.
+These selections results in 3,277 observations, for which 26 percent of wage observations are censored.
+Table LABEL:Table:Lee presents the estimates obtained using the IV Tobit model.
+Table 4. IV Tobit specification for Lee (1995)
 MLE
 Other household income Hours per year worked
-Husband’s occupation: manager or professional $198.3894^{***}$
+Husband’s occupation: manager or professional $120.802^{***}$
 $(10.813)$
-Other household income $0.7029415^{***}$
-(.$(.237801)$
-Constant $240.4661^{***}$ $752.9173^{***}$
+Other household income $-0.973^{***}$
+$(0.373)$
+Age $13.686^{***}$ $72.414^{***}$
 $(3.251)$ $(15.249)$
-$\rho$ $-0.3069913^{***}$
-(.$(.0572254)$
-$n$ $3,382$ $3,382$
-Standard errors (in parentheses); ***: significant at 1% level.
+Age squared $-0.105^{***}$ $-1.221^{***}$
+$(0.039)$ $(0.178)$
+Education $20.281^{***}$ $92.107^{***}$
+$(2.071)$ $(13.231)$
+Children under 5 $9.448$ $-500.332^{***}$
+$(6.482)$ $(28.163)$
+Children between 6 and 13 $3.204$ $-211.687^{***}$
+$(5.601)$ $(23.737)$
+Children between 14 and 17 $12.881$ $-16.878$
+$(9.342)$ $(39.152)$
+Nonwhite $-59.550^{***}$ $146.336^{***}$
+$(10.210)$ $(51.105)$
+Homeowner $60.591^{***}$ $13.830$
+$(15.461)$ $(69.142)$
+Has mortgage $24.192^{*}$ $254.660^{***}$
+$(13.954)$ $(61.036)$
+Local Unemployment $-8.165^{***}$ $-41.979^{***}$
+$(1.941)$ $(8.966)$
+Constant $-339.059^{***}$ $-337.347$
+$(63.387)$ $(314.207)$
+$\rho$ $0.042$
+$(0.094)$
+$n$ $3,377$ $3,377$
+Standard errors (in parentheses); ***: significant at 1% level; *significant at 10% level
 The first column presents the first-stage estimates indicating the relevance of the potential $IV$.
 The second column presents the structural equation reflecting the effect the treatment variable has on the outcome equation.
-The parameter $\rho$ shows evidence of a negative correlation between the unobservables driving the other household income and hours worked.
-The estimated coefficient of interest indicates that other household income has a positive effect on the wife’s labor supply.
-However, the model is rejected at conventional significance levels when we test for the IV-Tobit model’s validity, ($\hat{\theta}_{0.99}=0.0247848>0$, $\hat{\theta}_{0.95}=0.0270805>0$, and $\hat{\theta}_{0.90}=0.0283793>0$).
+The parameter $\rho$ shows evidence of no correlation between the unobservables driving the “other” household income and hours worked after controlling for covariates.
+The estimated coefficient of interest indicates that other household income negatively affects the wife’s labor supply after conditional on the household characteristics.
+In particular, for women working positive hours, an increase of one thousand dollars in household income from other sources is estimated to reduce hours worked by 9.7 hours per year.
+The direction of the impacts at the intensive margin of hours worked follow intuitive patterns and are qualitative similar to those in Lee (1995).
+However, the model is rejected at conventional significance levels when we test for the IV-Tobit model’s validity, ($\hat{\theta}_{0.99}=0.1335>0$, $\hat{\theta}_{0.95}=0.1368>0$, and $\hat{\theta}_{0.90}=0.1385>0$).
+This indicates that the assumptions underlying the IV-Tobit model are not compatible with the empirical distribution of the data, and caution is needed when relying on the results.
+Bounds under Assumption 6 and latent linear index
+Given the rejection of the IV-Tobit model in this case, we relax the distributional and exogeneity of treatment assumptions, and construct lower bounds on the treatment effect under the MTS assumption and latent linear index only.
+Assumption 6 imposes that average unobservables affecting women’s preferences related to hours worked away from home, are monotonically decreasing in characteristics leading to higher income of other sources in the household, such as husband’s income.
+In other words, households that prefer flexible schedules for women might similarly prioritize partner jobs that provide higher income.
+Table 5. Confidence sets for parameter of interest
+Parameter IV-Tobit estimates $\alpha_{1}$’s Lower Bound
+$\alpha_{1}$0.973 -0.419
+conf.: confidence; LB: lower bound
+The first column of Table 5 repeats the estimate for $\alpha_{1}$ from Table 4.
+The second column reports the estimated lower bound for $\alpha_{1}$, obtained under Assumption 6 and latent linear index only based on the nonparametric estimate of the average derivative of the conditional expectation for yearly hours worked with respect to the household income from other sources, as described in Remark 7.
+Even after relaxing the normality of errors and treatment exogeneity, the lower bound for the size of the effect of having higher household income from sources other than the wife’s labor on their labor supply indicates an effect larger than -4.19 hours worked per year for married women.
+Hence, we can rule out annual reductions of more than 4.2 hours in female labor supply for every one thousand dollars in other household income, but cannot reject that the effect is zero or positive.
 9. Conclusion
-In this paper, we develop sharp testable equalities for Tobit-type “two-part” models that can detect all observable violations of the model’s assumptions.
+In this paper, we develop sharp testable equalities for the classic Tobit and IV-Tobit models that can detect all observable violations of the model’s assumptions.
+The results are shown to extend to many other popular Tobit-type “two-part” models.
 By converting these sharp equalities into conditional moment inequalities, we propose a testing procedure that detects violations of the Tobit model assumptions on a grid on the joint support of the outcome (and treatment) variables leveraging inference results from Chernozhukov, Lee, and Rosen (2013) and the implementation from Chernozhukov et al.
@@
 Simulation results suggest the test performs well for reasonably sized samples (larger than 5000 observations).
-For smaller samples, the test is conservative, overejecting the null hypothesis of model validity.
+The test is conservative for smaller samples, overejecting the null hypothesis of model validity.
 Simulations indicate that the test is powerful to detect violations of the exogeneity assumption for the treatment/instrument that affect the point estimates and inference.
@@
 We provide a brief review of existing models that could be implemented under weak/alternative assumptions when the Tobit model is rejected.
-Furthermore, we propose a simple model that partially identifies the parameter of interest by relying solely on linear index and monotone treatment selection restrictions, a standard assumption from the partial identification literature (manski2000amonotone).
-We illustrate our methods on data from Olson1998 and lee1995semi.
-We conclude that stylized versions (excluding covariates) of both empirical strategies reject the null hypothesis of a Tobit model and an IV-Tobit model respectively.
-Furthermore when illustrating partial identification strategy on Olson1998 we obtain informative lower bounds for the effect of having employer-provided health insurance through their partner’s employment on the female labor supply.
-These estimates indicate an effect lower bound of $-0.8146$ hours worked per week for married women working positive hours - the intensive margin of labor supply.
-While we can rule out a reduction of one hour per week in female labor supply, we cannot rule out that the effect is zero.
+Furthermore, we propose a simple model that partially identifies the parameter of interest by relying solely on linear index and monotone treatment selection restrictions, a standard assumption from the partial identification literature (Manski and Pepper, 2000).
+We illustrate our methods on data from Lee (1995).
+We replicate qualitatively the results in the original paper and the proposed test for validity of the IV-Tobit model rejects the null hypothesis in this empirical application.
+We estimate our proposed lower bound for the effect of household income from sources other than the wife’s labor on their labor supply, which does not rely on the normality of latent errors or treatment exogeneity.
+While we can rule out that an extra 1,000 dollars in other household income reduce female labor supply by more than 4.2 hours per year, we cannot rule out that the effect is zero.
 Appendix A Sharpness for the classic Tobit case
@@
 Also note $f_{\tilde{U}}$ is the $N(0,1)$ density, thus assumption 2 holds.
-Considering the positive values for $\tilde{Y}$, for any constants $c_{0},c_{1}$ such that $0<c_{0}<c_{1}$:
-$P$ $(c_{0}\leq\tilde{Y}\leq c_{1}|D=d)=P(c_{0}\leq\tilde{Y}^{*}\leq c_{1}|D=d)=P(c_{0}\leq\alpha_{0}+\alpha_{1}d+\tilde{U}\leq c_{1}|D=d)$
-$=$ $P(c_{0}-\alpha_{0}-\alpha_{1}d\leq\tilde{U}\leq c_{1}-\alpha_{0}-\alpha_{1}d|D=d)=P(c_{0}-\alpha_{0}-\alpha_{1}d\leq\tilde{U}\leq c_{1}-\alpha_{0}-\alpha_{1}d)$
-$=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)=P(c_{0}\leq Y\leq c_{1}|D=d).$
+Considering the positive values for $\tilde{Y}$, for any constants $c_{0},c_{1}$ such that $0<c_{0}<c_{1}$, then,
+$P(c_{0}\leq\tilde{Y}\leq c_{1}|D=d)$ $=$ $P(c_{0}\leq\tilde{Y}^{*}\leq c_{1}|D=d)$
+$=$ $P(c_{0}\leq\alpha_{0}+\alpha_{1}d+\tilde{U}\leq c_{1}|D=d)$
+$=$ $P(c_{0}-\alpha_{0}-\alpha_{1}d\leq\tilde{U}\leq c_{1}-\alpha_{0}-\alpha_{1}d|D=d)$
+$=$ $P(c_{0}-\alpha_{0}-\alpha_{1}d\leq\tilde{U}\leq c_{1}-\alpha_{0}-\alpha_{1}d)$
+$=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)$
+$=$ $P(c_{0}\leq Y\leq c_{1}|D=d).$
 The first three equalities follow from the definitions of $\tilde{Y}^{*}$ and $\tilde{Y}$.
@@
 $=$ $P(c_{0}-\beta_{0}-\beta_{1}z\leq W\leq c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq V\leq d_{1}-\gamma_{0}-\gamma_{1}z,Y^{*}\geq 0|Z=z)$
-$=$ $P(c_{0}-\beta_{0}-\beta_{1}z\leq W\leq c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq V\leq d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
+$=$ $P(c_{0}-\beta_{0}-\beta_{1}z\leq W\leq c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq V\leq d_{1}-\gamma_{0}-\gamma_{1}z)$
 $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)-\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
@@
 The first equality follows from the law of total probability.
-The second through fourth equalities are in consequence of the model structure in 2.16.
+The second through fourth equalities are in consequence of the model structure in 2.12.
 The fifth step uses Assumption 3 and $c_{0}\geq 0$.
@@
 $P(Y=0,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $P(Y^{*}\leq 0,d_{0}\leq D\leq d_{1}|Z=z)$ (B.2)
-$=$ $P(W\leq-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}\leq V\leq d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
+$=$ $P(W\leq-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}\leq V\leq d_{1}-\gamma_{0}-\gamma_{1}z)$
 $=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
@@
 $=$ $P(c_{0}-\beta_{0}-\beta_{1}z\leq\tilde{W}\leq c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq\tilde{V}\leq d_{1}-\gamma_{0}-\gamma_{1}z|Z=z)$
-$=$ $P(c_{0}-\beta_{0}-\beta_{1}z\leq\tilde{W}\leq c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq\tilde{V}\leq d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
+$=$ $P(c_{0}-\beta_{0}-\beta_{1}z\leq\tilde{W}\leq c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq\tilde{V}\leq d_{1}-\gamma_{0}-\gamma_{1}z)$
 $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)-\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
@@
 $=$ $P(\tilde{W}\leq-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq\tilde{V}\leq d_{1}-\gamma_{0}-\gamma_{1}z|Z=z)$
-$=$ $P(\tilde{W}\leq-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq\tilde{V}\leq d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
+$=$ $P(\tilde{W}\leq-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq\tilde{V}\leq d_{1}-\gamma_{0}-\gamma_{1}z)$
 $=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)-\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
@@
 Appendix C Type 2 Tobit testable implications
-In this section we derive the results of the Type 2 tobit model also known as selection models or Heckman selection type of models (heckman1979sample).
+In this section we derive the results of the Type 2 tobit model also known as selection models or Heckman selection type of models (Heckman, 1979).
 The basic setup with no covariates (which can be extended in several directions and with different distributional assumptions as well as to incorporate treatment endogeneity, as we pointed out for the classic Tobit model in remark 3) is:
@@
 $=$ $P(V<-\gamma_{0}-\gamma_{1}z)$
-$=$ $\Phi_{v}(\frac{-\gamma_{0}-\gamma_{1}z}{\sigma_{V}})$
-Where the first and the second equality is due to the structure of the model described in equation C.8, the third one is due to Assumption 6 and the last one is due to the normalization for normal random variables and Assumption 7.
+$=$ $\Phi_{v}\left(\frac{-\gamma_{0}-\gamma_{1}z}{\sigma_{V}}\right)$
+Where the first and the second equalities are due to the structure of the model described in equation C.8.
+The third step is due to Assumption 7 and the last one follows from the normalization for normal random variables and Assumption 8.
 Additionally,
@@
 $=$ $\Phi_{U}(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi_{U}(c_{0}-\alpha_{0}-\alpha_{1}d)$
--- $\Phi_{U,V}(c_{1}-\alpha_{0}-\alpha_{1}d,-\gamma_{0}-\gamma_{1}z)$
-$+$ $\Phi_{U,V}(c_{0}-\alpha_{0}-\alpha_{1}d,-\gamma_{0}-\gamma_{1}z)$
+-- $\Phi_{U,V}(c_{1}-\alpha_{0}-\alpha_{1}d,-\gamma_{0}-\gamma_{1}z,\rho_{UV})$
+$+$ $\Phi_{U,V}(c_{0}-\alpha_{0}-\alpha_{1}d,-\gamma_{0}-\gamma_{1}z,\rho_{UV})$
 The first equality follows from the law of total probability.
-The second through fifth equalities follow from the model’s structure described in Equation C.8, namely $P(Y>0,Y^{*}<0|D=d)=0$ and $Y^{*}=Y,\text{not missing}$ for $S=1$ and the structure of $S$ as well as the latent linear model.
-The sixth one is due to Assumption 6 The seventh one is due to Assumption 7 as well as properties of normal random variables as well as bivariate cummulative distribution functions.
-Thus, we can test in a similar fashion as we did in the classic Tobit case and the IV-Tobit case but using the previous two type of equalities in addition to the ones of the form $P(Y\geq c_{2}|D=d,Z=z)$ which can be easily derived.
+The second through fifth equalities follow from the model’s structure described in Equation C.8, namely $P(Y>0,Y^{*}<0|D=d)=0$ and $Y^{*}=Y,\text{not missing}$ for $S=1$, the structure of $S$, and the latent linear model.
+The sixth step is due to Assumption 7.
+The last equality uses Assumption 8 as well as properties of normal random variables.
+Thus, we can construct a test similarly to the one in the main text, using these equalities in addition to the ones of the form $P(Y\geq c_{2}|D=d,Z=z)$.
 Appendix D Results from Remark 3
-In this appendix we derive the testable implications of the extensions discussed in Remark 3. We will discuss testable implications assuming identification of the relevant parameters or distributions.
-D.1. barros2018generalized and carson2007Tobit
-barros2018generalized proposes a variant of the tobit model with Elliptically contoured distributions and a non-zero threshold.
-At the same time carson2007Tobit proposes a tobit model with a non-zero threshold.
-In this section we combine both type of results and report the testable implications with a generic non-zero threshold and a generic known or identifiable parametric distribution function.
+In this appendix we derive the testable implications of the extensions to the classic Tobit and IV-Tobit discussed in Remark 3.
+We discuss testable implications assuming the identification of the relevant parameters or distributions.
+D.1. Barros et al. (2018) and Carson and Sun (2007)
+Barros et al.
+(2018) proposes a variant of the tobit model with elliptically contoured distributions and a non-zero threshold.
+At the same time Carson and Sun (2007) proposes a Tobit model with a non-zero threshold.
+In this section we combine both types of results and report the testable implications with a generic non-zero threshold and a generic known or identifiable parametric distribution function.
 In addition, extend the latent structure to be also known or an identifiable function up to a vector of parameters but invertible in $U$.
@@
 $P(Y=0|D=d)=P(Y^{*}\leq\tau|D=d)=F_{H}(g^{-1}(\alpha,d,\tau);\omega).$ (D.6)
-Which can then be used to test in a similar way as did for the classic tobit or iv-tobit by adding the type of equalities $P(Y\geq c_{2}|D=d)$ which can be derived in a similar fashion.
-Also, note it is easy to extend the previous development to cases where the truncation takes the form $Y=Y^{*}$ if $\tau_{0}\leq Y^{*}\leq\tau_{1}$, $Y=\tau_{0}$ if $\tau_{0}\geq Y^{*}$ and $Y=\tau_{1}$ if $Y^{*}\geq\tau_{1}$.
-D.2. wooldridge2005simple; honore2000estimation; honore1993orthogonality
-Consider the following dynamic version of the tobit model which is related to wooldridge2005simple; honore2000estimation; honore1993orthogonality.
-Here we specify the conditional behaviour of $c_{i}$ in the spirit of wooldridge2005simple and others.
+These equalities can be used to construct a test in a similar way as did for the classic Tobit and IV-Tobit by adding the type of equalities $P(Y\geq c_{2}|D=d)$ which can be derived in a similar fashion.
+Also, note it is possible to extend the previous development to cases where the truncation takes the form $Y=Y^{*}$ if $\tau_{0}\leq Y^{*}\leq\tau_{1}$, $Y=\tau_{0}$ if $\tau_{0}\geq Y^{*}$ and $Y=\tau_{1}$ if $Y^{*}\geq\tau_{1}$.
+D.2. Wooldridge (2005); Honore, Kyriazidou, and Powell (2000); Honoré (1993)
+Consider the following dynamic version of the Tobit model, which is related to Wooldridge (2005); Honore, Kyriazidou, and Powell (2000); Honoré (1993).
+Here, we specify the conditional behaviour of $c_{i}$ in the spirit of Wooldridge (2005) and others.
 $\left\{\begin{array}[]{lcl}Y_{i,t}&=&\max(0,Y_{i,t}^{*})\\ \\ Y_{i,t}^{*}&=&\alpha_{0}+\alpha_{1}D_{i,t}+\alpha_{2}g(Y_{i,t-1})+c_{i}+U_{i,t}\end{array}\right.$ (D.10)
@@
 Assumption 12.
-$U_{i,t}|Y_{i,t-1},c_{i}$ is distributed $N(0,1)$.
+$U_{i,t}|Y_{i,t-1},c_{i}$ follows a $N(0,1)$ distribution.
 Assumption 13.
@@
 […]
-Which can then be used to test in a similar way as did for the classic tobit or iv-tobit by adding the type of equalities $P(Y_{i,t}\geq c_{2}|D_{i,t}=d,Y_{i,t-1}=y)$, with the caveat that the right hand side will need to be numerically integrated or approximated by an estimator when no close form is available.
-Such estimator should assure that the whole left hand side converges at root-$N$ in order for the estimation of the the null model to be asymptotically negligble(Acerenza, Bartalotti, and Kédagni, 2023, Appendix B).
+Which can then be used to construct a test in a similar way as presented in the main text by adding the type of equalities $P(Y_{i,t}\geq c_{2}|D_{i,t}=d,Y_{i,t-1}=y)$, with the caveat that the right-hand side requires numerical integration or an approximation by an estimator when no closed form is available.
+Such estimator should ensure that the left-hand side of the equality converges at root-$N$ in order for the estimation step of the null model to be asymptotically negligible (Acerenza, Bartalotti, and Kédagni, 2023, Appendix B).
 Appendix E Non-learnability
@@
 Suppose that,
-$\left\{\begin{array}[]{lcl}\bar{Y}&=&\max(0,\bar{Y}^{*})\\ \\ \bar{Y}^{*}&=&\alpha_{0}+\alpha_{1}D+\bar{U}\end{array}\right.$ (E.4)
+$\left\{\begin{array}[]{lcl}\bar{Y}&=&\max(0,\bar{Y}^{*})\\ \\ \bar{Y}^{*}&=&G(\alpha_{0}+\alpha_{1}D+\bar{U})\end{array}\right.$ (E.4)
+Where $G(.)$ is some continuous and strictly increasing function such that $G^{-1}(0)=0$ and $\bar{U}$ is a random variable that is $N(0,1)$.
 For $c_{0}\leq Y\leq c_{1}$ for a $c_{1},c_{0}>0$ and for any value of the treatment variables $d\in\mathcal{D}$, assume that the equalities (3.1)-(3.3) hold:
@@
 $P(Y=0|D=d)$ $=$ $1-\Phi(\alpha_{0}+\alpha_{1}d).$
-Now define $\bar{Y}_{d}=\alpha_{0}+\alpha_{1}d+\bar{U}$ and define the following distribution function:
-$P(\bar{Y}_{d}\leq c_{1},\bar{U}\leq 0|D=d)=\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)\Lambda(d)$
-$P(\bar{Y}_{d}\leq c_{1},\bar{U}>0|D=d)=\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)(1-\Lambda(d))$
-Where $\Lambda(d)=\frac{e^{d}}{1+e^{d}}$. In the above proposed DGP, $\bar{U}$ is not normally distributed. However, the DGP is compatible with the data since,
-$P(\bar{Y}=0|D=d)$ $=$ $P(\bar{Y}^{*}\leq 0|D=d)=P(\alpha_{0}+\alpha_{1}d+\bar{U}\leq 0|D=d)$ (E.5)
-$=$ $P(\alpha_{0}+\alpha_{1}d+\bar{U}\leq 0,\bar{U}\leq 0|D=d)$
-$+$ $P(\alpha_{0}+\alpha_{1}d+\bar{U}\leq 0,\bar{U}>0|D=d)$
-$=$ $P(Y_{d}\leq 0,\bar{U}\leq 0|D=d)+P(Y_{d}\leq 0,\bar{U}>0|D=d)$
-$=$ $\Phi(-\alpha_{0}-\alpha_{1}d)\Lambda(d)+\Phi(-\alpha_{0}-\alpha_{1}d)(1-\Lambda(d))$
-$=$ $\Phi(-\alpha_{0}-\alpha_{1}d)=1-\Phi(\alpha_{0}+\alpha_{1}d)=P(Y=0|D=d)$
-where the last equality is due to the testable equalities holding. Similarly, for any $c_{1}>0$,
-$P(c_{1}<\bar{Y}|D=d)$ $=$ $P(c_{1}<\bar{Y}^{*}|D=d)=P(c_{1}<\alpha_{0}+\alpha_{1}d+\bar{U}|D=d)$ (E.6)
-$=$ $P(c_{1}<\bar{Y}_{d}|D=d)=1-P(\bar{Y}_{d}<c_{1}|D=d)$
-$=$ $1-P(\bar{Y}_{d}<c_{1},\bar{U}\leq 0|D=d)-P(\bar{Y}_{d}<c_{1}\leq 0,\bar{U}>0|D=d)$
-$=$ $1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)\Lambda(d)+\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)(1-\Lambda(d))$
-$=$ $1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)=P(c_{1}<Y|D=d).$
-Finally,
-$P(c_{0}\leq\bar{Y}\leq c_{1}|D=d)$ $=$ $P(c_{0}\leq\bar{Y}^{*}\leq c_{1}|D=d)=P(c_{0}<\alpha_{0}+\alpha_{1}d+\bar{U}<c_{1}|D=d)$ (E.7)
-$=$ $P(c_{0}<\bar{Y}_{d}<c_{1}|D=d)=P(\bar{Y}_{d}<c_{1}|D=d)P-(\bar{Y}_{d}<c_{0}|D=d)$
-$=$ $P(\bar{Y}_{d}<c_{1},\bar{U}\leq 0|D=d)+P(\bar{Y}_{d}<c_{1}\leq 0,\bar{U}>0|D=d)$
--- $P(\bar{Y}_{d}<c_{0},\bar{U}\leq 0|D=d)+P(\bar{Y}_{d}<c_{0}\leq 0,\bar{U}>0|D=d)$
-$=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)\Lambda(d)+\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)(1-\Lambda(d))$
--- $\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)\Lambda(d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)(1-\Lambda(d))$
-$=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)$
+Now define $\bar{Y}_{d}=G(\alpha_{0}+\alpha_{1}d+\bar{U})$ and define the following distribution function:
+$P(\bar{Y}_{d}\leq c_{1},\bar{U}\leq u|D=d)=\Phi(\min\{(G^{-1}(c_{1})-\alpha_{0}-\alpha_{1}d),u\})$
+In the above proposed DGP, the latent model is not linear. However, the DGP is compatible with the data since,
+$P(\bar{Y}_{d}\leq 0)$ $=$ $P(\tilde{U}\leq-\alpha_{0}-\alpha_{1}d)=\Phi(-\alpha_{0}-\alpha_{1}d)$ (E.5)
+$=$ $1-\Phi(\alpha_{0}+\alpha_{1}d)=P(Y=0|D=d)$
+The last equality is due to the testable equalities holding. Similarly, for any $c_{1}>0$, we can get that
+$P(G(c_{1})<\bar{Y}_{d})$ $=$ $1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)=P(c_{1}<Y|D=d).$ (E.6)
+And,
+$P(G(c_{0})\leq\bar{Y}_{d}\leq G(c_{1}))$ $=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)$ (E.7)
 $=$ $P(c_{0}\leq Y\leq c_{1}|D=d).$
-Hence, when the sharp equalities hold, there exists an alternative model to the classic tobit that can induce the observed data distribution.
+Hence, when the sharp equalities hold, there exists an alternative observationally equivalent model, to the classic tobit that can induce the observed data distribution.
 Indeed, the observed distribution of $Y$ and the non refuted marginal distribution of $U$, do not imply that we can learn the true distribution of $Y_{1},Y_{0},U$.
-A similar phenomena occurs with the iv-tobit mutatis mutandis.
+E.2. Non-Learnability for the IV- Tobit model
+As mentioned in Remark 5, the IV-Tobit model is non-verifiable, that is, we cannot learn if the maintained model is the true data generating process based on the sharp equalities proposed in Section 3.2.
+One can show that by finding an alternative model that is compatible with the equalities (3.4)-(3.12) in all cases in which the IV-Tobit model could not be disregarded.
+Suppose that,
+$\left\{\begin{array}[]{lcl}\tilde{Y}&=&\max(0,\tilde{Y}^{*})\\ \\ \tilde{Y}^{*}&=&G(\beta_{0}+\beta_{1}Z+\tilde{W})\\ \\ \tilde{D}&=&\gamma_{0}+\gamma_{1}Z+\tilde{V}\end{array}\right.$ (E.13)
+Where $G$ is the same function as in the previous subsection. Assume that equations 3.4-3.12 hold:
+$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
+-- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
+-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
+$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
+$P(c_{0}\leq Y\leq c_{1},D\leq d_{0}|Z=z)$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
+-- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
+$P(c_{0}\leq Y\leq c_{1},D\geq d_{1}|Z=z)$ $=$ $\Phi_{W}(c_{1}-\beta_{0}-\beta_{1}z)-\Phi_{W}(c_{0}-\beta_{0}-\beta_{1}z)$
+-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
+$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
+$P(c_{1}\leq Y,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z)-\Phi_{V}(d_{0}-\gamma_{0}-\gamma_{1}z)$
+-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
+$+$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
+$P(c_{1}\leq Y,D\leq d_{0}|Z=z)$ $=$ $\Phi_{V}(d_{0}-\gamma_{0}-\gamma_{1}z)$
+-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
+$P(c_{1}\leq Y,D\geq d_{1}|Z=z)$ $=$ $1-\Phi_{W}(c_{1}-\beta_{0}-\beta_{1}z)-\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z)$
+$+$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
+$P(Y=0,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
+-- $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
+$P(Y=0,D\leq d_{0}|Z=z)$ $=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho),$
+$P(Y=0,D\geq d_{1}|Z=z)$ $=$ $\Phi_{W}(-\beta_{0}-\beta_{1}z)-\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
+Define $\bar{Y}_{z}=G(\beta_{0}+\beta_{1}z+\tilde{W}),\tilde{D}_{z}=\gamma_{0}+\gamma_{1}z+\tilde{V}$ and define the following distribution function:
+$P(\bar{Y}_{d}\leq c_{1},\tilde{D}_{z}\leq d_{1},\bar{U}\leq u,\bar{V}\leq v|D=d)=\Phi(\min\{(G^{-1}(c_{1})-\beta_{0}-\beta_{1}z),u\},\min\{d_{1}-\gamma_{0}-\gamma_{1}z,v\};\rho)$
+Now note that, for example,
+$P(G(c_{0})\leq\tilde{Y}_{z}\leq G(c_{1}),d_{0}\leq\tilde{D}_{z}\leq d_{1})$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$ (E.14)
+-- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
+-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
+$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
+$=$ $P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z)$
+A similar logic holds for the other equalities. But the model is not the IV-tobit model.
+Appendix F Simulations for test’s power under violations of normality
+Turning our attention to the test’s power under violations of normality of the latent error, we consider the modified DGP in Equation F.7.
+In this DGP, $F(\cdot)$ denotes an alternative density for the latent error $U$, which we vary across scenarios.
+Table 6 present the test’s empirical coverage when $F(\cdot)$ is uniform, lognormal and t-student distributions.
+$\left\{\begin{array}[]{lcl}Y&=&\max\{Y^{*},0\}\\ Y^{*}&=&D+U\\ D&=&2Z-V\\ \begin{pmatrix}V\\ Z\end{pmatrix}&\sim&\mathcal{N}(\mathbf{0}_{2},\mathbf{\Sigma})\\ \mathbf{\Sigma}&=&\begin{pmatrix}1&0\\ 0&1\\ \end{pmatrix}\\ U\sim F(\cdot)\end{array}\right.$ (F.7)
+As expected, the rejection rate for the case of the t-student with 80 degrees of freedom is close to the nominal size of the test, since it represents a very mild violation of the normality assumption.
+In the other three cases, which deviate significantly from the normality assumption, the test rejects the null hypothesis in most instances.
+This example indicates that the proposed test is useful in detecting violations of the distributional assumptions in the classic Tobit model.
+Table 6. Test power for violations in error structure
+$N$ $U$ MSE$(\beta_{0})$ MSE$(\beta_{1})$ $\alpha=10\%$ $\alpha=5\%$ $\alpha=1\%$
+5000 t-student (df=80) 0.9987 1.0018 7.21% 4.01% 1.40%
+t-student (df=5) 1.027 1.0624 92.20% 81.80% 48.80%
+LogNormal 1.3336 0.0896 100.00% 100.00% 100.00%
+Uniform 0.892 0.785 100.00% 100.00% 100.00%
+8000 t-student (df=80) 0.9999 1.0034 7.00% 4.80% 1.20%
+t-student (df=5) 1.0251 1.0653 99.20% 97.80% 87.58%
+LogNormal 1.3385 0.0866 100.00% 100.00% 100.00%
+Uniform 0.8932 0.7896 100.00% 100.00% 100.00%
+10000 t-student (df=80) 1.0016 1.0059 5.42% 2.41% 1.00%
+t-student (df=5) 1.0241 1.0642 99.20% 98.40% 90.58%
+LogNormal 1.3348 0.0868 100.00% 100.00% 100.00%
+Uniform 0.8908 0.7862 100.00% 100.00% 100.00%
+Based on $500$ replications.

### v3

@@
 Abstract.
-This paper develops sharp testable implications for Tobit and IV-Tobit models’ identifying assumptions: linear index specification, (joint) normality of latent errors, and treatment (instrument) exogeneity and relevance.
+We develop sharp, testable implications for the identifying assumptions of Tobit and IV-Tobit models: linear index, (joint) normality of errors, treatment (instrument) exogeneity, and relevance.
 The new sharp testable equalities can detect all possible observable violations of the identifying conditions.
-We propose a testing procedure for the model’s validity using existing inference methods for intersection bounds.
-Simulation results suggest adequate test size for large samples and reasonable power in detecting violations of the exogeneity assumption and error structure.
-We review and propose new alternative paths to partially identify the parameters of interest under less restrictive assumptions.
-We revisit a study of married women’s labor supply in Lee (1995) to demonstrate the practical implementation and usefulness of the test.
-The present version is as of January 17, 2025, the first version is from December 2022.
+The proposed test procedure for the model’s validity uses existing inference methods for intersection bounds.
+Simulations suggest adequate test size and power in detecting exogeneity and error structure violations.
+We review and propose alternatives to partially identify the parameters of interest under less restrictive assumptions.
+We revisit a study of married women’s labor supply in lee1995semi to demonstrate the test’s practical implementation.
+The present version is as of December 15, 2025, the first version is from December 2022.
 All errors are ours.
@@
 1. Introduction
-Since the seminal work of Tobin (1958), Tobit models have earned attention in economics, business and social sciences.
+Since the seminal work of tobin_estimation_1958, Tobit models have earned attention in economics, business and social sciences.
 [^1] According to Google Scholar Tobin’s original paper has more than 10000 citations.
-Tobin (1958) analyzed household expenditure on durable goods using a regression model that specifically incorporated that expenditure (the dependent variable) cannot be negative.
+tobin_estimation_1958 analyzed household expenditure on durable goods using a regression model that specifically incorporated that expenditure (the dependent variable) cannot be negative.
 This approach is related to a broader class of censored or truncated regression models, depending on whether observations outside a specified range are lost or censored.
 When applied researchers are interested in modelling limited dependent variables, potentially with mass accumulation points, the Tobit family of models provides structure to identify parameters of interest, such as the average treatment effect (ATE).
-Identification relies on three primary sources: (i) instrument exogeneity (or exogeneity of the variable of interest itself), (ii) normality of the model’s latent variables, and in the case of an instrumental variable approach to endogeneity, (iii) the relevance condition for the instrument.
+Identification relies on three primary sources: (i) instrument exogeneity (or exogeneity of the variable of interest itself), (ii) normality of the model’s latent variables, and, in the case of an instrumental variable approach to endogeneity, (iii) the relevance condition for the instrument.
 While researchers recognize the model’s restrictive nature, it remains a valuable tool in the empirical literature.
 In this paper, we develop a test for the validity of the Tobit model’s structure and assumptions, providing three main contributions to the literature.
-The first is to provide sharp testable equalities that can detect all possible observable violations of the Tobit model.
+The first is to provide the set of sharp testable equalities that can detect all possible observable violations of the Tobit model.
 Second, we propose a test for the validity of the Tobit model’s identifying assumptions using the sharp equalities that characterize the model to check its falsifiability.
-Following recent literature, we convert the equalities into conditional moment inequalities and implement the test by existing inferential methods from Chernozhukov, Lee, and Rosen (2013).
-The Tobit family of models involves continuous outcomes with accumulation points and potentially continuous treatment, implying a large number of moment equalities to be tested.
+Following recent literature, we convert the equalities into conditional moment inequalities and implement the test by existing inferential methods from chernozhukov_intersection_2013.
+The Tobit model family is used for continuous outcomes with accumulation points.
+In the case of household expenditure, the outcome may exhibit a zero accumulation due to censoring.
+When evaluating causality, researchers may be interested in a treatment variable, not necessarily binary, which would result in a large number of moment equalities to be tested.
 This creates significant challenges to test implementation.
@@
 Most of the preceding work focuses on testing a particular assumption or feature of the model while maintaining other assumptions and structures as valid.
-Nelson (1981) constructs a Hausman-type test for misspecification of the classic Tobit model (that is, normality, linear index, and homoskedasticity) where the maximum likelihood estimates are compared with method of moment estimates.
+nelson_test_1981 constructs a Hausman-type test for misspecification of the classic Tobit model (that is, normality, linear index, and homoskedasticity) where the maximum likelihood estimates are compared with method of moment estimates.
 Nelson’s test compares the sample proportion of non-censored observations with the hypothesized probability of being non-censored in the Tobit model.
-Bera, Jarque, and Lee (1984) state that the test is equivalent to the Lagrange multiplier (LM) test of the Tobit model against Cragg’s model (Lin and Schmidt, 1984) and propose an alternative LM test for the normality assumption against other distributions of the Pearson family of distributions while the remaining assumptions are maintained.
-Newey (1987b) considers both exogenous and endogenous explanatory variables cases using symmetrically censored least squares estimators to construct specification tests of normality and homoskedasticity assumptions via a Hausman-type specification test.
-Holden (2004) examines several statistics proposed to test the normality assumption in the Tobit (censored regression) model and reinterpreted them as a version of the LM (score) test for a common null hypothesis.
-Reynolds and Shonkwiler (1991) use an information matrix misspecification test to detect violations of the distributional assumptions of the Tobit model.
-Other tests include Drukker (2002), which operationalized conditional moment tests developed by Newey (1985) and Tauchen (1985) to the case of misspecification of the distribution of the classic Tobit model.
+bera_testing_1984 state that the test is equivalent to the Lagrange multiplier (LM) test of the Tobit model against Cragg’s model (lin_test_1984) and propose an alternative LM test for the normality assumption against other distributions of the Pearson family of distributions while the remaining assumptions are maintained.
+newey_specification_1987 considers both exogenous and endogenous explanatory variables cases using symmetrically censored least squares estimators to construct specification tests of normality and homoskedasticity assumptions via a Hausman-type specification test.
+holden_testing_2004 examines several statistics proposed to test the normality assumption in the Tobit (censored regression) model and reinterprets them as a version of the LM (score) test for a common null hypothesis.
+reynolds_testing_1991 use an information matrix misspecification test to detect violations of the distributional assumptions of the Tobit model.
+Other tests include drukker_bootstrapping_2002, which operationalized conditional moment tests developed by newey_maximum_1985 and tauchen_diagnostic_1985 to the case of misspecification of the distribution of the classic Tobit model.
 With a similar intuition to our framework, their test writes down conditional moment restrictions, which should have zero conditional expected values under the null.
 Since the model was estimated by maximum likelihood, the assumed data-generating process specifies the moments of disturbances conditional on the covariates to be the ones of a normal distribution.
-Drukker (2002) use these moment-based methods based on the third and fourth moments of the normal distribution.
+drukker_bootstrapping_2002 use these moment-based methods based on the third and fourth moments of the normal distribution.
 While intuitively similar, our procedure detects all possible violations of the model, not only those evident from deviations in the third and fourth moments.
-Smith and Blundell (1986) propose a test of the treatment variable’s exogeneity in the IV-Tobit Model by a control function approach exploiting the joint normality of the latent variables.
-Most of these approaches focus on testing a particular subset of assumptions or a specific class of alternatives to consider.
+smith_exogeneity_1986 propose a test of the treatment variable’s exogeneity in the IV-Tobit Model by a control function approach exploiting the joint normality of the latent variables.
+Most of these approaches focus on testing a particular subset of assumptions or consider a specific class of alternatives.
 The developments proposed in this paper consider all observable violations of the general Tobit model structure and its assumptions, serving as a useful test for empirical researchers constructing models for censored/truncated data.
 This work contributes to the growing literature on the testability of the identifying assumptions in various econometric models.
-Those include tests for the validity of assumptions in instrumental variable models, local average treatment effects, regression discontinuity designs, bivariate probit, encouragement designs, among others.
-(Mourifié and Wan, 2017; Kitagawa, 2015; Huber and Mellace, 2015; Gunsilius, 2021; Arai et al., 2022; Acerenza, Bartalotti, and Kédagni, 2023; Bai and Tabord-Meehan, 2024).
-Our testing procedures are connected to Kédagni and Mourifié (2020), which provides tests for assumptions related to the instrumental variables in the model, while we also consider the implications of the parametric structure on the outcome and latent error structure, which are inherent to Tobit models.
-The two papers closest to ours are Acerenza, Bartalotti, and Kédagni (2023) and Goff, Kédagni, and Wu (2024).
+Those include sharp tests of the validity in instrumental variable models222gunsilius_nontestability_2021 proves that instrument validity cannot be tested in the case where the endogenous treatment is continuously distributed.
+and local average effects (mourifie_testing_2017; kitagawa_test_2015; huber_testing_2015; kedagni2020generalized), sharp tests of the validity of assumption in the regression discontinuity designs (arai_testing_2022), sharp tests of the assumptions of the bivariate probit (acerenza_testing_2023) and sharp tests in the context of encouragement designs (bai2024sharptestableimplicationsencouragement) among others.
+Our testing procedures are connected to kedagni2020generalized, which provides tests for assumptions related to the instrumental variables in the model, while we also consider the implications of the parametric structure on the outcome and latent error structure, which are inherent to Tobit models.
+The two papers closest to ours are acerenza_testing_2023 and goff2024testing.
 The first focuses on bivariate probit models, leading to a finite set of moment equalities as testable restrictions.
 As mentioned above, this study builds upon their work by tackling the case of the Tobit family of estimators, which provides additional technical and practical challenges given the continuous nature of outcome and treatment, including accumulation points.
-Goff, Kédagni, and Wu (2024) considers separable parametric instrumental variable models that induce a different set of moment equalities to be tested while our approach focuses on Tobit and IV-Tobit models, which can potentially be non-separable.
-[^2] While Goff, Kédagni, and Wu (2024) briefly mention the possibility of extending their results to non-separable models, they obtain different equalities from ours and do not discuss identification or sharpness.
+goff2024testing considers separable parametric instrumental variable models that induce a different set of sharp moment equalities to be tested while our approach focuses on Tobit and IV-Tobit models, which can potentially be non-separable.
+[^3] While goff2024testing briefly mentions the possibility of extending their results to non-separable models, they obtain different equalities from ours and do not discuss identification or sharpness in that context.
 Implementation would also differ from our approach.
@@
 Section 2 presents the model and identifying assumptions.
+Section 3 discusses heuristically the identification of the model’s parameters.
 Section 4 derives the sharp testable implications.
-Section 4 discusses heuristically the identification of the model’s parameters.
 Section 5 outlines the testing procedure.
-Section 6 include simulation evidence about the test’s size while discussions about power are relegated to the Appendix.
+Section 6 include simulation evidence about the test’s size, while discussions about power are relegated to the Appendix.
 Section 7 discusses how to relax the assumptions in case of rejection.
-Section 8 provides an empirical illustration revisiting the study of married women’s labor supply from Lee (1995) to demonstrate the practical implementation and usefulness of the test.
+Section 8 provides an empirical illustration revisiting the study of married women’s labor supply from lee1995semi to demonstrate the practical implementation and usefulness of the test.
 Finally, Section 9 concludes.
-Additional results and details are collected in Appendices A-F.
+Additional results and details are collected in Appendices A-E.
 2. Models
@@
 Both the treatment of interest, $D$, and the instrumental variable, $Z$, can be discrete or continuous and take values in $\mathcal{D}\subset\mathbb{R}$ and $\mathcal{Z}\subset\mathbb{R}$, respectively.
-We normalize the coefficients by the standard deviation of the convenient unobservable structure for the exposition.
+We consider the normalized coefficients by the standard deviation of the unobservable error term, which is convenient for the exposition.
 [^5] Identification for the normalized and non-normalized parameters is discussed in section 3.
@@
 The classic Tobit model considers the case in which the researcher is interested in the effect of an exogenous treatment on a non-negative outcome that has a mass point at zero:
-$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ \\ Y^{*}&=&\alpha_{0}+\alpha_{1}D+U\end{array}\right.$ (2.4)
+$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ \\ Y^{*}&=&\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+U\end{array}\right$. (2.4)
 where $U$ is an unobservable (latent) error. In addition to the model structure in system 2.4, the classic Tobit restricts the distribution of the error term and its relationship to the treatment.
@@
 Assumption 2.
-$U$ is distributed as $N(0,1)$.
+$U$ is distributed as $N(0,\sigma^{2})$.
 Assumption 1 states that the treatment of interest is independent of the model’s unobservables.
-Assumption 2 imposes that the latent error has a standard normal distribution.
-[^5] The error could take $N(0,\sigma^{2})$ but after scale normalizations we can model it as $N(0,1)$.
-Despite its simplicity, the classic Tobit model can accommodate continuous or discrete treatment, and the normality assumption for the latent error could be replaced by other distributional assumptions that preserve the identification of the parameters.
+Assumption 2 imposes that the latent error has a normal distribution.
+To simplify some of the derivations later in the manuscript, we defined the scaled version the model for which the transformed error term would have distribution $N(0,1)$.
+$\frac{Y^{*}}{\sigma}=\frac{\tilde{\alpha}_{0}}{\sigma}+\frac{\tilde{\alpha}_{1}}{{\sigma}}D+\frac{U}{\sigma}=\alpha_{0}+\alpha_{1}D+\frac{U}{\sigma}$
+We refer to the standardized parameters without the tilde, thus for example $\alpha_{1}=\frac{\tilde{\alpha}_{1}}{\sigma}$.
+Despite its simplicity, the classic Tobit model can accommodate continuous or discrete treatments, and the normality assumption for the latent error can be replaced with other distributional assumptions that preserve parameter identification.
 2.2. IV Tobit model
 The IV Tobit considers the case in which the treatment of interest is endogenous, but an instrumental variable is available to identify its effect.
+It can be traced back to heckman1977dummy; amemiya1979estimation; nelson1978specification among several approaches to consider censored variables with endogeneity.
+[^6] In Section 7, we discuss alternative models and how identification can still be achieved under different assumptions, including about the distribution of the latent error terms.
 Let the IV Tobit model be: [^7] Results for the IV-Tobit model are presented for continuous $D$ but hold for a binary treatment, $D=1\{\gamma_{0}+\gamma_{1}Z-V>0\}$.
 In that case, the parameters are identified only up to scale.
-$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\alpha_{0}+\alpha_{1}D+U\\ D&=&\gamma_{0}+\gamma_{1}Z+V\end{array}\right.$ (2.8)
-Alternatively, in its reduced form representation:
-$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\beta_{0}+\beta_{1}Z+W\\ D&=&\gamma_{0}+\gamma_{1}Z+V\end{array}\right.$ (2.12)
-Where $U,V$ are the latent structural error terms.
+$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+U\\ D&=&\tilde{\gamma}_{0}+\tilde{\gamma}_{1}Z+\tilde{V}\end{array}\right$. (2.8)
+Where $U,\tilde{V}$ are the latent structural error terms. Alternatively, in its reduced form representation:
+$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\tilde{\beta}_{0}+\tilde{\beta}_{1}Z+\tilde{W}\\ D&=&\tilde{\gamma}_{0}+\tilde{\gamma}_{1}Z+\tilde{V}\end{array}\right$. (2.12)
+Where $\tilde{\beta}_{0}=\tilde{\alpha}_{0}+\tilde{\alpha}_{1}\tilde{\gamma}_{0},\tilde{\beta}_{1}=\tilde{\alpha}_{1}\tilde{\gamma}_{1}$ and $\tilde{W},\tilde{V}$ are the reduce form error terms where $\tilde{W}=\tilde{\alpha}_{1}\tilde{V}+U$.
 To that model structure, restrictions on the joint distribution of the latent variables and their relationship to the instrumental variable are added.
 Assumption 3.
-$Z$ is independent of $(U,V)$ and $\gamma_{1}\neq 0$.
+[…]
 Assumption 4.
-Let $U,V$ follow a bivariate normal distribution with covariance $\rho$, i.e., $\begin{pmatrix}U\\ V\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma=\begin{pmatrix}\sigma^{2}_{U}&\rho_{UV}\\ \rho_{UV}&\sigma^{2}_{V}\end{pmatrix}$.
+Let $U,\tilde{V}$ follow a bivariate normal distribution with covariance $\rho_{U\tilde{V}}$, i.e., $\begin{pmatrix}U\\ \tilde{V}\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma=\begin{pmatrix}\sigma^{2}_{U}&\rho_{U\tilde{V}}\\ \rho_{U\tilde{V}}&\sigma^{2}_{\tilde{V}}\end{pmatrix}$.
 Assumption 3(i) states that the instrumental variable is independent of the model’s structural latent variables.
 The second part of Assumption 3 is the usual instrument relevance in determining the treatment.
-This assumption also implies that $Z$ is independent of $W,V$, since $W=\alpha_{1}V+U$.
-Assumption 4 characterizes the distribution of the latent vector of error terms as bivariate normal.
-Then, $W,V$ also follows a bivariate normal distribution and $\alpha_{0},\alpha_{1},\gamma_{0},\gamma_{1}$ could be scale normalized so $(W,V)^{\prime}$ follows the standard bivariate normal distribution with covariance $\rho$, i.e., $\begin{pmatrix}W\\ V\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\end{pmatrix}$, and $\Sigma=\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}$.
-[^7] Note that if $U,V$ follow a bivariate normal distribution, $\begin{pmatrix}U\\ V\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma=\begin{pmatrix}\sigma^{2}_{U}&\rho_{UV}\\ \rho_{UV}&\sigma^{2}_{V}\end{pmatrix}$, then $W,V$ follow a bivariate normal distribution, $\begin{pmatrix}W\\ V\end{pmatrix}\sim\mathcal{N}(\mu_{1},\Sigma_{1})$, where $\mu_{1}=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma_{1}=\begin{pmatrix}\alpha_{1}^{2}\sigma^{2}_{V}+\sigma^{2}_{U}+2\rho_{UV}\alpha_{1}\sigma_{V}\sigma_{U}&\alpha_{1}\sigma_{V}^{2}+\rho_{UV}\\ \alpha_{1}\sigma_{V}^{2}+\rho_{UV}&\sigma^{2}_{V}\end{pmatrix}$.
-Then the reduced form system can be scale normalized.
-The latter characterization aids the development of the sharp testable implications.
+This assumption also implies that $Z$ is independent of $\tilde{W},\tilde{V}$, since $\tilde{W}=\tilde{\alpha}_{1}\tilde{V}+U$.
+[…]
+3. Identification
+In this section, we provide heuristic arguments for the identification of the parameters of the respective models, expanding on han_identification_2017 and acerenza_testing_2023.
+These are known in the literature and discussed here for completeness and intuition.
+The testable implications that are our main result are presented in Section 4.
+3.1. Classic Tobit model
+Note that,
+$1-P(Y=0|D)=1-P(Y^{*}\leq 0|D)=\Phi(\alpha_{0}+\alpha_{1}D).$
+The equality follows from the model’s structure in Equation 2.4 and assumptions 1 and 2.
+Since the standard normal CDF, $\Phi(.)$, is monotonic, we use its inverse to obtain:
+$\Phi^{-1}(1-P(Y=0|D))=\alpha_{0}+\alpha_{1}D.$
+Thus, $\alpha_{1}=\frac{Cov(\Phi^{-1}(1-P(Y=0|D)),D)}{Var(D)}$, and $\alpha_{0}=E(\Phi^{-1}(1-P(Y=0|D)))-\alpha_{1}E(D)$.
+The coefficients have been normalized by dividing by the latent variable’s ($U$) square root of the variance, denoted by $\sigma$.
+This implies that the ratio of the original parameters is identified since $\frac{\alpha_{1}}{\alpha_{0}}=\frac{\frac{\tilde{\alpha}_{1}}{\sigma}}{\frac{\tilde{\alpha}_{0}}{\sigma}}=\frac{\tilde{\alpha}_{1}}{\tilde{\alpha}_{0}}$.
+Recall that,
+$F_{U|U<c}(u|U<c)=P(U\leq u|U<c)$ $=$ $\frac{P(U\leq\min(u,c))}{P(U<c)}=\begin{cases}\frac{P(U\leq c)}{P(U\leq c)}=1,\text{ for }u\geq c\\ \frac{P(U\leq u)}{P(U\leq c)}=\frac{\Phi(u)}{\Phi(c)},\text{ for }u<c\end{cases}$
+Then, $f_{U|U<c}(u|U<c)=\begin{cases}0,\text{ for }u\geq c\\ \frac{\phi(u)}{\Phi(c)},\text{ for }u<c\end{cases}$, and $f_{U|U>c}(u|U>c)=\begin{cases}0,\text{ for }u\leq c\\ \frac{\phi(u)}{1-\Phi(c)},\text{ for }u>c\end{cases}$.
+Note then that,
+$E(Y|D,Y>0)$ $=$ $E(\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+U|D,Y>0)$
+$=$ $\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+E(U|D,Y>0)$
+$=$ $\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+E(U|D,\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+U>0)$
+$=$ $\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+E(U|U>-\tilde{\alpha}_{0}-\tilde{\alpha}_{1}D)$
+$=$ $\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+\sigma\frac{\phi(-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}D)}{1-\Phi(-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}D)}$
+where the fourth equality uses independence, and the fifth equality uses the properties of standard truncated normal distributions.
+Since $\alpha_{0},\alpha_{1}$ are identified, then the inverse Mills ratio is identified,
+$\frac{\phi(-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}D)}{1-\Phi(-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}D)}$ $=$ $\lambda(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D)=\lambda(\alpha_{0}+\alpha_{1}D)$
+Thus,
+$E(Y|D,Y>0)=\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+\sigma\lambda(\alpha_{0}+\alpha_{1}D)$
+Which in turn implies (by the property that we can express $X=E(X|W)+e$, where $E(e|W)=0$):
+$Y=\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+\sigma\lambda(\alpha_{0}+\alpha_{1}D)+e$
+with $e$ mean independent of $D,\lambda(\alpha_{0}+\alpha_{1}D)$.
+Since $D$ is known and $\lambda(\cdot)$ is identified, we get a three-by-three linear system with a unique solution for $\tilde{\alpha}_{0},\tilde{\alpha}_{1},\sigma$.
+In other words, the normality assumption allows identification by leveraging the relationship of the variable of interest ($D$) and the outcome at the accumulation point ($Y=0$) to obtain the scaled parameters, and the relationship of treatment with the distribution of $Y$ (when $Y>0$) to overcome the normalization.
+However, the reliance on the normality and linearity assumptions underlines the importance of adequately testing the model’s structure and assumptions in empirical applications.
+3.2. IV Tobit model
+Turning the focus to the IV-Tobit model, the first stage is readily identified from the linear structure, $\tilde{\gamma}_{1}=\frac{Cov(Z,D)}{Var(Z)}$, $\tilde{\gamma}_{0}=E(D)-\tilde{\gamma}_{1}E(Z)$.
+Furthermore, $\tilde{V}=D-E(D)+\frac{Cov(Z,D)}{Var(Z)}E(Z)-\frac{Cov(Z,D)}{Var(Z)}Z$, thus $\tilde{V}$ and $\sigma_{\tilde{V}}^{2}$ are identified.
+Combining these results implies that we can identify $\gamma_{0}=\frac{\tilde{\gamma}_{0}}{\sigma_{\tilde{V}}}$ and $\gamma_{1}=\frac{\tilde{\gamma}_{1}}{\sigma_{\tilde{V}}}$.
+Now note that calculating $P(Y=0|Z)$ from the reduce form:
+$1-P(Y=0|Z)=\Phi(-\beta_{0}-\beta_{1}Z)$
+Where $\beta_{0}=\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}$ and $\beta_{1}=\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}$.
+[…]
+To identify $\rho$, let $s=c-\beta_{0}-\beta_{1}z$ and $t=d-\gamma_{0}-\gamma_{1}z$. Then, by using
+$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}\left(\frac{c_{1}}{\sigma_{\tilde{W}}}-\beta_{0}-\beta_{1}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\gamma_{0}-\gamma_{1}z;\rho\right)$
+- $\Phi_{W,V}\left(\frac{c_{0}}{\sigma_{\tilde{W}}}-\beta_{0}-\beta_{1}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\gamma_{0}-\gamma_{1}z;\rho\right)$
+- $\Phi_{W,V}\left(\frac{c_{1}}{\sigma_{\tilde{W}}}-\beta_{0}-\beta_{1}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\gamma_{0}-\gamma_{1}z;\rho\right)$
+$+$ $\Phi_{W,V}\left(\frac{c_{0}}{\sigma_{\tilde{W}}}-\beta_{0}-\beta_{1}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\gamma_{0}-\gamma_{1}z;\rho\right)$
+$\equiv$ $\Phi^{1}_{c_{1},c_{0},d_{1},d_{0}},$
+We can obtain
+$P(Y\leq c,D\leq d|Z=z)$ $=$ $\Phi_{W,V}(s,t;\rho),$
+which by classic results of bivariate normal random variables if we differentiate w.r.t. $\rho$ is:
+$\frac{\partial P(Y\leq c,D\leq d|Z=z)}{\partial\rho}$ $=$ $\phi_{W,V}(s,t;\rho)$
+Where $\phi_{a,b}$ is the bivariate normal probability density, which is positive for any $s,t$. Hence, $\Phi_{W,V}(.,.;\rho)$ is monotonic in $\rho$ and thus invertible. So $\rho$ can be identified.
+Recall the previous coefficients are normalized, as noted by (wooldridge_econometric_2010, p.
+683-84) we can identify the original and reduced form covariance matrices, as well as the true coefficients without the normalization.
+To explain this in detail, recall that we have identified $\rho$, the covariance between $W,V$ the parameters $\gamma_{1},\gamma_{0},\beta_{1},\beta_{0}$ and the variance of $\tilde{V}$, $\sigma^{2}_{\tilde{V}}$.
+Similarly to the classic tobit model, we can leverage normality of $W$, independence of $Z$, and exploit the reduced form representation of $Y$ to obtain:
+$E(Y|Z,Y>0)$ $=$ $\tilde{\beta}_{0}+\tilde{\beta}_{1}Z+\sigma_{\tilde{W}}\lambda(\beta_{0}+\beta_{1}Z)$
+Which in turn implies
+$Y=\tilde{\beta}_{0}+\tilde{\beta}_{1}Z+\sigma_{\tilde{W}}\lambda(\beta_{0}+\beta_{1}Z)+e$
+with $e$ mean independent of $(Z,\lambda(\beta_{0}+\beta_{1}Z))$.
+Since $Z$ is known and $\lambda(\beta_{0}+\beta_{1}z)$ is identified for any given $Z=z$, we can get a solution for $\sigma_{\tilde{W}}$ as one of the coefficients for the linear projection of $Y$ on $(Z,\lambda(\beta_{0}+\beta_{1}Z))$, identifying the non-standardized reduced form error for $Y$.
+Finally, since $\beta_{0}=\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}=\frac{\tilde{\alpha}_{0}+\tilde{\alpha}_{1}\tilde{\gamma}_{0}}{\sigma_{\tilde{W}}}$, $\beta_{1}=\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}=\frac{\tilde{\alpha}_{1}\tilde{\gamma}_{0}}{\sigma_{\tilde{W}}}$ and we have identified $\sigma_{\tilde{V}},\gamma_{0},\gamma_{1},\tilde{\gamma}_{0},\tilde{\gamma}_{1}$ and $\sigma_{\tilde{W}}$ we can recover $\tilde{\alpha}_{0},\tilde{\alpha}_{1}$.
+It now only remains to recover the original variance-covariance structure.
+Under bivariate normality of the latent variables in the structural model, we can express the following relationships between $U,\tilde{V}$ with $\tilde{W}=\tilde{\alpha}_{1}U+\tilde{V},V$ and $W=\frac{\tilde{\alpha}_{1}U+\tilde{V}}{\sigma_{\tilde{W}}},V=\frac{\tilde{V}}{\sigma_{\tilde{V}}}$
+$\begin{pmatrix}U\\ \tilde{V}\end{pmatrix}\sim\mathcal{N}\Bigg(\begin{pmatrix}0\\ 0\ \end{pmatrix},\begin{pmatrix}\sigma^{2}_{U}&\rho_{U\tilde{V}}\\ \rho_{U\tilde{V}}&\sigma^{2}_{\tilde{V}}\end{pmatrix}\Bigg)$
+$\begin{pmatrix}\tilde{W}\\ \tilde{V}\end{pmatrix}\sim\mathcal{N}\Bigg(\begin{pmatrix}0\\ 0\ \end{pmatrix},\begin{pmatrix}\tilde{\alpha}_{1}^{2}\sigma^{2}_{\tilde{V}}+\sigma^{2}_{U}+2\rho_{U\tilde{V}}\tilde{\alpha}_{1}\sigma_{\tilde{V}}\sigma_{U}&\tilde{\alpha}_{1}\sigma_{\tilde{V}}^{2}+\rho_{U\tilde{V}}\\ \tilde{\alpha}_{1}\sigma_{\tilde{V}}^{2}+\rho_{U\tilde{V}}&\sigma^{2}_{\tilde{V}}\end{pmatrix}\Bigg)$
+$\begin{pmatrix}W\\ V\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\ \end{pmatrix},\begin{pmatrix}1&\frac{\tilde{\alpha}_{1}\sigma_{\tilde{V}}^{2}+\rho_{U\tilde{V}}}{\big(\tilde{\alpha}_{1}^{2}\sigma^{2}_{\tilde{V}}+\sigma^{2}_{U}+2\rho_{U\tilde{V}}\tilde{\alpha}_{1}\sigma_{\tilde{V}}\sigma_{U}\big)\sigma_{\tilde{V}}}\\ \frac{\tilde{\alpha}_{1}\sigma_{\tilde{V}}^{2}+\rho_{U\tilde{V}}}{\big(\tilde{\alpha}_{1}^{2}\sigma^{2}_{\tilde{V}}+\sigma^{2}_{U}+2\rho_{U\tilde{V}}\tilde{\alpha}_{1}\sigma_{\tilde{V}}\sigma_{U}\big)\sigma_{\tilde{V}}}&1\end{pmatrix}\right)$
+Since we have identified $\tilde{\alpha}_{1},\sigma_{\tilde{V}},\sigma_{\tilde{W}},\rho$ from the following equations:
+$\rho=\frac{\tilde{\alpha}_{1}\sigma_{\tilde{V}}^{2}+\rho_{U\tilde{V}}}{\big(\tilde{\alpha}_{1}^{2}\sigma^{2}_{\tilde{V}}+\sigma^{2}_{U}+2\rho_{U\tilde{V}}\tilde{\alpha}_{1}\sigma_{\tilde{V}}\sigma_{U}\big)\sigma_{\tilde{V}}}$
+$\sigma^{2}_{\tilde{W}}=\tilde{\alpha}_{1}^{2}\sigma^{2}_{\tilde{V}}+\sigma^{2}_{U}+2\rho_{U\tilde{V}}\tilde{\alpha}_{1}\sigma_{\tilde{V}}\sigma_{U}$
+We can identify $\sigma_{U}$ and $\rho_{U\tilde{V}}$, and thus, recover all the structural parameters.
 4. Sharp testable equalities
 This section presents the sharp testable equalities for the models described in Section 2.
-Deviations from these equalities imply violations of the “null hypothesis” that the Tobit model (linear latent index, independence, instrument validity, and normality) is valid.
+Deviations from these equalities imply violations of the “null hypothesis” that the Tobit model (linear latent index, treatment/instrument independence, instrument relevance, and normality) is valid.
 The equalities are conditional on parameters being identified under the model’s assumptions, and a discussion on identification is postponed to Section 3.
@@
 $=$ $P(c_{0}\leq Y^{*}\leq c_{1}|D=d)$
-$=$ $P(c_{0}-\alpha_{0}-\alpha_{1}d\leq U\leq c_{1}-\alpha_{0}-\alpha_{1}d|D=d)$
-$=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)$
+$=$ $P(c_{0}-\tilde{\alpha}_{0}-\tilde{\alpha}_{1}d\leq U\leq c_{1}-\tilde{\alpha}_{0}-\tilde{\alpha}_{1}d|D=d)$
+$=$ $\Phi\left(\frac{c_{1}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}d\right)-\Phi\left(\frac{c_{0}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}d\right)$
 where $\Phi(.)$ is the standard normal CDF.
@@
 The fifth one is given by the latent linear model structure of $Y^{*}$, and finally, the last step follows from assumptions 1 and 2.
-To cover all possible cases of positive outcomes, note that for any $c_{1}>0$:
-$P(c_{1}\leq Y|D=d)$ $=$ $1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d).$ (3.2)
-Turning to the accumulation point, the observed event of $Y=0$ has probability,
-$P(Y=0|D=d)=P(Y^{*}\leq 0|D=d)=\Phi(-\alpha_{0}-\alpha_{1}d).$ (3.3)
-The first equality follows from the model’s structure in Equation 2.4, and the second one applies assumptions 1 and 2.
-The equalities described in equations 3.1-3.3 fully characterize the distribution of $Y$ conditional on $D$, connecting the probabilities in the observed data to those implied by the Tobit model.
-Remark 1 (Sharpness).
-In the context of model 2.4, equalities 3.1-3.3 are sharp in the sense that whenever they hold, it is possible to construct a vector of $(\tilde{Y},D,\tilde{U})$ that satisfies model 2.4, Assumptions 1 and 2, and induces the observed distribution on the data $(Y,D)$.
-For the proof, see Appendix A.
-Then, the equalities 3.1-3.3 are sharp testable implications for the validity of the classic Tobit Model in 2.4 coupled with assumptions 1-2.
+Recalling the accumulation point, the observed event of $Y=0$ has a probability,
+$P(Y=0|D=d)=P(Y^{*}\leq 0|D=d)=\Phi\left(-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}d\right).$ (4.2)
+The equalities described in equations 4.1-4.2 fully characterize the distribution of $Y$ conditional on $D$, connecting the probabilities in the observed data to those implied by the Tobit model.
+We collect these results more formally in Theorem 1.
+Theorem 1.
+[…]
+The proof for the equalities follows from the discussion above, and further details about sharpness are presented in Appendix A.
+Then, the equalities (4.1)-(4.2) are sharp testable implications for the validity of the classic Tobit Model in 2.4 coupled with assumptions 1-2.
 They serve as the basis for the test procedure described in Section 5.
-Remark 2 (Non-learnability).
+Proposition 1 (Non-learnability).
 The testable implications and sharpness discussed above show that the classic Tobit model is generally refutable.
-However, the model is non-verifiable in the sense that we can always construct a joint probability law of $(\bar{Y},D,\bar{U})$ that violates the Tobit model validity but satisfies equalities 3.1-3.3.
+However, the model is non-verifiable in the sense that we can always construct a joint probability law of $(\bar{Y},D,\bar{U})$ that violates the Tobit model validity but satisfies equalities (4.1)-(4.2).
 See Appendix E for the proof.
-Remark 3 (Extensions).
-The previous derivation can be adjusted for different variations of Tobit models, such as generalizations of the distributional assumptions (Barros et al., 2018), different thresholds (Carson and Sun, 2007), dynamic Tobit models, or including individual specific effects (Wooldridge, 2005; Honore, Kyriazidou, and Powell, 2000; Honoré, 1993).
+Proposition 2 (Extensions).
+The previous derivation can be adjusted for different variations of Tobit models, such as generalizations of the distributional assumptions (barros2018generalized), different thresholds (carson2007Tobit), dynamic Tobit models, or including individual-specific effects (wooldridge2005simple; honore2000estimation; honore1993orthogonality).
+In particular, note that similar approaches to the ones proposed in Theorem 1 can be used even if the latent errors don’t follow a normal distribution and $Y^{*}$ does not have a linear index form, as long as the model is identified.
 In appendix C, we derive the equalities for the type 2 Tobit model, and similar logic can be applied to other two-part models.
@@
 4.2. IV Tobit model
-We turn our attention to the instrumental variable Tobit case and propose testable implications that can be used to test the model described in 2.8 and the associated assumptions 3 and 4.
+We turn our attention to the IV Tobit case and propose testable implications that can be used to test the model described in 2.8 and the associated assumptions 3 and 4.
 The observed data includes $(Y,D,Z)$, and we characterize the joint distribution of $(Y,D)\in\mathcal{Y}\times\mathcal{D}$ conditional on the instrument, $Z$, to obtain the model’s testable implications.
@@
 Consider any $0<c_{0}<c_{1},d_{0}<d_{1}$, and $z\in\mathcal{Y}\times\mathcal{D}\times\mathcal{Z}$,
-$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.4)
--- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
--- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
-$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
+$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$ (4.3)
+- $\Phi_{W,V}(\frac{c_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$
+- $\Phi_{W,V}(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$
+$+$ $\Phi_{W,V}(\frac{c_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$
 $\equiv$ $\Phi^{1}_{c_{1},c_{0},d_{1},d_{0}},$
-$P(c_{0}\leq Y\leq c_{1},D\leq d_{0}|Z=z)$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.5)
--- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
+$P(c_{0}\leq Y\leq c_{1},D\leq d_{0}|Z=z)$ $=$ $\Phi_{W,V}\left(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho\right)$ (4.4)
+- $\Phi_{W,V}\left(\frac{c_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho\right)$
 $\equiv$ $\Phi^{1}_{c_{1},c_{0},d_{0}},$
-$P(c_{0}\leq Y\leq c_{1},D\geq d_{1}|Z=z)$ $=$ $\Phi_{W}(c_{1}-\beta_{0}-\beta_{1}z)-\Phi_{W}(c_{0}-\beta_{0}-\beta_{1}z)$ (3.6)
--- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
-$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
+$P(c_{0}\leq Y\leq c_{1},D\geq d_{1}|Z=z)$ $=$ $\Phi_{W}(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z)-\Phi_{W}(\frac{c_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z)$ (4.5)
+- $\Phi_{W,V}(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$
+$+$ $\Phi_{W,V}(\frac{c_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$
 $\equiv$ $\Phi^{1}_{c_{1},c_{0},d_{1}},$
-$P(c_{1}\leq Y,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z)-\Phi_{V}(d_{0}-\gamma_{0}-\gamma_{1}z)$ (3.7)
--- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
-$+$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
+$P(c_{1}\leq Y,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{V}(\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z)-\Phi_{V}(\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z)$ (4.6)
+- $\Phi_{W,V}(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$
+$+$ $\Phi_{W,V}(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$
 $\equiv$ $\Phi^{1}_{c_{1},d_{1},d_{0}},$
-$P(c_{1}\leq Y,D\leq d_{0}|Z=z)$ $=$ $\Phi_{V}(d_{0}-\gamma_{0}-\gamma_{1}z)$ (3.8)
--- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
+$P(c_{1}\leq Y,D\leq d_{0}|Z=z)$ $=$ $\Phi_{V}(\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z)$ (4.7)
+- $\Phi_{W,V}(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$
 $\equiv$ $\Phi^{1}_{c_{1},d_{0}},$
-$P(c_{1}\leq Y,D\geq d_{1}|Z=z)$ $=$ $1-\Phi_{W}(c_{1}-\beta_{0}-\beta_{1}z)-\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z)$ (3.9)
-$+$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
+$P(c_{1}\leq Y,D\geq d_{1}|Z=z)$ $=$ $1-\Phi_{W}(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z)-\Phi_{V}(\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z)$ (4.8)
+$+$ $\Phi_{W,V}(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$
 $\equiv$ $\Phi^{1}_{c_{1},d_{1}}.$
 where $\Phi_{W,V}(w,v;\rho)$ denotes the joint c.d.f. of $(W,V)$, a standard bivariate normal with coefficient of correlation $\rho$. Similarly, considering the case that the observed $Y$ equals zero,
-$P(Y=0,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.10)
--- $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)\equiv\Phi^{2}_{d_{1},d_{0}},$
-$P(Y=0,D\leq d_{0}|Z=z)$ $=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)\equiv\Phi^{2}_{d_{0}},$ (3.11)
-$P(Y=0,D\geq d_{1}|Z=z)$ $=$ $\Phi_{W}(-\beta_{0}-\beta_{1}z)-\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.12)
+$P(Y=0,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}(-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$ (4.9)
+- $\Phi_{W,V}(-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)\equiv\Phi^{2}_{d_{1},d_{0}},$
+$P(Y=0,D\leq d_{0}|Z=z)$ $=$ $\Phi_{W,V}(-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)\equiv\Phi^{2}_{d_{0}},$ (4.10)
+$P(Y=0,D\geq d_{1}|Z=z)$ $=$ $\Phi_{W}(-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z)-\Phi_{W,V}(-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$ (4.11)
 $\equiv$ $\Phi^{2}_{d_{1}}.$
-Remark 4 (Sharpness).
-In the context of model 2.8 or 2.12, equalities 3.4-3.12 for all $c,d,z\in\mathcal{Y}\times\mathcal{D}\times\mathcal{Z}$ are sharp in the sense that whenever they hold, it is possible to construct a vector of $(\tilde{Y},\tilde{D},\tilde{U},\tilde{V},Z)$ or equivalently $(\tilde{Y},\tilde{D},\tilde{W},\tilde{V},Z)$ that satisfies model 2.8 and 2.12, Assumptions 3 and 4, and induces the observed distribution on the data $(Y,D,Z)$.
-See Appendix B for the proof.
-Remark 5 (Non-learnability).
+Then, the equalities (4.3)-(4.11) are sharp testable implications for the validity of the instrumental variable Tobit Model in equations (2.8) coupled with assumptions 3-4.
+We collect these results more formally in Theorem 2.
+Theorem 2.
+Suppose that the IV Tobit model (2.8) or its reduced form representation (2.12) along with assumptions 3-4 hold.
+Then, the parameters $\alpha_{0}$, $\alpha_{1}$, $\gamma_{0}$, $\gamma_{1}$, $\sigma^{2}_{U}$, $\rho_{UV}$ are identified, and equalities (4.3)-(4.11) must hold for all $c,d,z\in\mathcal{Y}\times\mathcal{D}\times\mathcal{Z}$.
+Furthermore, these equalities are sharp, that is, whenever they hold, it is possible to construct a vector of $(\tilde{Y},\tilde{D},\tilde{U},\tilde{V},Z)$ or equivalently $(\tilde{Y},\tilde{D},\tilde{W},\tilde{V},Z)$ that satisfies model 2.8 and 2.12, Assumptions 3 and 4, and induces the observed distribution on the data $(Y,D,Z)$.
+The proof for the equalities follows from the discussion above, and further details about sharpness are presented in Appendix B. They serve as the basis for the test procedure described in Section 5.
+Proposition 3 (Non-learnability).
 The testable implications and sharpness discussed above show that the IV-Tobit model is generally refutable.
@@
 See Appendix E for the proof.
-Then, the equalities 3.4-3.12 are sharp testable implications for the validity of the instrumental variable Tobit Model in 2.8 coupled with assumptions 3-4.
-They serve as the basis for the test procedure described in Section 5.
-4. Identification
-In this section, we provide heuristic arguments for the identification of the parameters of the respective models, expanding on Han and Vytlacil (2017) and Acerenza, Bartalotti, and Kédagni (2023).
-4.1. Classic Tobit model
-Equation (3.3) established that in the classic Tobit model,
-$1-P(Y=0|D)=\Phi(\alpha_{0}+\alpha_{1}D).$
-Since the standard normal CDF, $\Phi(.)$, is monotonic, we use its inverse to obtain:
-$\Phi^{-1}(1-P(Y=0|D))=\alpha_{0}+\alpha_{1}D.$
-Thus, $\alpha_{1}=\frac{Cov(\Phi^{-1}(1-P(Y=0|D)),D)}{Var(D)}$, and $\alpha_{0}=E(\Phi^{-1}(1-P(Y=0|D)))-\alpha_{1}E(D)$.
-The coefficients have been normalized by dividing by the latent variable’s ($U$) square root of the variance, denoted by $\sigma$.
-Naturally, this implies that the ratio of the original parameters is identified since $P(Y=0|D=d)=\Phi(\alpha_{0}+\alpha_{1}D)=\Phi(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D)$.
-This allows the identification of the inverse mills ratio, which is a known function of the standard normal density, $\lambda(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D))$.
-Then,
-$E(Y|D,Y>0)=\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D+\sigma\lambda(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D))$
-Which in turn implies:
-$Y=\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D+\sigma\lambda(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D))+e$
-with $e$ mean independent of $D,\lambda(\cdot)$.
-Since $D$ is known and $\lambda(\cdot)$ was already identified then by considering the moment conditions for $Cov(.,D)$, $Cov(.,\lambda(\cdot))$ and $E(.)$ we get a three by three linear system with a unique solution for $\tilde{\alpha}_{0},\tilde{\alpha}_{1},\sigma$.
-In other words, the normality assumption allows identification by leveraging the relationship of the variable of interest and the outcome at the accumulation point ($Y=0$) to obtain the scaled parameters, and the same relationship when $Y>0$ to overcome the normalization.
-However, the reliance on the normality and linearity assumptions underlines the importance of adequately testing the model’s structure and assumptions in empirical applications.
-4.2. IV Tobit model
-Turning the focus to the IV-Tobit model, the first stage is readily identified from the linear structure, $\gamma_{1}=\frac{Cov(Z,D)}{Var(Z)}$, $\gamma_{0}=E(D)-\gamma_{1}E(Z)$.
-Now note that calculating $P(Y=0|Z)$:
-$1-P(Y=0|Z)=\Phi(-\beta_{0}-\beta_{1}Z)$
-Similarly to subsection 4.1, by inverting the normal CDF we obtain $-\beta_{1}=\frac{Cov(\Phi^{-1}(1-P(Y=0|Z)),Z)}{Var(Z)}$, $-\beta_{0}=E(\Phi^{-1}(1-P(Y=0|Z)))+\beta_{1}E(Z)$.
-The relationship between $\beta_{0},\beta_{1},\gamma_{0},\gamma_{1}$ identifies $(\alpha_{0},\alpha_{1})$.
-Finally, to identify $\rho$, let $s=c-\beta_{0}-\beta_{1}z$ and $t=d-\gamma_{0}-\gamma_{1}z$.
-Then, by using equation 3.7:
-$P(Y\leq c,D\leq d|Z=z)$ $=$ $\Phi_{W,V}(s,t;\rho)$
-Which by classic results of bivariate normal random variables if we differentiate with respect to $\rho$ is:
-$\frac{\partial P(Y\leq c,D\leq d|Z=z)}{\partial\rho}$ $=$ $\phi_{W,V}(s,t;\rho)$
-Where $\phi_{a,b}$ is the bivariate normal probability density, which is positive for any $s,t$. Hence, $\Phi_{W,V}(.,.;\rho)$ is monotonic in $\rho$ and thus invertible. So $\rho$ can be identified.
-Recall the previous coefficients are normalized, a two-step procedure can be followed to identify the full original covariance matrix and the reduced form covariance matrix (Wooldridge, 2010, p.
-683-84).
+Similarly to the discussion in Proposition 2, we conjecture that the approaches proposed in Theorem 2 can be adapted to models in which the researcher is willing to assume a known joint distribution for the latent errors, $(U,V)$, replacing the normal distribution, as long as the model is identified.
 5. Testing procedure
-To test the sharp equalities, we rewrite each of these equalities as two inequalities (Mourifié and Wan, 2017; Acerenza, Bartalotti, and Kédagni, 2023). For a concrete example, we first note that
-$P(c_{1}\leq Y|D=d)=1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)\iff E[1\{c_{1}\leq Y\}-1+\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]=0.$
+To test the sharp equalities, we rewrite each of these equalities as two inequalities (mourifie_testing_2017; acerenza_testing_2023). For a concrete example, we first note that
+$P(c_{1}\leq Y|D=d)=1-\Phi\left(\frac{c_{1}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}d\right)\iff E\left[1\{c_{1}\leq Y\}-1+\Phi\left(\frac{c_{1}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}D\right)|D\right]=0.$
 We rewrite each of these moment equalities implied by the restrictions on the empirical distribution as moment inequalities
-$E[1\{c_{1}\leq Y\}-1+\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]\leq 0,E[-1\{c_{1}\leq Y\}+1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]\leq 0.$
-In doing this for all the restrictions on the empirical distribution, we can implement a test relying on existing intersection bounds inferential methods such as Chernozhukov, Lee, and Rosen (2013), which is specifically suited to test conditional moment inequalities.
+$E\left[1\{c_{1}\leq Y\}-1+\Phi\left(\frac{c_{1}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}D\right)|D\right]\leq 0,E\left[-1\{c_{1}\leq Y\}+1-\Phi\left(\frac{c_{1}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}D\right)|D\right]\leq 0.$
+In doing this for all the restrictions on the empirical distribution, we can implement a test relying on existing intersection bounds inferential methods such as chernozhukov_intersection_2013, which is specifically suited to test conditional moment inequalities.
 Note that the equalities for the classic Tobit hold for any pair of constants $(c_{0},c_{1})$, and the ones from the IV Tobit hold for pairs of $(c_{0},c_{1})$ and $(d_{0},d_{1})$.
@@
 Classic Tobit model
-The sharp testable equalities for every $c_{0},c_{1}\in\mathcal{Y}$ are given by equations 3.1 and 3.3.
-For a partition of the support of $Y$ into $K$ arbitrary chosen sets $C_{k}=(0,c_{k})$ such that $C_{k}\in C_{k+1}$, the following set of sufficient inequalities are, for some chosen values of $c_{k},c_{k+1}\in\mathcal{Y}$, related to the components of 3.1-3.3.
+The sharp testable equalities for every $c_{0},c_{1}\in\mathcal{Y}$ are given by equations (4.1) and (4.2).
+For a partition of the support of $Y$ into $K$ arbitrary chosen sets $C_{k}=(0,c_{k})$ such that $C_{k}\in C_{k+1}$, the following set of sufficient inequalities are, for some chosen values of $c_{k},c_{k+1}\in\mathcal{Y}$, related to the components of equations (4.1)-(4.2).
 The formulation of the inequalities considered will depend on each partition’s location on the support of the outcome variable $Y$. Let $c_{1}=0$ and $W_{k}$ be
-$W_{k}=\left\{\begin{array}[]{cc}1\{Y=0\}-(1-\Phi(\alpha_{0}+\alpha_{1}D)),&\text{ if }k=0\\ 1\{c_{k}<Y\leq c_{k+1}\}-\Phi(c_{k+1}-\alpha_{0}-\alpha_{1}D)+\Phi(c_{k}-\alpha_{0}-\alpha_{1}D),&\text{ if }1\leq k<K\\ 1\{c_{K}<Y\}-(1-\Phi(c_{K}-\alpha_{0}-\alpha_{1}D)),&\text{ if }k=K\\ \end{array}\right.$
+[…]
 The intersection bounds framework considers the following $2(K+1)$ inequalities
-$\sup_{d}E[W_{k}|D=d]$$\leq$ 00 0
+$\sup_{d}E[W_{k}|D=d]$ $\leq$ $0$
 $\sup_{d}E[-W_{k}|D=d]$ $\leq$ $0,\text{ for }k=0,\dots,K$
@@
 $\max_{k}\sup_{D}\theta_{k}(D)\leq 0$ (5.2)
-where $\theta_{k}(D)$ collects all the inequalities being tested. The decision rule for the test is given by Chernozhukov, Lee, and Rosen (2013), we reject $H_{0}$ if
+where $\theta_{k}(D)$ collects all the inequalities being tested. The decision rule for the test is given by chernozhukov_intersection_2013, we reject $H_{0}$ if
 $\hat{\theta}_{1-\alpha}\equiv\max_{k}\sup_{D}\left\{\hat{\theta}(D,k)-\kappa_{1-\alpha}\hat{s}(D,k)\right\}>0,$ (5.3)
@@
 For the instrumental variable Tobit model, the continuous support for both the outcome and treatment poses challenges to the implementation of the test.
-[^8] Note that a discrete treatment can also be accommodated as an intermediate step between the current derivation and the derivations from Acerenza, Bartalotti, and Kédagni (2023).
-The sharp testable equalities for every $c_{0},c_{1}\in\mathcal{Y}$ and $d_{0},d_{1}\in\mathcal{D}$ are given by equations 3.4-3.12.
+[^9] Note that a discrete treatment can also be accommodated as an intermediate step between the current derivation and the derivations from acerenza_testing_2023.
+The sharp testable equalities for every $c_{0},c_{1}\in\mathcal{Y}$ and $d_{0},d_{1}\in\mathcal{D}$ are given by equations (4.3)-(4.11).
 Consider a partition of the support of $Y$ into $K$ arbitrary chosen sets $C_{k}=(0,c_{k})$ such that $C_{k}\in C_{k+1}$ and of the support of $D$ into $Q$ arbitrary chosen sets $D_{q}=(0,d_{q})$ such that $D_{q}\in D_{q+1}$, the following set of sufficient inequalities are related to the components of (4.3)-(4.11).
@@
 Let $W_{kq}$ be given by,
-$W_{kq}=\left\{\begin{array}[]{cc}1\{Y=0\}1\{D\leq d_{0}\}-\Phi^{2}_{d_{0}},&\text{ if }k=0,q=0\\ 1\{Y=0\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{2}_{d_{q},d_{q+1}},&\text{ if }k=0,1\leq q<Q\\ 1\{Y=0\}1\{D\geq d_{Q}\}-\Phi^{2}_{d_{Q}},&\text{ if }k=0,q=Q\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{D\leq d_{0}\}-\Phi^{1}_{c_{k+1},c_{k},d_{0}},&\text{ if }1\leq k<K,q=0\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{1}_{c_{k+1},c_{k},d_{q+1},d_{q}},&\text{ if }1\leq k<K,1\leq q<Q\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{D\geq d_{Q}\}-\Phi^{1}_{c_{k+1},c_{k},d_{Q}},&\text{ if }1\leq k<K,q=Q\\ 1\{Y\geq c_{K}\}1\{D\leq d_{0}\}-\Phi^{1}_{c_{K},d_{0}},&\text{ if }k=K,q=0\\ 1\{Y\geq c_{K}\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{1}_{c_{K},d_{q+1},d_{q}},&\text{ if }k=K,1\leq q<Q\\ 1\{Y\geq c_{K}\}1\{D\geq d_{Q}\}-\Phi^{1}_{c_{K},d_{Q}},&\text{ if }k=K,q=Q.\\ \end{array}\right.$
+[…]
 Where we used the simplifying notation defined in equations (4.3)-(4.11) for each partition of the support for $Y$ and $D$.
 The intersection bounds framework considers the following $2(K+1)(Q+1)$ inequalities
-$\sup_{z}E[W_{kq}|Z=z]$$\leq$ 00 0
+$\sup_{z}E[W_{kq}|Z=z]$ $\leq$ $0$
 $\sup_{z}E[-W_{kq}|Z=z]$ $\leq$ $0,\text{ for }k=0,\dots,K;q=0,\dots,Q.$
@@
 $\max_{k,q}\sup_{Z}\theta_{k}q(Z)\leq 0$ (5.5)
-where $\theta_{k}q(Z)$ collects all the inequalities being tested. The decision rule for the test is given by Chernozhukov, Lee, and Rosen (2013), we reject $H_{0}$ if
+where $\theta_{k}q(Z)$ collects all the inequalities being tested. The decision rule for the test is given by chernozhukov_intersection_2013, we reject $H_{0}$ if
 $\hat{\theta}_{1-\alpha}\equiv\max_{k,q}\sup_{Z}\left\{\hat{\theta}(Z,k,q)-\kappa_{1-\alpha}\hat{s}(D,k,q)\right\}>0,$ (5.6)
 where $\hat{\theta}(Z,k,q)$ is a nonparametric estimator for $\theta_{kq}(Z)$, $\hat{s}(D,k,q)$ its standard error, and $\kappa_{1-\alpha}$ is a critical value at the significance level $\alpha$.
-To implement the test within the Chernozhukov, Lee, and Rosen (2013) intersection bounds inferential method, we use the CLR Stata package described in Chernozhukov et al.
-(2015).
-The parameters in the relevant model, for example, $\beta_{0},\beta_{1},\gamma_{0},\gamma_{1},\rho$, are replaced by their maximum likelihood estimators (MLE), and asymptotic validity of this “plug-in” test follows from the argument described by (Acerenza, Bartalotti, and Kédagni, 2023, Appendix B).
+To implement the test within the chernozhukov_intersection_2013 intersection bounds inferential method, we use the CLR Stata package described in chernozhukov_implementing_2015.
+The parameters in the relevant model, for example, $\beta_{0},\beta_{1},\gamma_{0},\gamma_{1},\rho$, are replaced by their maximum likelihood estimators (MLE), and asymptotic validity of this “plug-in” test follows from the argument described by (acerenza_testing_2023, Appendix B).
 Some additional details on the test implementation are discussed in Section 6.
@@
 3.
-Violations due to the empirical distributions of the residuals differing from the implied distributions in certain values of the treatment (Classic Tobit) or instrument (IV Tobit), which indicate violations of the exogeneity of treatment or instrument (Kédagni and Mourifié, 2020).
+Violations due to the empirical distributions of the residuals differing from the implied distributions in certain values of the treatment (Classic Tobit) or instrument (IV Tobit), which indicate violations of the exogeneity of treatment or instrument (kedagni_generalized_2020).
+Remark 2.
+Alternative approaches to test the classic and IV Tobit could be considered.
+For example, an intuitive approach would be calculate the residuals $\hat{U}=Y-\hat{\alpha}_{0}-\hat{\alpha}_{1}D$ and compare its distribution to the (truncated) std.
+normal through one of the usual normality tests in the literature.
+[^10] We received this suggestion in several seminars and from anonymous referees, to whom we are thankful.
+One of the challenges in doing so is that the latent variable is recoverable (as estimated by the residuals) only if $Y=\alpha_{0}+\alpha_{1}D+U>0$, while for all observations for which $Y=0$ the only information available is that $\alpha_{0}+\alpha_{1}D\leq-U$.
+Thus, in our case, all observations that are “at the corner” would only give us information about U being below a certain truncation value, which depends on $D$ (and any other covariates included in the model, see below), which is a non-pivotal quantity and is not associated with any well-established distributional test.
+A second approach would follow li_nonparametric_2023 by testing for correct parametric functional forms for $E[1\{c_{0}\leq Y\leq c_{1}\}|D]$ against $\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}D)$, which is similar in spirit to our test if one incorporates all the equalities proposed in Theorem 1-2.
+That approach would also require partitioning the support of the variables being considered.
+A nonparametric test statistic could be used, following li_nonparametric_2023, the details of which are beyond the scope of this manuscript.
+Both approaches would present difficulties of implementation at least as important as the ones faced by proposed test based in the inferential approach by chernozhukov_implementing_2015 discussed above.
+While recognizing that CLR may lead to a conservative test, the relatively easy implementation of the test using available statistical packages is an attractive feature.
+Further research on different approaches and their properties could lead to valuable alternatives.
 Including Covariates
@@
 [^11] A theoretically interesting approach would partition the joint support of all the exogenous variables and compute the test statistics across all parts of the grid.
-However, this approach would entail significant implementation challenges since the statistical package for Chernozhukov et al.
-(2015) allows for only one exogenous covariate.
+However, this approach would entail significant implementation challenges since the statistical package for chernozhukov_implementing_2015 allows for only one exogenous covariate.
 This procedure can be cumbersome when there are many covariates or when they are continuous.
-Another route that is less computationally intensive follows Acerenza, Bartalotti, and Kédagni (2023).
+Another route that is less computationally intensive follows acerenza_testing_2023.
 We illustrate it for the $IV$-tobit case, and a similar logic could be applied to the classic Tobit case.
 Thus, with covariates added to the linear index function, we have,
-$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\beta_{0}+\beta_{1}Z+\beta_{2}X+W\\ D&=&\gamma_{0}+\gamma_{1}Z+\gamma_{2}X+V\end{array}\right.$ (5.10)
+$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\tilde{\beta}_{0}+\tilde{\beta}_{1}Z+\tilde{\beta}_{2}X+\tilde{W}\\ D&=&\tilde{\gamma}_{0}+\tilde{\gamma}_{1}Z+\tilde{\gamma}_{2}X+\tilde{V}\end{array}\right$.
+(5.10)
 We extend Assumption 3 to formalize covariates’ exogeneity:
@@
 Under Assumptions 4-5, by a similar argument to that used in Section 4, any given generic testable implication becomes
-$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z,X=x)$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{1}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
--- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{1}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
--- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{0}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
-$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{0}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
+$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z,$ $X$ $=x)=\Phi_{W,V}\left(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z-\frac{\tilde{\beta}_{2}}{\sigma_{\tilde{W}}}x,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z-\frac{\tilde{\gamma_{2}}}{\sigma_{\tilde{V}}}x;\rho\right)$ (5.11)
+- $\Phi_{W,V}\left(\frac{c_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z-\frac{\tilde{\beta}_{2}}{\sigma_{\tilde{W}}}x,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z-\frac{\tilde{\gamma_{2}}}{\sigma_{\tilde{V}}}x;\rho\right)$
+- $\Phi_{W,V}\left(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z-\frac{\tilde{\beta}_{2}}{\sigma_{\tilde{W}}}x,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z-\frac{\tilde{\gamma_{2}}}{\sigma_{\tilde{V}}}x;\rho\right)$
+$+$ $\Phi_{W,V}\left(\frac{c_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z-\frac{\tilde{\beta}_{2}}{\sigma_{\tilde{W}}}x,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z-\frac{\tilde{\gamma_{2}}}{\sigma_{\tilde{V}}}x;\rho\right)$
 for all $z\in\mathcal{Z}$ and $x\in\mathcal{X}$, where $x$ can be a vector. Equivalent conditions for the case that $Y=0$ and other parts of the support of $Y$ and $D$ can be similarly obtained.
 Implementation
-Since the test is carried out nonparametrically, we face challenges when the dimension of $X$ is large, especially with continuous covariates.
+Since the test is nonparametric, we face challenges when $X$ is high-dimensional, especially with continuous covariates.
 Another operational limitation is that the Stata’s clrbound package only allows for one conditioning variable at a time.
 For those practical reasons, we propose implementing a weaker, non-sharp, version of the testable equalities.
-Continuing with the example of Equality 5, we can integrate over the covariates $X$ (or alternatively, $Z$), taking advantage of the fact that $\mathbb{E}[W|Z=z,X=x]=0$ implies $\mathbb{E}[W|Z=z]=0$ and $\mathbb{E}[W|X=x]=0$ for all random variables $W$.
+Continuing with the example of Equality 5.11, we can integrate over the covariates $X$ (or alternatively, $Z$), taking advantage of the fact that $\mathbb{E}[W|Z=z,X=x]=0$ implies $\mathbb{E}[W|Z=z]=0$ and $\mathbb{E}[W|X=x]=0$ for all random variables $W$.
 Implementation becomes very similar to the IV-Tobit discussed above by redefining the simplifying notation in equations (4.3)-(4.11), with the only difference being the inclusion of $X$ on the linear indexes in $\Phi_{W}(\cdot)$ and $\Phi_{W,V}(\cdot,\cdot,\cdot)$.
-For example,
-$\Phi^{1}_{c_{1},c_{0},d_{1},d_{0}}$$\equiv$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{1}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
--- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{1}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
--- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{0}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
-$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{0}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho),$
-and similarly for the other terms.
+For example, $\Phi^{1}_{c_{1},c_{0},d_{1},d_{0}}$ is given by equation (5.11) and similarly for the other terms.
 Then, compute the new $W_{kq}$ in the same manner as in Section 5, and the intersection bounds framework considers the similar $2(K+1)(Q+1)$ inequalities, based on the partition of the support of $Y$ and $D$.
-$\sup_{z}E[W_{kq}|Z=z]$$\leq$ 00 0
+$\sup_{z}E[W_{kq}|Z=z]$ $\leq$ $0$
 $\sup_{z}E[-W_{kq}|Z=z]$ $\leq$ $0,\text{ for }k=0,\dots,K;q=0,\dots,Q.$
 This is the test procedure that we implement for the empirical application in Section 8.
-Under Assumption 5, one could base the test in conditioning on a particular covariate $X_{C}$ instead of $Z$ by integrating the sharp Equality 5 over $Z$ and then implementing the intersection bounds procedure using $X_{C}$ as the sole conditioning variable.
+Under Assumption 5, one could base the test on conditioning on a particular covariate $X_{C}$ instead of $Z$ by integrating the sharp Equality (5.11) over $Z$ and then implementing the intersection bounds procedure using $X_{C}$ as the sole conditioning variable.
 [^12] Assumption 5 constraints the relationship between the covariates and the latent error terms in such a way that one could combine the information on all $X$ and $Z$ in an index and construct non-sharp testable equalities that would hold conditional on this index.
@@
 When both the outcome and treatment are continuous, evaluating all possible equalities is technically challenging.
-We focus on a non-sharp set of the equalities by evaluating them at different grid partitions of their support, in a similar spirit to Honoré and Hu (2020) which is particularly well suited for continuous supports.
+We focus on a non-sharp set of the equalities by evaluating them at different grid partitions of their support, in a similar spirit to honore2020selection which is particularly well suited for continuous supports.
 Naturally, this choice makes the test less powerful as we don’t consider the continuum of equalities derived in Section 3, but is justified by the ease of implementation of the procedure based on intersection bounds and the performance of the test on the simulations below.
@@
 [^14] The simulations presented in Section 6 have limited sample sizes and a relatively small number of equalities being tested due to the long-running time and computational constraints when repeating the test procedure thousands of times.
-When implementing the intersection bounds in STATA using the package clrbound Chernozhukov et al.
-(2015), the researcher must determine the range of values of the exogenous variable for which each equality will be evaluated.
-To guarantee the feasibility of the procedure, we adjust the evaluation points for $D$ ($Z$) to the first and $99^{th}$ percentiles of the exogenous variable in each partition of the support for the outcome $Y$ [^13] See details on implementation on the simulation replication STATA code.
+When implementing the intersection bounds in STATA using the package clrbound chernozhukov_implementing_2015, the researcher must determine the range of values of the exogenous variable for which each equality will be evaluated.
+To guarantee the feasibility of the procedure, we adjust the evaluation points for $D$ ($Z$) to the first and $99^{th}$ percentiles of the exogenous variable in each partition of the support for the outcome YY.
+[^15] See details on implementation on the simulation replication STATA code.
 In evaluating the finite sample performance of the proposed tests, we consider a continuous treatment $D$ following the data generating process described in Equation 6.8.
-$\left\{\begin{array}[]{lcl}Y&=&\max\{Y^{*},0\}\\ Y^{*}&=&D+U\\ D&=&2Z-V\\ \begin{pmatrix}U\\ V\\ Z\end{pmatrix}&\sim&\mathcal{N}(\mathbf{0}_{3},\mathbf{\Sigma})\\ \mathbf{\Sigma}&=&\begin{pmatrix}1&\rho_{uv}&\rho_{uz}\\ \rho_{uv}&1&\rho_{vz}\\ \rho_{uz}&\rho_{vz}&1\end{pmatrix}\\ \rho_{uz}&=&\rho_{vz}=0\\ \rho_{uv}&=&\rho\end{array}\right.$ (6.8)
+$\left\{\begin{array}[]{lcl}Y&=&\max\{Y^{*},0\}\\ Y^{*}&=&D+U\\ D&=&2Z-V\\ \begin{pmatrix}U\\ V\\ Z\end{pmatrix}&\sim&\mathcal{N}(\mathbf{0}_{3},\mathbf{\Sigma})\\ \mathbf{\Sigma}&=&\begin{pmatrix}1&\rho_{uv}&\rho_{uz}\\ \rho_{uv}&1&\rho_{vz}\\ \rho_{uz}&\rho_{vz}&1\end{pmatrix}\\ \rho_{uz}&=&\rho_{vz}=0\\ \rho_{uv}&=&\rho\end{array}\right$.
+(6.8)
 Where $\mathbf{0}_{p}$ is a $p\times 1$ vector of zeroes.
@@
 Larger values of $\rho$ produce more acute violations of the null hypothesis.
-Table 2 presents the results for the Classic Tobit test: columns 3-4 report the MSE, while the remaining columns show empirical rejection rates.
+Table 2 presents the results for the Classic Tobit test.
 As expected, the power of the test increases with larger $\rho$ and bigger sample sizes.
-[^14] Simulations for test performance when violations of normality of the latent errors occur are presented in Appendix F.
 Table 2. Classic Tobit test power for violations in exogeneity
-$N$ $\rho$ MSE$(\beta_{0})$ MSE$(\beta_{1})$ $\alpha=10\%$ $\alpha=5\%$ $\alpha=1\%$
-5000 0.10 0.0005 0.0005 8.60% 6.00% 3.20%
-0.50 0.0005 0.0104 16.03% 10.82% 5.21%
-0.75 0.0005 0.0227 36.00% 27.80% 14.20%
-0.80 0.0005 0.0259 42.00% 30.60% 16.80%
-0.90 0.0004 0.0326 62.40% 48.40% 27.60%
-0.95 0.0004 0.0360 71.34% 59.32% 35.87%
-8000 0.10 0.0003 0.0005 8.40% 5.00% 2.00%
-0.50 0.0003 0.0102 17.43% 12.22% 5.61%
-0.75 0.0003 0.0225 45.58% 36.14% 20.08%
-0.80 0.0003 0.0256 53.80% 39.80% 18.00%
-0.90 0.0003 0.0323 79.00% 65.40% 36.40%
-0.95 0.0002 0.0363 85.57% 75.15% 49.70%
-10000 0.10 0.0003 0.0005 10.04% 7.43% 2.41%
-0.50 0.0002 0.0101 15.80% 10.00% 3.80%
-0.75 0.0002 0.0226 44.80% 32.80% 16.00%
-0.80 0.0002 0.0256 56.60% 41.00% 21.60%
-0.90 0.0002 0.0326 79.60% 66.60% 36.60%
-0.95 0.0002 0.0361 90.56% 79.52% 51.61%
+NN $\rho$ $\alpha=10\%$ $\alpha=5\%$ $\alpha=1\%$
+5000 0.10 8.60% 6.00% 3.20%
+0.50 16.03% 10.82% 5.21%
+0.75 36.00% 27.80% 14.20%
+0.80 42.00% 30.60% 16.80%
+0.90 62.40% 48.40% 27.60%
+0.95 71.34% 59.32% 35.87%
+8000 0.10 8.40% 5.00% 2.00%
+0.50 17.43% 12.22% 5.61%
+0.75 45.58% 36.14% 20.08%
+0.80 53.80% 39.80% 18.00%
+0.90 79.00% 65.40% 36.40%
+0.95 85.57% 75.15% 49.70%
+10000 0.10 10.04% 7.43% 2.41%
+0.50 15.80% 10.00% 3.80%
+0.75 44.80% 32.80% 16.00%
+0.80 56.60% 41.00% 21.60%
+0.90 79.60% 66.60% 36.60%
+0.95 90.56% 79.52% 51.61%
+Based on $500$ replications.
+To consider a violation with respect to the violation of the normality of errors.
+We consider a modified DGP from Equation 6.8, by seting $(V,Z)^{\top}\sim\mathcal{N}(\mathbf{0}_{2},\mathbf{I}_{2})$ and $U\sim F(\cdot)$.
+Table 3 presents the power results when $F(\cdot)$ is uniform, lognormal, and t-student distributions.
+As expected, the rejection rate for the t-student test with 80 degrees of freedom is close to the nominal size of the test, since it represents a very mild violation of the normality assumption.
+In the other three cases, which deviate significantly from the normality assumption, the test rejects the null hypothesis in most instances.
+This example demonstrates that the proposed test is effective in identifying violations of the distributional assumptions in the classic Tobit model.
+Table 3. Classic Tobit test power for violations in the error structure
+NN $U$ $\alpha=10\%$ $\alpha=5\%$ $\alpha=1\%$
+5000 t-student (df=80) 7.21% 4.01% 1.40%
+t-student (df=5) 92.20% 81.80% 48.80%
+LogNormal 100.00% 100.00% 100.00%
+Uniform 100.00% 100.00% 100.00%
+8000 t-student (df=80) 7.00% 4.80% 1.20%
+t-student (df=5) 99.20% 97.80% 87.58%
+LogNormal 100.00% 100.00% 100.00%
+Uniform 100.00% 100.00% 100.00%
+10000 t-student (df=80) 5.42% 2.41% 1.00%
+t-student (df=5) 99.20% 98.40% 90.58%
+LogNormal 100.00% 100.00% 100.00%
+Uniform 100.00% 100.00% 100.00%
 Based on $500$ replications.
@@
 N $\rho_{uv}$ $\alpha=10\%$ $\alpha=5\%$ $\alpha=1\%$
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+1000 0 30.20% 25.60% 18.00%
+0.5 34.80% 29.00% 20.00%
+0.8 29.60% 23.40% 15.40%
+2000 0 20.20% 15.60% 8.20%
+0.5 21.40% 16.60% 9.00%
+0.8 22.60% 15.20% 9.40%
+5000 0 10.80% 7.00% 3.20%
+0.5 11.40% 9.80% 3.40%
+0.8 14.40% 9.60% 4.60%
+8000 0 7.40% 3.80% 1.80%
+0.5 9.20% 5.60% 1.80%
+0.8 15.20% 10.00% 3.60%
+10000 0 8.68% 5.21% 1.74%
+0.5 10.00% 6.60% 1.80%
+0.8 12.00% 8.60% 3.40%
 Based on $500$ replications.
@@
 The second assumption is the exogeneity of the treatment or potential instrument.
-Here, we provide a non-exhausting survey of existing work.
-Cragg (1971) maintains the normality of the errors, linearity of the index and exogeneity but relaxes the way censoring occurs in comparison to the latent structure of the censored outcome.
+Here, we provide a non-exhaustive survey of existing work.
+cragg1971some maintains the normality of the errors, linearity of the index and exogeneity but relaxes the way censoring occurs in comparison to the latent structure of the censored outcome.
 Specifically, while the latent outcome is still modelled by $Y^{*}=\alpha_{0}+\alpha_{1}D+U$, they allow for the censoring to depend on a different linear index, $P(Y^{*}>0)=P(\gamma_{0}+\gamma_{1}D+e)$, increasing the model’s flexibility.
-Powell (1984) relaxes the parametric structure of the errors while maintaining the latent linear index and treatment exogeneity, and estimates the parameters of interest by least absolute deviations.
-Powell (1986) also maintains linearity of the latent index and treatment exogeneity, but relaxes normality by imposing symmetrical distributions to the latent errors, which leads to estimation by symmetrically censored least squares.
-Newey (1987a) relaxes exogeneity of the treatment, relying on normality and an instrumental variable to identify the model, which is estimated by generalized least squares.
-Honoré and Powell (1994) relax linearity and mean independence of the unobservable with respect to the treatment to exploit the idea that, although $Y^{*}_{i}-\alpha_{0}-\alpha_{1}D_{i}$ is not mean-independent of $D_{i}$, one can trim any pair of residuals $Y^{*}_{i}-\alpha_{0}-\alpha_{1}D_{i}$ and $Y^{*}_{j}-\alpha_{0}-\alpha_{1}D_{j}$, and the trimmed residuals are independent and identically distributed conditional on $D_{i},D_{j}$.
+POWELL1984303 relaxes the parametric structure of the errors while maintaining the latent linear index and treatment exogeneity, and estimates the parameters of interest by least absolute deviations.
+powell1986symmetrically also maintains linearity of the latent index and treatment exogeneity, but relaxes normality by imposing symmetrical distributions to the latent errors, which leads to estimation by symmetrically censored least squares.
+newey1987efficient relaxes exogeneity of the treatment, relying on normality and an instrumental variable to identify the model, which is estimated by generalized least squares.
+honore1994pairwise relax linearity and mean independence of the unobservable with respect to the treatment to exploit the idea that, although $Y^{*}_{i}-\alpha_{0}-\alpha_{1}D_{i}$ is not mean-independent of $D_{i}$, one can trim any pair of residuals $Y^{*}_{i}-\alpha_{0}-\alpha_{1}D_{i}$ and $Y^{*}_{j}-\alpha_{0}-\alpha_{1}D_{j}$, and the trimmed residuals are independent and identically distributed conditional on $D_{i},D_{j}$.
 They estimate the model by identically censored least absolute deviations and identically censored least squares (ICLS).
-Das (2002) estimates a model using symmetrically censored least squares that relaxes exogeneity of the treatment and normality of the errors.
+das2002estimators estimates a model using symmetrically censored least squares that relaxes exogeneity of the treatment and normality of the errors.
 To achieve that they rely on instrumental variables, linearity of the mean of the structural error conditional on the reduced form error, and mean independence of the reduced form error.
-Blundell and Powell (2007) proposes a control variable approach that relaxes exogeneity and normality but maintains the latent linear structure ($\alpha_{0}+\alpha_{1}D+U$).
+blundell2007censored proposes a control variable approach that relaxes exogeneity and normality but maintains the latent linear structure ($\alpha_{0}+\alpha_{1}D+U$).
 Crucially, they impose that the distribution (or quantiles) of the latent error conditional on the treatment and instrument is only a function of the control variable $V=D-\pi(Z)$, which isolates the endogenous variation on the treatment.
@@
 This allows them to estimate the effect of the treatment by censored quantile instrumental variable regression augmented by a control variable given by the quantiles of $U$ conditional on $V$ at the quantile of interest.
-In a similar spirit, Chernozhukov, Fernández-Val, and Kowalski (2015) focuses on conditional quantile functions and flexible approaches to estimate the control variable in the first stage.
-Finally, Chesher, Kim, and Rosen (2023) provides partial identification results for a general alternative by relaxing the exogeneity of the treatment and instrument, linearity of the latent index and imposing no parametric structure of the error term.
-They characterize the identified set for the parameters of interest following the Generalized Instrumental Variables framework (Chesher and Rosen, 2017), relying on the assumption that the relationship of $Y^{*}$ to treatment and errors is continuous and monotonic in the errors.
+In a similar spirit, chernozhukov2015quantile focuses on conditional quantile functions and flexible approaches to estimate the control variable in the first stage.
+Finally, chesher2023iv provides partial identification results for a general alternative by relaxing the exogeneity of the treatment and instrument, linearity of the latent index and imposing no parametric structure of the error term.
+They characterize the identified set for the parameters of interest following the Generalized Instrumental Variables framework (chesher2017generalized), relying on the assumption that the relationship of $Y^{*}$ to treatment and errors is continuous and monotonic in the errors.
 Their approach uses the residual sets associated with the structure of the latent function and conditional probability of the error term given potential instruments.
@@
 In this subsection, we present an approach that partially identifies the effect of an endogenous treatment variable by replacing the normality and exogeneity assumptions with a monotonicity in treatment selection constraint.
-While less general than Chesher, Kim, and Rosen (2023), this approach is easy to implement and could be useful to empirical researchers.
+While less general than chesher2023iv, this approach is easy to implement and could be useful to empirical researchers.
 Consider the model that maintains linearity (or a known structure of $Y^{*}$ up to a finite number of parameters),
@@
 Let $E(U|D=d,Y>0)\equiv\Gamma(d)$. Then, for any $d>d^{*}$ we either have $\Gamma(d)<\Gamma(d^{*})$ or $\Gamma(d)>\Gamma(d^{*})$.
-Assumption 6 is common in the partial identification literature (Jiang, Chiba, and VanderWeele, 2014; Manski, 1997; Manski and Pepper, 2000; Okumura and Usui, 2014).
+Assumption 6 is common in the partial identification literature (jiang2014monotone; manski1997monotone; manski2000amonotone; okumura2014concave).
 In this context, we restrict the latent selection to be monotonic with respect to the treatment.
@@
 Since it is not as restrictive as imposing a parametric structure and independence, we can only partially identify the parameter of interest.
-Under Assumption 6 and the model described in 7.3, treatment and outcome are not independent. Note that,
+Under Assumption 6 and the model described in equations (7.3), treatment and outcome are not independent. Note that,
 $E(Y|D=d,Y>0)$ $=$ $\alpha_{0}+\alpha_{1}d+\Gamma(d)$
-$E(Y|D=d,Y>0)$displaystyle-- $\alpha_{0}-\alpha_{1}d=\Gamma(d)$
+$E(Y|D=d,Y>0)$ - $\alpha_{0}-\alpha_{1}d=\Gamma(d)$
 Then, for any two $d>d^{*}$ we have, by Assumption 6:
@@
 The bound can be more informative in the case of a continuous or multi-valued treatment. If $\Gamma(D)$ is differentiable we have:
-$\alpha_{1}$ $>$ $\frac{\partial E(Y|D=d,Y>0)}{\partial d}\text{, for all d.}$. (7.12)
+$\alpha_{1}$ $>$ $\frac{\partial E(Y|D=d,Y>0)}{\partial d}\text{, for all d.}$ (7.12)
 Since the inequality holds for any $d$ in the continuous case or for any $d,d^{*}$ for multi-valued discrete treatment, the linear index structure with constant parameters implies that tighter bounds for $\alpha_{1}$ are given by the maximum value of $\frac{\partial E(Y|D=d,Y>0)}{\partial d}$ across all possible points in the support for $D$.
@@
 In this case, the estimator is asymptotically normal and converges at a nonparametric rate that depends on the bandwidth $h$.
-Recent developments in Calonico, Cattaneo, and Farrell (2018, 2022) for estimation and optimal coverage error bandwidth and kernel selection methods that are nonparametric robust bias-corrected (RBC) can be used through their convenient implementation using the package nprobust (Calonico, Cattaneo, and Farrell, 2019).
+Recent developments in calonico2018effect; calonico2022coverage for estimation and optimal coverage error bandwidth and kernel selection methods that are nonparametric robust bias-corrected (RBC) can be used through their convenient implementation using the package nprobust (calonico2019nprobust).
+Alternatively, since $\alpha_{1}>sup_{d}\frac{\partial E(Y|D=d,Y>0)}{\partial d}$, one could consider obtaining confidence regions for $\alpha_{1}$ using a CLR approach similar to that described in Section 5 by considering nonparametric estimates of $\frac{\partial E(Y|D=d,Y>0)}{\partial d}$ or its discrete counterpart.
+[^17] We thank an anonymous referee for this suggestion.
 Remark 3 (On including covariates).
@@
 Tighter bounds are achieved by considering the maximum value of $\frac{\partial E(Y|D=d,X=x,Y>0)}{\partial d}$ across all possible points in the support for $D$ and $X$.
-8. Empirical Illustration: Lee (1995)
-In this section, we implement the proposed test to the data from Lee (1995).
+In practice, nonparametric estimates of these derivatives can be difficult, even for moderate numbers of covariates, and particularly challenging when multiple continuous covariates are in the conditioning set.
+The estimated values might be unstable, in which case using the maximum estimated value can lead to unreasonable bounds for $\alpha_{1}$.
+An easier to implement conservative alternative is to use the average derivatives with respect to $D$, $E\left[\frac{\partial E(Y|D=d,X=x,Y>0)}{\partial d}\right]$ and rely on bootstrapped standard errors for inference.
+8. Empirical Illustration: lee1995semi
+In this section, we implement the proposed test to the data from lee1995semi.
+[^18] Data availability statement: The data that support the findings of this study are openly available in the Journal of Applied Econometrics Data Archive at http://dx.doi.org/10.15456/jae.2022313.1130270920.
 Using the 1987 cross-section of the Michigan Panel Study of Income Dynamics, the authors study the responses of married women’s labor supply ($Y$) - measured in hours per year - to hundreds of dollars in “other” household income ($D$), which is endogenous.
@@
 Census, and this dummy variable seems to include workers listed in the categories “1-195 Professional, Technical, and Kindred Workers,” and “201–245 Managers and Administrators, except Farm.”
-Following Lee (1995), we proceed with the analysis focusing on the data for married couples with non-negative family total income or “other” household income and where the wife was of working age (18-64) and not self-employed.
+Following lee1995semi, we proceed with the analysis focusing on the data for married couples with non-negative family total income or “other” household income and where the wife was of working age (18-64) and not self-employed.
 These selections results in 3,277 observations, for which 26 percent of wage observations are censored.
 Table LABEL:Table:Lee presents the estimates obtained using the IV Tobit model.
-Table 4. IV Tobit specification for Lee (1995)
+Table 5. IV Tobit specification for lee1995semi
 MLE
@@
 $(10.813)$
-Other household income $-0.973^{***}$
+Other household income $0.973^{***}$
 $(0.373)$
@@
 $(3.251)$ $(15.249)$
-Age squared $-0.105^{***}$ $-1.221^{***}$
+Age squared $0.105^{***}$ $1.221^{***}$
 $(0.039)$ $(0.178)$
@@
 $(2.071)$ $(13.231)$
-Children under 5 $9.448$ $-500.332^{***}$
+Children under 5 $9.448$ $500.332^{***}$
 $(6.482)$ $(28.163)$
-Children between 6 and 13 $3.204$ $-211.687^{***}$
+Children between 6 and 13 $3.204$ $211.687^{***}$
 $(5.601)$ $(23.737)$
-Children between 14 and 17 $12.881$ $-16.878$
+Children between 14 and 17 $12.881$ 16.878-16.878
 $(9.342)$ $(39.152)$
-Nonwhite $-59.550^{***}$ $146.336^{***}$
+Nonwhite $59.550^{***}$ $146.336^{***}$
 $(10.210)$ $(51.105)$
@@
 $(13.954)$ $(61.036)$
-Local Unemployment $-8.165^{***}$ $-41.979^{***}$
+Local Unemployment $8.165^{***}$ $41.979^{***}$
 $(1.941)$ $(8.966)$
-Constant $-339.059^{***}$ $-337.347$
+Constant $39.059^{***} 33$7.347-337.347
 $(63.387)$ $(314.207)$
@@
 In particular, for women working positive hours, an increase of one thousand dollars in household income from other sources is estimated to reduce hours worked by 9.7 hours per year.
-The direction of the impacts at the intensive margin of hours worked follow intuitive patterns and are qualitative similar to those in Lee (1995).
+The direction of the impacts at the intensive margin of hours worked follows intuitive patterns and is qualitatively similar to those in lee1995semi.
 However, the model is rejected at conventional significance levels when we test for the IV-Tobit model’s validity, ($\hat{\theta}_{0.99}=0.1335>0$, $\hat{\theta}_{0.95}=0.1368>0$, and $\hat{\theta}_{0.90}=0.1385>0$).
@@
 Parameter IV-Tobit estimates $\alpha_{1}$’s Lower Bound
-$\alpha_{1}$0.973 -0.419
-conf.: confidence; LB: lower bound
+$\alpha_{1}$ -0.973 -0.419
+(0.373) (0.125)
+Note: Standard errors in parentheses.
 The first column of Table 6 repeats the estimate for $\alpha_{1}$ from Table 5.
-The second column reports the estimated lower bound for $\alpha_{1}$, obtained under Assumption 6 and latent linear index only based on the nonparametric estimate of the average derivative of the conditional expectation for yearly hours worked with respect to the household income from other sources, as described in Remark 7.
+The second column reports the estimated lower bound for $\alpha_{1}$, obtained under Assumption 6 and latent linear index only based on the nonparametric estimate of the average derivative of the conditional expectation for yearly hours worked with respect to the household income from other sources.
+As described in Remark 3, we opt to use the estimate for the average derivatives due to instability of the estimated derivatives across the different values of covariates and treatment.
+This is a conservative approach regarding the bounds for $\alpha$.
+Inference for the average derivative can be obtained by bootstrapping.
 Even after relaxing the normality of errors and treatment exogeneity, the lower bound for the size of the effect of having higher household income from sources other than the wife’s labor on their labor supply indicates an effect larger than -4.19 hours worked per year for married women.
@@
 The results are shown to extend to many other popular Tobit-type “two-part” models.
-By converting these sharp equalities into conditional moment inequalities, we propose a testing procedure that detects violations of the Tobit model assumptions on a grid on the joint support of the outcome (and treatment) variables leveraging inference results from Chernozhukov, Lee, and Rosen (2013) and the implementation from Chernozhukov et al.
-(2015).
+By converting these sharp equalities into conditional moment inequalities, we propose a testing procedure that detects violations of the Tobit model assumptions on a grid on the joint support of the outcome (and treatment) variables, leveraging inference results from chernozhukov_intersection_2013 and the implementation from chernozhukov_implementing_2015.
 Simulation results suggest the test performs well for reasonably sized samples (larger than 5000 observations).
-The test is conservative for smaller samples, overejecting the null hypothesis of model validity.
+The test is conservative for smaller samples, over-rejecting the null hypothesis of model validity.
 Simulations indicate that the test is powerful to detect violations of the exogeneity assumption for the treatment/instrument that affect the point estimates and inference.
@@
 We provide a brief review of existing models that could be implemented under weak/alternative assumptions when the Tobit model is rejected.
-Furthermore, we propose a simple model that partially identifies the parameter of interest by relying solely on linear index and monotone treatment selection restrictions, a standard assumption from the partial identification literature (Manski and Pepper, 2000).
-We illustrate our methods on data from Lee (1995).
+Furthermore, we propose a simple model that partially identifies the parameter of interest by relying solely on linear index and monotone treatment selection restrictions, a standard assumption from the partial identification literature (manski2000amonotone).
+We illustrate our methods on data from lee1995semi.
 We replicate qualitatively the results in the original paper and the proposed test for validity of the IV-Tobit model rejects the null hypothesis in this empirical application.
@@
 Appendix A Sharpness for the classic Tobit case
-Let $\alpha_{1},\alpha_{0}$ be identified.
-Suppose 3.1- 3.3 hold.
-Let $f_{\tilde{U}|D}=\frac{1}{\sqrt{2\pi}}e^{-1/2\tilde{U}^{2}}$.
-Let $\tilde{Y}^{*}=\alpha_{0}+\alpha_{1}D+\tilde{U}$ and $\tilde{Y}=\max(0,\tilde{Y}^{*})$.
-From $f_{\tilde{U}|D}$ we can see assumption 1 holds, since $f_{\tilde{U}|D}=f_{\tilde{U}}$.
-Also note $f_{\tilde{U}}$ is the $N(0,1)$ density, thus assumption 2 holds.
-Considering the positive values for $\tilde{Y}$, for any constants $c_{0},c_{1}$ such that $0<c_{0}<c_{1}$, then,
+[…]
 $P(c_{0}\leq\tilde{Y}\leq c_{1}|D=d)$ $=$ $P(c_{0}\leq\tilde{Y}^{*}\leq c_{1}|D=d)$
-$=$ $P(c_{0}\leq\alpha_{0}+\alpha_{1}d+\tilde{U}\leq c_{1}|D=d)$
-$=$ $P(c_{0}-\alpha_{0}-\alpha_{1}d\leq\tilde{U}\leq c_{1}-\alpha_{0}-\alpha_{1}d|D=d)$
-$=$ $P(c_{0}-\alpha_{0}-\alpha_{1}d\leq\tilde{U}\leq c_{1}-\alpha_{0}-\alpha_{1}d)$
-$=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)$
+$=$ $P(c_{0}\leq\tilde{\alpha}_{0}+\tilde{\alpha}_{1}d+\tilde{U}\leq c_{1}|D=d)$
+$=$ $P(c_{0}-\tilde{\alpha}_{0}-\tilde{\alpha}_{1}d\leq\tilde{U}\leq c_{1}-\tilde{\alpha}_{0}-\tilde{\alpha}_{1}d|D=d)$
+$=$ $P\left(\frac{c_{0}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}d\leq\frac{\tilde{U}}{\sigma}\leq\frac{c_{1}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}d\right)$
+$=$ $\Phi\left(\frac{c_{1}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}d\right)-\Phi\left(\frac{c_{0}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}d\right)$
 $=$ $P(c_{0}\leq Y\leq c_{1}|D=d).$
@@
 The fourth and fifth steps use the specific choice for the probability density of $\tilde{U}$, which implies independence from $D$ and normality respectively.
-The last step uses the relationship between the observed data and the model, established in 3.1.
-Similar reasoning applies for $P(c_{1}\leq\tilde{Y}|D=d)=P(c_{1}\leq\alpha_{0}+\alpha_{1}d+\tilde{U}|D=d)=1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)=P(c_{1}\leq Y|D=d)$.
+The last step uses the relationship between the observed data and the model, established in Equation (4.1).
 Furthermore, for $P(\tilde{Y}=0|D=d)$:
-$P(\tilde{Y}=0|D=d)$ $=$ $P(\tilde{Y}^{*}\leq 0|D=d)=P(\alpha_{0}+\alpha_{1}d+\tilde{U}\leq 0|D=d)$
-$=$ $P(\tilde{U}\leq-\alpha_{0}-\alpha_{1}d|D=d)=P(\tilde{U}\leq-\alpha_{0}-\alpha_{1}d)$
-$=$ $\Phi(-\alpha_{0}-\alpha_{1}d)=1-\Phi(\alpha_{0}+\alpha_{1}d)=P(Y=0|D=d)$
-The last step uses the equality established in 3.3.
+$P(\tilde{Y}=0|D=d)$ $=$ $P(\tilde{Y}^{*}\leq 0|D=d)=P(\tilde{\alpha}_{0}+\tilde{\alpha}_{1}d+\tilde{U}\leq 0|D=d)$
+$=$ $P(\tilde{U}\leq-\tilde{\alpha}_{0}-\tilde{\alpha}_{1}d|D=d)=P(\tilde{U}\leq-\tilde{\alpha}_{0}-\tilde{\alpha}_{1}d)$
+$=$ $\Phi\left(-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}d\right)=1-\Phi\left(\frac{\tilde{\alpha}_{0}}{\sigma}+\frac{\tilde{\alpha}_{1}}{\sigma}d\right)=P(Y=0|D=d)$
+The last step uses the equality established in Equation 4.2.
 Thus, we characterized the distribution of $\tilde{Y}|D$, which is equal to the one of $Y|D$, with $D$ given.
@@
 Appendix B Details of the IV-Tobit case
-B.1. Derivation of equation 3.4
-$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z)$ (B.1)
+B.1. Derivation of equation (4.3)
+$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z)$
 $=$ $P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1},Y^{*}\geq 0|Z=z)+P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1},Y^{*}<0|Z=z)$
 $=$ $P(c_{0}\leq Y^{*}\leq c_{1},d_{0}\leq D\leq d_{1},Y^{*}\geq 0|Z=z)$
-$=$ $P(c_{0}\leq\beta_{0}+\beta_{1}z+W\leq c_{1},d_{0}\leq\gamma_{0}+\gamma_{1}z+V\leq d_{1},Y^{*}\geq 0|Z=z)$
-$=$ $P(c_{0}-\beta_{0}-\beta_{1}z\leq W\leq c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq V\leq d_{1}-\gamma_{0}-\gamma_{1}z,Y^{*}\geq 0|Z=z)$
-$=$ $P(c_{0}-\beta_{0}-\beta_{1}z\leq W\leq c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq V\leq d_{1}-\gamma_{0}-\gamma_{1}z)$
-$=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)-\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
--- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)+\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
+$=$ $P(c_{0}\leq\tilde{\beta}_{0}+\tilde{\beta}_{1}z+\tilde{W}\leq c_{1},d_{0}\leq\tilde{\gamma}_{0}+\tilde{\gamma}_{1}z+\tilde{V}\leq d_{1},Y^{*}\geq 0|Z=z)$
+$=$ $P(c_{0}-\tilde{\beta}_{0}-\tilde{\beta}_{1}z\leq\tilde{W}\leq c_{1}-\tilde{\beta}_{0}-\tilde{\beta}_{1}z,d_{0}-\tilde{\gamma}_{0}-\tilde{\gamma}_{1}z\leq\tilde{V}\leq d_{1}-\tilde{\gamma}_{0}-\tilde{\gamma}_{1}z,Y^{*}\geq 0|Z=z)$
+$=$ $P(c_{0}-\tilde{\beta}_{0}-\tilde{\beta}_{1}z\leq\tilde{W}\leq c_{1}-\tilde{\beta}_{0}-\tilde{\beta}_{1}z,d_{0}-\tilde{\gamma}_{0}-\tilde{\gamma}_{1}z\leq\tilde{V}\leq d_{1}-\tilde{\gamma}_{0}-\tilde{\gamma}_{1}z)$
+$=$ $\Phi_{W,V}\left(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{1}}{\sigma_{\tilde{V}}}z;\rho\right)-\Phi_{W,V}\left(\frac{c_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{1}}{\sigma_{\tilde{V}}}z;\rho\right)$
+- $\Phi_{W,V}\left(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{1}}{\sigma_{\tilde{V}}}z;\rho\right)+\Phi_{W,V}\left(\frac{c_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{1}}{\sigma_{\tilde{V}}}z;\rho\right)$
 The first equality follows from the law of total probability.
@@
 The final equality is by the properties of probabilities and the joint normality for $W,V$ (Assumption 4).
-B.2. Derivation of equation 3.10
-By a similar approach to the derivation of 3.4:
+B.2. Derivation of equation (4.9)
+By a similar approach to the derivation of (4.3):
 $P(Y=0,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $P(Y^{*}\leq 0,d_{0}\leq D\leq d_{1}|Z=z)$ (B.1)
-$=$ $P(W\leq-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}\leq V\leq d_{1}-\gamma_{0}-\gamma_{1}z)$
-$=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
--- $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
+$=$ $P(\tilde{W}\leq-\tilde{\beta}_{0}-\tilde{\beta}_{1}z,d_{0}-\tilde{\gamma}_{0}-\tilde{\gamma}_{1}\leq\tilde{V}\leq d_{1}-\tilde{\gamma}_{0}-\tilde{\gamma}_{1}z)$
+$=$ $\Phi_{W,V}\left(-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{1}}{\sigma_{\tilde{V}}}z;\rho\right)$
+- $\Phi_{W,V}\left(-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{1}}{\sigma_{\tilde{V}}}z;\rho\right)$
 B.3. Proof of Sharpness
-Let $\beta_{0},\beta_{1},\gamma_{0},\gamma_{1},\rho$ be identified and equalities 3.4-3.12 hold. Define the joint density of $\left(\tilde{W},\tilde{V}\right)$ conditional on $Z$ as
+[…]
 $f_{\left(\tilde{W},\tilde{V}|Z\right)}(w,v|z)=\frac{1}{\sqrt{1-\rho^{2}}}\phi\left(\frac{w-\rho v}{\sqrt{1-\rho^{2}}}\right)\phi(v),$
@@
 $\left\{\begin{array}[]{lcl}\tilde{Y}&=&\max(0,\tilde{Y}^{*})\\ \\ \tilde{Y}^{*}&=&\beta_{0}+\beta_{1}Z+\tilde{W}\\ \\ \tilde{D}&=&\gamma_{0}+\gamma_{1}Z+\tilde{V}\end{array}\right$.
-Note that $f_{\left(\tilde{W},\tilde{V}|Z\right)}=f_{\left(\tilde{W},\tilde{V}\right)}$, thus assumption 3 holds.
-Similarly, $\left(\tilde{W},\tilde{V}\right)$ follow a bivariate normal distribution as $\begin{pmatrix}\tilde{W}\\ \tilde{V}\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\end{pmatrix}$, and $\Sigma=\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}$.
-Let $\tilde{U}=\tilde{W}-\alpha_{1}\tilde{V}$, which implies $\tilde{U},\tilde{V}$ satisfies assumption 4, given the scale-location normalizations.
+[…]
 Then, for any constants $0<c_{0}\leq c_{1}$ and $d_{0}<d_{1}$:
@@
 $=$ $P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z)$
-The first through third equalities follow from the definitions of $\tilde{Y},\tilde{Y}^{*}$ and $\tilde{D}$.
-The fourth and fifth steps are consequences of the particular choice for joint density for $\left(\tilde{W},\tilde{V}\right)$ conditional on $Z$.
-The final equality is given by the relationship between observable data and the latent model structure in Equation 3.4.
-Similar derivations hold for 3.5-3.9.
+[…]
 For the accumulation point, at $Y=0$:
@@
 The steps of the proof are similar to previous cases and the last equality is given by the relationship between observable data and the latent model structure in Equation 4.9.
-Similar derivations hold for equalities 3.11-3.12.
+Similar derivations hold for equalities (4.10)-(4.11).
 Thus, we characterized the distribution of $(\tilde{Y},\tilde{D})|Z$, which coincides with the joint distribution of $(Y,D)|Z$, for given $Z$.
@@
 Appendix C Type 2 Tobit testable implications
-In this section we derive the results of the Type 2 tobit model also known as selection models or Heckman selection type of models (Heckman, 1979).
-The basic setup with no covariates (which can be extended in several directions and with different distributional assumptions as well as to incorporate treatment endogeneity, as we pointed out for the classic Tobit model in remark 3) is:
-$\left\{\begin{array}[]{lcl}Y&=&Y^{*}\quad\text{if}\quad S=1\\ \\ Y&=&\text{missing}\quad\text{if}\quad S=0\\ \\ Y^{*}&=&\alpha_{0}+\alpha_{1}D+U\\ \\ S&=&1\{\gamma_{0}+\gamma_{1}Z+V\geq 0\}\end{array}\right.$ (C.8)
+In this section, we derive the results of the Type 2 tobit model, also known as selection models or Heckman selection-type models (heckman1979sample).
+The basic setup with no covariates (which can be extended in several directions and with different distributional assumptions as well as to incorporate treatment endogeneity, as we pointed out for the classic Tobit model in proposition 2) is:
+$\left\{\begin{array}[]{lcl}Y&=&Y^{*}\quad\text{if}\quad S=1\\ \\ Y&=&\text{missing}\quad\text{if}\quad S=0\\ \\ Y^{*}&=&\alpha_{0}+\alpha_{1}D+U\\ \\ S&=&1\{\gamma_{0}+\gamma_{1}Z+V\geq 0\}\end{array}\right$.
+(C.8)
 Where $U,V$ are the latent structural error terms.
@@
 Let $U,V$ follow a bivariate normal distribution with covariance $\rho$, i.e., $\begin{pmatrix}U\\ V\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma=\begin{pmatrix}\sigma^{2}_{U}&\rho_{UV}\\ \rho_{UV}&\sigma^{2}_{V}\end{pmatrix}$.
-Note that $Y$ is missing at $S=0$. Thus, fully characterizing the distribution implies observing the behaviour at the missing point and beyond it.
+Note that $Y$ is missing at $S=0$. Thus, fully characterizing the distribution implies observing the behavior at the missing point and beyond it.
 From the observed data the conditional probabilities of $c_{0}\leq Y\leq c_{1}$ for an $c_{1},c_{0}$ can be computed. Then note for any $d\in\mathcal{D}$ and $z\in\mathcal{Z}$ :
@@
 $=$ $\Phi_{v}\left(\frac{-\gamma_{0}-\gamma_{1}z}{\sigma_{V}}\right)$
-Where the first and the second equalities are due to the structure of the model described in equation C.8.
+Where the first and the second equalities are due to the structure of the model described in equation (C.8).
 The third step is due to Assumption 7 and the last one follows from the normalization for normal random variables and Assumption 8.
@@
 $=$ $P(c_{0}-\alpha_{0}-\alpha_{1}d\leq U\leq c_{1}-\alpha_{0}-\alpha_{1}d,V\geq-\gamma_{0}-\gamma_{1}z)$
-$=$ $\Phi_{U}(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi_{U}(c_{0}-\alpha_{0}-\alpha_{1}d)$
--- $\Phi_{U,V}(c_{1}-\alpha_{0}-\alpha_{1}d,-\gamma_{0}-\gamma_{1}z,\rho_{UV})$
-$+$ $\Phi_{U,V}(c_{0}-\alpha_{0}-\alpha_{1}d,-\gamma_{0}-\gamma_{1}z,\rho_{UV})$
+$=$ $\Phi_{U}\left(\frac{c_{1}-\alpha_{0}-\alpha_{1}d}{\sigma_{U}}\right)-\Phi_{U}\left(\frac{c_{0}-\alpha_{0}-\alpha_{1}d}{\sigma_{U}}\right)$
+- $\Phi_{U,V}\left(\frac{c_{1}-\alpha_{0}-\alpha_{1}d}{\sigma_{U}},\frac{-\gamma_{0}-\gamma_{1}z}{\sigma_{V}},\rho_{UV}\right)$
+$+$ $\Phi_{U,V}\left(\frac{c_{0}-\alpha_{0}-\alpha_{1}d}{\sigma_{U}},\frac{-\gamma_{0}-\gamma_{1}z}{\sigma_{V}},\rho_{UV}\right)$
 The first equality follows from the law of total probability.
-The second through fifth equalities follow from the model’s structure described in Equation C.8, namely $P(Y>0,Y^{*}<0|D=d)=0$ and $Y^{*}=Y,\text{not missing}$ for $S=1$, the structure of $S$, and the latent linear model.
+The second through fifth equalities follow from the model’s structure described in Equation (C.8), namely $P(Y>0,Y^{*}<0|D=d)=0$ and $Y^{*}=Y,\text{not missing}$ for $S=1$, the structure of $S$, and the latent linear model.
 The sixth step is due to Assumption 7.
@@
 Thus, we can construct a test similarly to the one in the main text, using these equalities in addition to the ones of the form $P(Y\geq c_{2}|D=d,Z=z)$.
-Appendix D Results from Remark 3
-In this appendix we derive the testable implications of the extensions to the classic Tobit and IV-Tobit discussed in Remark 3.
+Appendix D Results from Proposition 2
+In this appendix, we derive the testable implications of the extensions to the classic Tobit and IV-Tobit discussed in Proposition 2.
 We discuss testable implications assuming the identification of the relevant parameters or distributions.
-D.1. Barros et al. (2018) and Carson and Sun (2007)
-Barros et al.
-(2018) proposes a variant of the tobit model with elliptically contoured distributions and a non-zero threshold.
-At the same time Carson and Sun (2007) proposes a Tobit model with a non-zero threshold.
+D.1. barros2018generalized and carson2007Tobit
+barros2018generalized proposes a variant of the tobit model with elliptically contoured distributions and a non-zero threshold.
+At the same time carson2007Tobit proposes a Tobit model with a non-zero threshold.
 In this section we combine both types of results and report the testable implications with a generic non-zero threshold and a generic known or identifiable parametric distribution function.
@@
 Also, note it is possible to extend the previous development to cases where the truncation takes the form $Y=Y^{*}$ if $\tau_{0}\leq Y^{*}\leq\tau_{1}$, $Y=\tau_{0}$ if $\tau_{0}\geq Y^{*}$ and $Y=\tau_{1}$ if $Y^{*}\geq\tau_{1}$.
-D.2. Wooldridge (2005); Honore, Kyriazidou, and Powell (2000); Honoré (1993)
-Consider the following dynamic version of the Tobit model, which is related to Wooldridge (2005); Honore, Kyriazidou, and Powell (2000); Honoré (1993).
-Here, we specify the conditional behaviour of $c_{i}$ in the spirit of Wooldridge (2005) and others.
+D.2. wooldridge2005simple; honore2000estimation; honore1993orthogonality
+Consider the following dynamic version of the Tobit model, which is related to wooldridge2005simple; honore2000estimation; honore1993orthogonality.
+Here, we specify the conditional behaviour of $c_{i}$ in the spirit of wooldridge2005simple and others.
 $\left\{\begin{array}[]{lcl}Y_{i,t}&=&\max(0,Y_{i,t}^{*})\\ \\ Y_{i,t}^{*}&=&\alpha_{0}+\alpha_{1}D_{i,t}+\alpha_{2}g(Y_{i,t-1})+c_{i}+U_{i,t}\end{array}\right$. (D.10)
@@
 $=$ $\int P(c_{0}-\alpha_{0}-\alpha_{1}d-\alpha_{2}g(y)-c_{i}\leq U_{i,t}\leq c_{1}-\alpha_{0}-\alpha_{1}d-\alpha_{2}g(y)-c_{i}|Y_{i,t-1}=y,c_{i})f(c_{i}|d,y)dc_{i}$
-[…]
+$=$ $\int[\Phi(c_{1}-\alpha_{0}-\alpha_{1}d-\alpha_{2}g(y)-c_{i})-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d-\alpha_{2}g(y)-c_{i})]\phi(c_{i})dc_{i}$
 Similarly,
@@
 $=$ $\int P(U_{i,t}\leq-\alpha_{0}-\alpha_{1}d-\alpha_{2}g(y)-c_{i}|Y_{i,t-1}=y,c_{i})f(c_{i}|d,y)dc_{i}$
-[…]
+$=$ $\int[\Phi(-\alpha_{0}-\alpha_{1}d-\alpha_{2}g(y)-c_{i})]\phi(c_{i})dc_{i}$
 Which can then be used to construct a test in a similar way as presented in the main text by adding the type of equalities $P(Y_{i,t}\geq c_{2}|D_{i,t}=d,Y_{i,t-1}=y)$, with the caveat that the right-hand side requires numerical integration or an approximation by an estimator when no closed form is available.
-Such estimator should ensure that the left-hand side of the equality converges at root-$N$ in order for the estimation step of the null model to be asymptotically negligible (Acerenza, Bartalotti, and Kédagni, 2023, Appendix B).
+Such estimator should ensure that the left-hand side of the equality converges at root-$N$ in order for the estimation step of the null model to be asymptotically negligible (acerenza_testing_2023, Appendix B).
 Appendix E Non-learnability
@@
 One can show that by finding an alternative model that is compatible with the equalities (4.1)-(4.2) in all cases in which the Tobit model could not be disregarded.
-Suppose that,
-$\left\{\begin{array}[]{lcl}\bar{Y}&=&\max(0,\bar{Y}^{*})\\ \\ \bar{Y}^{*}&=&G(\alpha_{0}+\alpha_{1}D+\bar{U})\end{array}\right.$ (E.4)
-Where $G(.)$ is some continuous and strictly increasing function such that $G^{-1}(0)=0$ and $\bar{U}$ is a random variable that is $N(0,1)$.
-For $c_{0}\leq Y\leq c_{1}$ for a $c_{1},c_{0}>0$ and for any value of the treatment variables $d\in\mathcal{D}$, assume that the equalities (3.1)-(3.3) hold:
+Thus the model is non-verifiable in the sense that we can always construct a joint probability law of $(\tilde{Y},D,\tilde{U}$) that violates the Tobit model validity but satisfies the equalities.
+Concretely, for the classic Tobit model suppose that,
+$\left\{\begin{array}[]{lcl}\bar{Y}&=&\max(0,\bar{Y}^{*})\\ \\ \bar{Y}^{*}&=&\alpha_{0}+\alpha_{1}D+\tilde{U}\end{array}\right$. (E.4)
+$\tilde{U}$ is a random variable that has the following mixed discrete-continuous “density”:
+$f_{\tilde{U}|D}(\tilde{u}|d)=\begin{cases}\Phi(\hat{u})(\frac{e^{d}}{1+e^{d}})\text{ if }\tilde{u}=\hat{u}\\ \Phi(-\alpha_{0}-\alpha_{1}d)-\Phi(\hat{u})(\frac{e^{d}}{1+e^{d}})\text{ if }\tilde{u}=-\alpha_{0}-\alpha_{1}d\\ \phi(\tilde{u})\text{ if }\tilde{u}\geq-\alpha_{0}-\alpha_{1}d\\ 0,\text{ otherwise}\end{cases}$
+[…]
+For $c_{0}\leq Y\leq c_{1}$ for a $c_{1},c_{0}>0$ and for any value of the treatment variables $d\in\mathcal{D}$, assume that the sharp equalities hold:
 $P(c_{0}\leq Y\leq c_{1}|D=d)$ $=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)$
-$P(c_{1}\leq Y|D=d)$ $=$ $1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)$
 $P(Y=0|D=d)$ $=$ $1-\Phi(\alpha_{0}+\alpha_{1}d).$
-Now define $\bar{Y}_{d}=G(\alpha_{0}+\alpha_{1}d+\bar{U})$ and define the following distribution function:
-$P(\bar{Y}_{d}\leq c_{1},\bar{U}\leq u|D=d)=\Phi(\min\{(G^{-1}(c_{1})-\alpha_{0}-\alpha_{1}d),u\})$
-In the above proposed DGP, the latent model is not linear. However, the DGP is compatible with the data since,
-$P(\bar{Y}_{d}\leq 0)$ $=$ $P(\tilde{U}\leq-\alpha_{0}-\alpha_{1}d)=\Phi(-\alpha_{0}-\alpha_{1}d)$ (E.5)
-$=$ $1-\Phi(\alpha_{0}+\alpha_{1}d)=P(Y=0|D=d)$
-The last equality is due to the testable equalities holding. Similarly, for any $c_{1}>0$, we can get that
-$P(G(c_{1})<\bar{Y}_{d})$ $=$ $1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)=P(c_{1}<Y|D=d).$ (E.6)
-And,
-$P(G(c_{0})\leq\bar{Y}_{d}\leq G(c_{1}))$ $=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)$ (E.7)
+Now define $\bar{Y}_{d}=\alpha_{0}+\alpha_{1}d+\tilde{U}$
+In the above proposed DGP, the latent model is not normal. However, the DGP is compatible with the data and the testable implications derived in Section 4 since,
+$P(\bar{Y}_{d}<0)$ $=$ $P(\tilde{U}\leq-\alpha_{0}-\alpha_{1}d)$ (E.5)
+$=$ $\Phi(\hat{u})\left(\frac{e^{d}}{1+e^{d}}\right)+\Phi(-\alpha_{0}-\alpha_{1}d)-\Phi(\hat{u})\left(\frac{e^{d}}{1+e^{d}}\right)$ (E.6)
+$=$ $\Phi(-\alpha_{0}-\alpha_{1}d)=1-\Phi(\alpha_{0}+\alpha_{1}d)$ (E.7)
+$=$ $P(Y=0|D=d)$ (E.8)
+The last equality is due to the testable equalities holding. Similarly for any $c_{0},c_{1}>0$,
+$P(c_{0}\leq\bar{Y}_{d}\leq c_{1})$ $=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)$ (E.9)
 $=$ $P(c_{0}\leq Y\leq c_{1}|D=d).$
 Hence, when the sharp equalities hold, there exists an alternative observationally equivalent model, to the classic tobit that can induce the observed data distribution.
-Indeed, the observed distribution of $Y$ and the non refuted marginal distribution of $U$, do not imply that we can learn the true distribution of $Y_{1},Y_{0},U$.
+Indeed, the observed distribution of $Y$ and the non refuted marginal distribution of $U$, do not imply that we can learn the true distribution of ($Y_{1},Y_{0},U$).
 E.2. Non-Learnability for the IV- Tobit model
-As mentioned in Remark 5, the IV-Tobit model is non-verifiable, that is, we cannot learn if the maintained model is the true data generating process based on the sharp equalities proposed in Section 3.2.
+As mentioned in proposition 3, the IV-Tobit model is non-verifiable, that is, we cannot learn if the maintained model is the true data generating process based on the sharp equalities proposed in Section 4.2.
 One can show that by finding an alternative model that is compatible with the equalities (4.3)-(4.11) in all cases in which the IV-Tobit model could not be disregarded.
 Suppose that,
-$\left\{\begin{array}[]{lcl}\tilde{Y}&=&\max(0,\tilde{Y}^{*})\\ \\ \tilde{Y}^{*}&=&G(\beta_{0}+\beta_{1}Z+\tilde{W})\\ \\ \tilde{D}&=&\gamma_{0}+\gamma_{1}Z+\tilde{V}\end{array}\right.$ (E.13)
-Where $G$ is the same function as in the previous subsection. Assume that equations 3.4-3.12 hold:
-$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
--- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
--- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
-$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
-$P(c_{0}\leq Y\leq c_{1},D\leq d_{0}|Z=z)$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
--- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
-$P(c_{0}\leq Y\leq c_{1},D\geq d_{1}|Z=z)$ $=$ $\Phi_{W}(c_{1}-\beta_{0}-\beta_{1}z)-\Phi_{W}(c_{0}-\beta_{0}-\beta_{1}z)$
--- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
-$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
-$P(c_{1}\leq Y,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z)-\Phi_{V}(d_{0}-\gamma_{0}-\gamma_{1}z)$
--- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
-$+$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
-$P(c_{1}\leq Y,D\leq d_{0}|Z=z)$ $=$ $\Phi_{V}(d_{0}-\gamma_{0}-\gamma_{1}z)$
--- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
-$P(c_{1}\leq Y,D\geq d_{1}|Z=z)$ $=$ $1-\Phi_{W}(c_{1}-\beta_{0}-\beta_{1}z)-\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z)$
-$+$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
-$P(Y=0,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
--- $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
-$P(Y=0,D\leq d_{0}|Z=z)$ $=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho),$
-$P(Y=0,D\geq d_{1}|Z=z)$ $=$ $\Phi_{W}(-\beta_{0}-\beta_{1}z)-\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
-Define $\bar{Y}_{z}=G(\beta_{0}+\beta_{1}z+\tilde{W}),\tilde{D}_{z}=\gamma_{0}+\gamma_{1}z+\tilde{V}$ and define the following distribution function:
-$P(\bar{Y}_{d}\leq c_{1},\tilde{D}_{z}\leq d_{1},\bar{U}\leq u,\bar{V}\leq v|D=d)=\Phi(\min\{(G^{-1}(c_{1})-\beta_{0}-\beta_{1}z),u\},\min\{d_{1}-\gamma_{0}-\gamma_{1}z,v\};\rho)$
-Now note that, for example,
-$P(G(c_{0})\leq\tilde{Y}_{z}\leq G(c_{1}),d_{0}\leq\tilde{D}_{z}\leq d_{1})$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$ (E.14)
--- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
--- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
-$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
-$=$ $P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z)$
-A similar logic holds for the other equalities. But the model is not the IV-tobit model.
-Appendix F Simulations for test’s power under violations of normality
-Turning our attention to the test’s power under violations of normality of the latent error, we consider the modified DGP in Equation F.7.
-In this DGP, $F(\cdot)$ denotes an alternative density for the latent error $U$, which we vary across scenarios.
-Table 6 present the test’s empirical coverage when $F(\cdot)$ is uniform, lognormal and t-student distributions.
-$\left\{\begin{array}[]{lcl}Y&=&\max\{Y^{*},0\}\\ Y^{*}&=&D+U\\ D&=&2Z-V\\ \begin{pmatrix}V\\ Z\end{pmatrix}&\sim&\mathcal{N}(\mathbf{0}_{2},\mathbf{\Sigma})\\ \mathbf{\Sigma}&=&\begin{pmatrix}1&0\\ 0&1\\ \end{pmatrix}\\ U\sim F(\cdot)\end{array}\right.$ (F.7)
-As expected, the rejection rate for the case of the t-student with 80 degrees of freedom is close to the nominal size of the test, since it represents a very mild violation of the normality assumption.
-In the other three cases, which deviate significantly from the normality assumption, the test rejects the null hypothesis in most instances.
-This example indicates that the proposed test is useful in detecting violations of the distributional assumptions in the classic Tobit model.
-Table 6. Test power for violations in error structure
-$N$ $U$ MSE$(\beta_{0})$ MSE$(\beta_{1})$ $\alpha=10\%$ $\alpha=5\%$ $\alpha=1\%$
-5000 t-student (df=80) 0.9987 1.0018 7.21% 4.01% 1.40%
-t-student (df=5) 1.027 1.0624 92.20% 81.80% 48.80%
-LogNormal 1.3336 0.0896 100.00% 100.00% 100.00%
-Uniform 0.892 0.785 100.00% 100.00% 100.00%
-8000 t-student (df=80) 0.9999 1.0034 7.00% 4.80% 1.20%
-t-student (df=5) 1.0251 1.0653 99.20% 97.80% 87.58%
-LogNormal 1.3385 0.0866 100.00% 100.00% 100.00%
-Uniform 0.8932 0.7896 100.00% 100.00% 100.00%
-10000 t-student (df=80) 1.0016 1.0059 5.42% 2.41% 1.00%
-t-student (df=5) 1.0241 1.0642 99.20% 98.40% 90.58%
-LogNormal 1.3348 0.0868 100.00% 100.00% 100.00%
-Uniform 0.8908 0.7862 100.00% 100.00% 100.00%
-Based on $500$ replications.
+$\left\{\begin{array}[]{lcl}\tilde{Y}&=&\max(0,\tilde{Y}^{*})\\ \\ \tilde{Y}^{*}&=&\beta_{0}+\beta_{1}Z+\tilde{W}\\ \\ \tilde{D}&=&\gamma_{0}+\gamma_{1}Z+\tilde{V}\end{array}\right$. (E.15)
+$\tilde{W},\tilde{V}$ are random variables that have the next mixed discrete-continuous behavior:
+[…]
+[…]
+$f_{\tilde{W}|\tilde{V},Z}(\tilde{w}|\tilde{v},z)=\begin{cases}\Phi(\hat{w}|\tilde{v};\rho)(\frac{e^{z}}{1+e^{z}})\text{ if }\tilde{w}=\hat{w}\\ \Phi(-\beta_{0}-\beta_{1}z|\tilde{v};\rho)-\Phi(\hat{w}|\tilde{v};\rho)(\frac{e^{z}}{1+e^{z}})\text{ if }\tilde{w}=-\beta_{0}-\beta_{1}z\\ \phi(\tilde{w}|\tilde{v};\rho)\text{ if }\tilde{w}\geq-\beta_{0}-\beta_{1}z\\ 0,\text{ otherwise.}\end{cases}$
+Where $\hat{w}$ is some real number strictly less than $\beta_{0}-\beta_{1}z$, $\phi(\tilde{w}|\tilde{v},\rho)$ is the normal p.d.f.
+conditional on $\tilde{v}$ with correlation coefficient $\rho$ and $\Phi(\tilde{w}|\tilde{v};\rho)$ is the normal c.d.f.
+analog.
+Note that, $\Phi(\hat{w}|\tilde{v};\rho)(\frac{e^{z}}{1+e^{z}})>0,\phi(\tilde{w}|\tilde{v};\rho)>0$ and $\Phi(-\beta_{0}-\beta_{1}z|\tilde{v};\rho)-\Phi(\hat{w}|\tilde{v};\rho)(\frac{e^{z}}{1+e^{z}})>0$.
+This distribution assigns probability mass to two points, $\hat{w}$ and $\beta_{0}-\beta_{1}z$ while assigning the conditional normal density above $\beta_{0}-\beta_{1}z$.
+Thus, above $\beta_{0}-\beta_{1}z$, $\tilde{W}$ behaves like a conditional normal random variable, but below behaves like a discrete random variable with two mass points.
+Note that, with the previous structure, for values of $\tilde{w}$ less than $\hat{w}$:
+[…]
+[…]
+$=$ $\int_{-\infty}^{\tilde{v}}P(\tilde{W}\leq\hat{w}|\tilde{V}=v)\phi(v)dv\frac{e^{z}}{1+e^{z}}$
+Furthermore, note that:
+$\int_{-\infty}^{\tilde{v}}\Phi(\hat{w}|\tilde{v};\rho)\phi(\tilde{v})d\tilde{v}$ $=$ $\int_{-\infty}^{\tilde{w}}\int_{-\infty}^{\tilde{v}}\phi(\hat{w}|\tilde{v};\rho)\phi(\tilde{v})d\tilde{v}d\tilde{w}=\int_{-\infty}^{\tilde{w}}\int_{-\infty}^{\tilde{v}}\phi(\hat{w},\tilde{v};\rho)d\tilde{v}d\tilde{w}$
+$=$ $\Phi(\tilde{w},\tilde{v};\rho)$
+With this, we can then say that:
+$F_{\tilde{W},\tilde{V}|Z}(\tilde{w},\tilde{v}|z)=\begin{cases}\Phi(\hat{w},\tilde{v};\rho)\frac{e^{z}}{1+e^{z}}\text{ if }\tilde{w}=\hat{w},-\infty<\tilde{v}\\ \Phi(-\beta_{0}-\beta_{1}z,\tilde{v};\rho)\text{ if }\hat{w}<\tilde{w}\leq-\beta_{0}-\beta_{1}z,-\infty<\tilde{v}\\ \Phi(\tilde{w},\tilde{v};\rho)\text{ if }\tilde{w}>-\beta_{0}+\beta_{1}z,-\infty<\tilde{v}\\ 0,\text{ otherwise}\end{cases}$
+This joint c.d.f., which integrates to $1$ when $\tilde{w}$ and $\tilde{v}$ goes to $\infty$ is associated with the following mixed discrete-continuous ”density”:
+$f_{\tilde{W},\tilde{V}|Z}(\tilde{w},\tilde{v}|z)=\begin{cases}\Phi(\hat{w},\tilde{v};\rho)\frac{e^{z}}{1+e^{z}}\text{ if }\tilde{w}=\hat{w},\forall\tilde{v}\\ \Phi(-\beta_{0}-\beta_{1}z,\tilde{v};\rho)-\Phi(\hat{w},\tilde{v};\rho)\frac{e^{z}}{1+e^{z}}\text{ if }\tilde{w}=-\beta_{0}-\beta_{1}z,\forall\tilde{v}\\ \phi(\tilde{w},\tilde{v};\rho)\text{ if }\tilde{w}>-\beta_{0}+\beta_{1}z,\forall\tilde{v}\end{cases}$
+Assume that equations 4.3-4.11 hold. Then, define $\bar{Y}_{z}=\beta_{0}+\beta_{1}z+\tilde{W},\tilde{D}_{z}=\gamma_{0}+\gamma_{1}z+\tilde{V}$
+$P(\bar{Y}_{z}<0,D_{z}\leq d_{0})$ $=$ $P(\tilde{W}\leq-\beta_{0}-\beta_{1}z,\tilde{V}\leq d_{0}-\gamma_{0}-\gamma_{1}z)$ (E.16)
+$=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$ (E.17)
+$=$ $P(Y=0,D\leq d_{0}|Z=z)$ (E.18)
+The last equality is due to the testable equalities holding.
+A similar display holds for any $c_{0},c_{1}>0,d_{1},d_{0}$.
+Thus a similar logic holds for the other equalities.
+So this is an example of a model that satisfies all the equalities, and is observationally equivalent to the IV-Tobit, even though the underlying latent error distributions does not satisfy the conditions for the IV-Tobit.
+Hence, we cannot falsify the assumed IV-Tobit model.